### v1

# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

Abstract
We establish a local topological obstruction to flattening Berry curvature in spin-orbit-coupled Bose-Einstein condensates (SOC BECs), valid even when the global Chern number vanishes. For a generic two-component SOC BEC, the extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ carries a natural metric connection $\nabla^{C}$ whose torsion 3-form encodes the synthetic gauge fields. Its harmonic part defines a mixed cohomology class
$\omega]\in\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
whose mixed tensor rank equals one. Using a general geometric bound for connections with skew torsion on product manifolds, we show that the obstruction kernel $\mathcal{K}$ vanishes, yielding the sharp inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$. This forces at least one off-diagonal curvature operator, preventing complete gauging-away of Berry phases even when the total Chern number is zero. This provides the first cohomological lower bound certifying locally irremovable curvature in SOC BECs beyond the Chern-number paradigm.
Keywords: Berry curvature; spin–orbit-coupled Bose–Einstein condensates; synthetic gauge fields; holonomy with torsion; mixed cohomology; Kaluza–Klein geometry.
MSC (2020): 53C29, 53C80, 81V45, 82B10, 51P05, 55S20, 57R19, 57R56, 58A12.
1 Introduction
The study of spin–orbit-coupled (SOC) Bose–Einstein condensates (BECs) stands at a rich confluence of differential geometry, quantum many-body physics, and synthetic gauge theory. This interdisciplinary effort is part of the broader pursuit of artificial gauge fields in engineered quantum systems, spanning ultracold atoms, photonics, and solid-state platforms [1]. Advances in generating artificial gauge fields with Raman lasers have made these systems a versatile laboratory for exploring Berry curvature, synthetic magnetic fields, and non-Abelian textures with unprecedented control [6, 7, 14]. These studies are often conducted in toroidal traps, which can serve as magnetic storage rings for BECs [3], providing a well-controlled environment for synthetic gauge field experiments. Remarkable achievements include the observation of Dirac monopoles in synthetic magnetic fields [10], demonstrating the high degree of geometric control now attainable. The geometric phase underpinning these phenomena traces back to the foundational work of Simon [13], while the experimental realization of artificial gauge fields has been comprehensively reviewed in [4].
A persistent theoretical challenge is to identify local topological obstructions that constrain the curvature even when global invariants—such as the first Chern number—vanish. While the Chern number quantizes the net flux, its vanishing does not guarantee that the curvature can be locally flattened; geometric obstructions may prevent its simultaneous annihilation along independent directions in the parameter space.
This work addresses that challenge by bringing a recent geometric result into physical focus. In [12], a lower bound was established for the off-diagonal holonomy of metric connections with totally skew-symmetric torsion on product manifolds. The bound is expressed in terms of a mixed de Rham cohomology class $\omega]\in H^{p}(M_{1})\otimes H^{q}(M_{2})$ associated with the torsion; its mixed tensor rank $r$ counts the minimal number of simple tensors needed to represent the class, and after subtracting contributions from parallel forms (which do not generate off-diagonal curvature) one obtains a corrected integer $r^{\sharp}$ that bounds the dimension of the off-diagonal holonomy algebra:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim\mathcal{K}.$
Here we show how this abstract geometric theorem materializes in a concrete physical setting: a generic two-component SOC BEC confined in a quasi-two-dimensional toroidal trap. Such a system possesses three universal, experimentally accessible degrees of freedom: the two crystal momenta forming a Brillouin-zone torus $T^{2}_{\mathrm{BZ}}$, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin components [9, 11]. We therefore take as the extended parameter space the compact product manifold.
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}.$
The first main task of the paper is to construct, in this space, a natural geometric structure to which the theorem of [12] applies. We introduce a metric $g_{M}$ of Kaluza–Klein type, which couples the base metric on $T^{2}_{\mathrm{BZ}}$ to the $U(1)$-valued connection forms that encode the synthetic gauge fields of the SOC BEC. On $(M,g_{M})$ we define a metric connection $\nabla^{C}$ whose torsion three-form is precisely the physical three-form
$T^{\flat}=\operatorname{Tr}(\mathcal{F}\wedge\mathbf{A}_{\mathrm{syn}}),$
built from the synthetic Berry curvature $\mathcal{F}$ and connection $\mathbf{A}_{\mathrm{syn}}$. The harmonic part of $T^{\flat}$ with respect to $g_{M}$ then defines a mixed cohomology class
$\omega]\in H^{3}(M)\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr).$
The second main task is to compute the topological data that enters the general bound. We prove that, under the generic condition that the two nearly degenerate bands carry non-zero Berry curvature classes $F^{(+)}],[F^{(-)}]\in H^{2}(T^{2}_{\mathrm{BZ}})$, the mixed tensor rank of $\omega]$ attains its maximal possible value within this geometry. Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, any two non-zero classes in this space are cohomologically proportional. Consequently, the mixed class $\omega]$ can be written as a single simple tensor
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
and its mixed tensor rank is therefore $r=1$. Moreover, owing to the specific form of the metric $g_{M}$, the spaces of parallel one-forms on the phase circles are trivial, which forces the obstruction kernel $\mathcal{K}$ to vanish. Consequently the reduced rank is $r^{\sharp}=1$, and the general theorem yields the sharp inequality
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
This lower bound has immediate physical consequences. It implies that the curvature of the synthetic gauge connection cannot be made block-diagonal with respect to the splitting between momentum and phase directions; at least one independent off-diagonal operator persists, mixing the Brillouin zone with the two phase directions. Crucially, the bound detects the topological coupling of the momentum torus to the two distinct $U(1)$ phases, and therefore remains non-trivial even when the total Chern number of the lowest band is tuned to zero. Thus $r^{\sharp}$ provides a cohomological certificate of locally irremovable Berry curvature, invisible to conventional global invariants. The result suggests interferometric protocols that separately probe the Berry phases associated with $\phi_{+}$ and $\phi_{-}$, offering a way to detect this obstruction even in regimes of vanishing net vorticity.
The paper is organized as follows. Section 2 constructs the extended parameter space $M$, the Kaluza–Klein metric $g_{M}$, and the metric connection $\nabla^{C}$ with torsion $T^{\flat}$. Section 3 analyses the cohomology of $M$, computes the mixed tensor rank $r$, and shows that $\mathcal{K}=0$ in the present geometry. Section 4 extends the Pigazzini–Toda theorem to Kaluza–Klein metrics, ensuring the bound applies to the present geometry. Section 5 applies the general theorem of [12] to obtain the lower bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ and discusses its physical interpretation. Section 6 elaborates the physical implications of the bound, including its persistence when the total Chern number vanishes and its robustness under metric deformations. Section 7 illustrates how the same cohomological data reproduce, in a particular Rashba–Dresselhaus texture, the familiar Chern-number bound on the vortex number, and demonstrates with a tuned example that the holonomy bound remains valid even when the Chern number vanishes. Section 8 outlines prospective interferometric tests for detecting the locally irremovable curvature and indicates directions for future work. Section 9 summarises the results and discusses broader implications for topological quantum matter.
2 Geometric Construction of the Extended Parameter Space
A generic two-component spin–orbit-coupled Bose–Einstein condensate (SOC BEC) confined in a quasi-two-dimensional toroidal trap is characterized by three independent, experimentally accessible degrees of freedom: the crystal momenta $(k_{x},k_{y})$ spanning a Brillouin zone, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin states [9, 11].
We therefore take as the extended parameter space the smooth, compact, oriented product manifold
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}.$
Our goal in this section is to endow $M$ with a Riemannian structure that naturally incorporates the synthetic gauge fields of the SOC BEC, and to define on it a metric connection whose torsion three-form coincides with the physical Berry-curvature three-form.
2.1 Kaluza–Klein metric on the extended space
Let $g_{\mathrm{BZ}}=dk_{x}^{2}+dk_{y}^{2}$ be the flat metric on the Brillouin-zone torus $T^{2}_{\mathrm{BZ}}$. On each phase circle we introduce the standard angular coordinate with period $2\pi$; the canonical $U(1)$-invariant metric on $S^{1}$ is $d\phi^{2}$. The synthetic gauge fields of the SOC BEC provide, for each of the two independent $U(1)$ directions, a Berry connection. Denote by $A^{(+)},A^{(-)}\in\Omega^{1}(T^{2}_{\mathrm{BZ}})$ the corresponding real one-forms (the components of the synthetic $\mathfrak{su}(2)$ connection along the two independent $U(1)$ generators). Their curvatures are the Berry curvature two-forms $F^{(\pm)}=dA^{(\pm)}$.
Following the Kaluza–Klein prescription we define a metric $g_{M}$ on $M$ that couples the base metric to the connection forms:
$g_{M}=\pi^{*}g_{\mathrm{BZ}}+\bigl(d\phi_{+}+\pi^{*}A^{(+)}\bigr)^{2}+\bigl(d\phi_{-}+\pi^{*}A^{(-)}\bigr)^{2},$ (2.1)
where $\pi:M\to T^{2}_{\mathrm{BZ}}$ is the projection onto the momentum coordinates. The one-forms
$\Theta^{(+)}:=d\phi_{+}+\pi^{*}A^{(+)},\qquad\Theta^{(-)}:=d\phi_{-}+\pi^{*}A^{(-)}$
are the horizontal lifts of the phase coordinates; they are globally defined on $M$ and are orthogonal to the base directions with respect to $g_{M}$.
Lemma 2.1 (Properties of the metric $g_{M}$).
The metric (2.1) satisfies the following.
  1. 1.
    $g_{M}$ is smooth, positive definite and makes $\pi:(M,g_{M})\to(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ a Riemannian submersion.
  2. 2.
    The fibres $\pi^{-1}(k)\cong S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ are flat $2$-tori with the induced metric $(d\phi_{+})^{2}+(d\phi_{-})^{2}$.
  3. 3.
    The horizontal distribution $\mathcal{H}=\ker(d\phi_{+}+\pi^{*}A^{(+)})\cap\ker(d\phi_{-}+\pi^{*}A^{(-)})$ is orthogonal to the vertical distribution $\mathcal{V}=\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$.
  4. 4.
    If $A^{(\pm)}=0$ then $g_{M}$ reduces to the trivial product metric $g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$.
Proof.
Smoothness follows from the smoothness of $g_{\mathrm{BZ}}$ and $A^{(\pm)}$. Positive definiteness is clear because each term in (2.1) is non‑negative and their sum vanishes only when all differentials vanish.
For any $k\in T^{2}_{\mathrm{BZ}}$ the restriction of $g_{M}$ to the fibre $\pi^{-1}(k)$ is $(d\phi_{+})^{2}+(d\phi_{-})^{2}$, independent of $k$; hence the fibres are isometric to a flat $2$-torus. The map $\pi$ is a Riemannian submersion because, by construction, $g_{M}$ restricted to $\mathcal{H}$ coincides with $\pi^{*}g_{\mathrm{BZ}}$.
Orthogonality $\mathcal{H}\perp\mathcal{V}$ follows directly from $g_{M}(\partial_{\phi_{\pm}},X)=0$ for every $X\in\mathcal{H}$, which is a consequence of the definition of $\Theta^{(\pm)}$.
The last statement is immediate from (2.1). ∎
Remark 2.2.
Equation (2.1) is physically natural: it measures distances in the extended parameter space by combining the kinetic contribution from momentum changes with the “phase cost” induced by the synthetic gauge potentials. If $A^{(\pm)}=0$, the metric reduces to the trivial product metric $g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$.
2.2 Metric connection with prescribed torsion
On the Riemannian manifold $(M,g_{M})$ we now construct a metric connection $\nabla^{C}$ whose torsion three-form is built from the physical gauge fields. Let $\nabla^{LC}$ be the Levi‑Civita connection of $g_{M}$.
Define the three‑form
$T^{b}:=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}\in\Omega^{3}(M),$ (2.2)
where $F^{(\pm)}\wedge A^{(\pm)}$ are pulled back to $M$ via $\pi$ (we keep the same notation for the pull‑backs). $T^{b}$ is the trace $\operatorname{Tr}(\mathcal{F}\wedge\mathcal{A})$ with respect to the Killing form on $\mathfrak{u}(1)_{+}\oplus\mathfrak{u}(1)_{-}$, where $\mathcal{A}=A^{(+)}\otimes t_{+}+A^{(-)}\otimes t_{-}$ and $\mathcal{F}=F^{(+)}\otimes t_{+}+F^{(-)}\otimes t_{-}$.
Introduce the (2,1)(2,1)‑tensor $K$ by
$K(X,Y,Z):=\frac{1}{2}\,T^{b}(X,Y,Z)\qquad(X,Y,Z\in\mathfrak{X}(M)).$
The connection
$\nabla^{C}_{X}Y:=\nabla^{LC}_{X}Y+K(X,Y,\cdot)^{\sharp}$
is by construction metric ($\nabla^{C}g_{M}=0$) and has totally skew‑symmetric torsion $T^{b}$. Indeed, for any vector fields $X,Y$,
$T^{b}(X,Y,\cdot)^{\sharp}=\nabla^{C}_{X}Y-\nabla^{C}_{Y}X-[X,Y],$
which is precisely the definition of the torsion tensor of $\nabla^{C}$.
Because $T^{b}$ is not parallel with respect to $\nabla^{LC}$ unless the gauge fields are flat, the connection $\nabla^{C}$ satisfies the non‑parallelism condition $\nabla^{LC}T^{b}\neq 0$ required for the applicability of the general theorem [12].
Remark 2.3.
The three‑form $T^{b}$ is not closed in general; its exterior derivative contains terms quadratic in the curvatures $F^{(\pm)}$. Its de Rham cohomology class $T^{b}]\in H^{3}(M;\mathbb{R})$, however, depends only on the harmonic part of $T^{b}$ with respect to $g_{M}$, as will be computed in Section 3.
2.3 Parallel forms in the Kaluza–Klein geometry
A key ingredient of the topological bound is the space of parallel forms on the factors. For the base torus with the flat metric $g_{\mathrm{BZ}}$ we have the well‑known result
$\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})=H^{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}},\qquad\mathrm{vol}_{\mathrm{BZ}}=dk_{x}\wedge dk_{y}.$
The situation for the phase circles is more subtle. On a circle endowed with the standard metric $d\phi^{2}$, the one‑form $d\phi$ is parallel, hence $\mathcal{P}_{1}(S^{1},d\phi^{2})=\mathbb{R}\,d\phi$. In the Kaluza–Klein metric (2.1) the geometry of the fibres is “twisted” by the connection forms $A^{(\pm)}$. The vector fields $\partial_{\phi_{+}},\partial_{\phi_{-}}$ are Killing fields of $g_{M}$, but their dual one‑forms are $\Theta^{(\pm)}$, not $d\phi_{\pm}$.
Lemma 2.4 (Non‑parallelism of $d\phi_{\pm}$).
Assume $A^{(\pm)}$ are not identically zero. Then the one‑forms $d\phi_{+},d\phi_{-}$ are not parallel with respect to $g_{M}$.
Proof.
We compute the covariant derivative $\nabla^{LC}d\phi_{+}$ in local coordinates $(k_{x},k_{y},\phi_{+},\phi_{-})$. The non‑zero Christoffel symbols involving the index $\phi_{+}$ are
$\Gamma^{\phi_{+}}_{k_{x}\phi_{+}}=\frac{1}{2}g^{\phi_{+}\rho}\bigl(\partial_{k_{x}}g_{\phi_{+}\rho}+\partial_{\phi_{+}}g_{k_{x}\rho}-\partial_{\rho}g_{k_{x}\phi_{+}}\bigr).$
Since $g_{k_{x}\phi_{+}}=A^{(+)}_{k_{x}}$ is independent of $\phi_{+}$, we have $\partial_{\phi_{+}}g_{k_{x}\phi_{+}}=0$. Moreover $g_{\phi_{+}\phi_{+}}=1+(A^{(+)}_{k_{x}})^{2}+(A^{(+)}_{k_{y}})^{2}$ and $g_{\phi_{+}\phi_{-}}=A^{(+)}_{k_{x}}A^{(-)}_{k_{x}}+A^{(+)}_{k_{y}}A^{(-)}_{k_{y}}$. Consequently,
[…]
which is non‑zero for generic non‑constant potentials $A^{(\pm)}$. Hence $(\nabla^{LC}_{\partial_{k_{x}}}d\phi_{+})(\partial_{\phi_{+}})=-\Gamma^{\phi_{+}}_{k_{x}\phi_{+}}\neq 0$, showing that $d\phi_{+}$ is not parallel. The same argument applies to $d\phi_{-}$. ∎
Corollary 2.5 (Trivial parallel‑form spaces on the fibres).
For a generic SOC BEC with non‑trivial Berry curvature $F^{(\pm)}\neq 0$,
$\mathcal{P}_{1}\bigl(S^{1}_{\phi_{\pm}},g_{M}|_{\mathrm{fibre}}\bigr)=\{0\}.$
More precisely, any fibre‑tangent 1‑form that is parallel with respect to the full metric $g_{M}$ must be of the form $\xi=c_{+}d\phi_{+}+c_{-}d\phi_{-}$ with constants $c_{+},c_{-}$. However, Lemma 2.4 shows that for non‑constant potentials $A^{(\pm)}$, $\nabla^{LC}d\phi_{\pm}$ has non‑vanishing components along the base directions; consequently no non‑zero linear combination can be parallel on the whole manifold $M$. If $A^{(\pm)}$ are constant but $F^{(\pm)}\neq 0$, they are not closed, hence still not harmonic and therefore not parallel (on a compact Riemannian manifold every parallel form is harmonic).
2.4 Summary of the geometric setup
We have constructed:
  1. 1.
    an extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$;
  2. 2.
    a Kaluza–Klein metric $g_{M}$ on $M$ given by (2.1), coupling the momentum directions to the two independent $U(1)$ phases via the synthetic Berry connections $A^{(\pm)}$;
  3. 3.
    a metric connection $\nabla^{C}$ on $(M,g_{M})$ whose torsion three‑form is the physical quantity $T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}$.
This construction provides the precise Riemannian framework needed to apply the general theorem of [12]. In the following section we analyze the cohomological data that arise from this geometry.
3 Cohomology of the Extended Space and the Mixed Tensor Rank
With the geometric structure $(M,g_{M},\nabla^{C})$ established in Section 2, we now analyze the cohomological data that enter the general lower bound. Recall that the torsion three-form $T^{\flat}\in\Omega^{3}(M)$ is not necessarily closed, but its de Rham cohomology class $T^{\flat}]\in H^{3}(M;\mathbb{R})$ is well defined. Let $\omega_{h}$ be the unique $g_{M}$-harmonic representative of this class; then
$\omega]:=[\omega_{h}]=[T^{\flat}]\in H^{3}(M;\mathbb{R}).$
3.1 Künneth decomposition of $H^{3}(M)$
Since $M$ is a product $T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ and each factor is oriented and compact, the Künneth theorem gives an isomorphism
$H^{3}(M;\mathbb{R})\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
because $H^{3}(T^{2}_{\mathrm{BZ}})=H^{2}(S^{1}_{\phi_{\pm}})=0$ and $H^{1}(T^{2}_{\mathrm{BZ}})$ does not contribute to degree $3$. Consequently $\dim H^{3}(M)=2$, and every class $\eta]\in H^{3}(M)$ can be written uniquely as
$\eta]=[\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}],$ (3.1)
where $\alpha_{\pm}]\in H^{2}(T^{2}_{\mathrm{BZ}})$ and $d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}})$ are the fundamental cohomology classes of the factors.
For the harmonic representative $\omega_{h}$ of our torsion class, the Hodge decomposition with respect to $g_{M}$ yields
$\omega_{h}=\alpha_{+}\wedge d\phi_{+}+\alpha_{-}\wedge d\phi_{-}+d\eta,$
where $\alpha_{\pm}\in\mathcal{H}^{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ are $g_{\mathrm{BZ}}$-harmonic two-forms on the Brillouin torus, and $\eta\in\Omega^{2}(M)$ is a coexact two-form. The term $d\eta$ is exact and therefore orthogonal to harmonic forms in the $L^{2}$-inner product; it does not contribute to the cohomology class $\omega]=[\omega_{h}]$.
Remark 3.1.
The representatives $d\phi_{\pm}$ are harmonic with respect to the standard metric on the circles, but they are not the harmonic representatives of the fundamental classes with respect to the full Kaluza–Klein metric $g_{M}$; the $g_{M}$-harmonic representatives are the horizontal forms $\Theta^{(\pm)}=d\phi_{\pm}+A^{(\pm)}$ introduced in (2.1). Because $d\phi_{\pm}]=[\Theta^{(\pm)}]$ in $H^{1}(S^{1}_{\phi_{\pm}})$, the cohomological decomposition (3.1) is unchanged.
3.2 Mixed tensor rank of the torsion class
The mixed tensor rank of a cohomology class $\eta]\in H^{3}(M)$ written as in (3.1) is defined as the minimal number of simple tensors required to represent it:
$\operatorname{rank}_{\mathbb{R}}([\eta]):=\min\Bigl\{m\;\Big|\;[\eta]=\sum_{i=1}^{m}[\beta_{i}]\otimes[\gamma_{i}],\;[\beta_{i}]\in H^{2}(T^{2}_{\mathrm{BZ}}),\;[\gamma_{i}]\in H^{1}(S^{1}_{\phi_{+}})\oplus H^{1}(S^{1}_{\phi_{-}})\Bigr\}.$
For the class $\omega]$ coming from the physical torsion, the two-forms $\alpha_{+},\alpha_{-}$ in (3.1) are precisely the (cohomology classes of the) Berry curvature two-forms $F^{(+)},F^{(-)}$ of the two independent $U(1)$ sectors. A key observation is that $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$. Therefore, any two non-zero classes in this space are necessarily cohomologically proportional. In a generic two-component SOC BEC, both Berry curvature classes $F^{(+)}]$ and $F^{(-)}]$ are non-zero; consequently, there exists a non-zero constant $\lambda\in\mathbb{R}$ such that
$F^{(-)}]=\lambda[F^{(+)}]\quad\text{in }H^{2}(T^{2}_{\mathrm{BZ}}).$
Substituting this relation into (3.1) gives
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+\lambda[F^{(+)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
which is a single simple tensor. Hence the mixed tensor rank attains the value
$r=1.$
Remark 3.2.
Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the maximal possible mixed tensor rank for a class of the form $\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}]$ with both $\alpha_{\pm}]\neq 0$ is $1$. All experimentally prominent SOC geometries (Rashba–Dresselhaus, Hofstadter, Raman-induced couplings) satisfy the condition $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$, thereby realizing the generic case $r=1$.
Remark 3.3.
Since $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ is a product, its first cohomology decomposes as $H^{1}(M_{2})\cong H^{1}(S^{1}_{+})\oplus H^{1}(S^{1}_{-})$. Therefore an element of $H^{2}(M_{1})\otimes H^{1}(M_{2})$ can be written as a sum of terms of the form $\alpha]\otimes[d\phi_{+}]+[\beta]\otimes[d\phi_{-}]$. The mixed tensor rank $r$ is then the minimal number of simple tensors needed to represent this sum in the space $H^{2}(M_{1})\otimes\bigl(H^{1}(S^{1}_{+})\oplus H^{1}(S^{1}_{-})\bigr)$.
3.3 Vanishing of the obstruction kernel $\mathcal{K}$
The general bound of [12] involves a correction term $\dim\mathcal{K}$, where
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\pi_{2,1}([\omega])\cap\bigl(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})\bigr),$
and $\ker\widetilde{\Psi}_{p}$ is defined analogously with the roles of the two circles exchanged. Here $\pi_{2,1}$ denotes the projection onto the $(2,1)$-Künneth component.
In our geometry, we have $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}}$ (the space of parallel two-forms on the flat torus). The crucial observation, already made in Section 2.3, is that for the Kaluza–Klein metric $g_{M}$ the spaces of parallel one-forms on the phase circles are trivial:
$\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},\,g_{M}|_{\text{\rm fibre}})=\{0\},$
because any $g_{M}$-parallel one-form tangent to a fibre would have to be a constant multiple of $\Theta^{(\pm)}=d\phi_{\pm}+A^{(\pm)}$, and $\Theta^{(\pm)}$ is not closed (hence not of the form $d\phi_{\pm}$) unless $F^{(\pm)}=0$. Since the Berry curvatures $F^{(\pm)}$ are generically non-zero, we indeed have $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})=\{0\}$ in the physical geometry.
Consequently the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(S^{1}_{\phi_{+}})\otimes\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ (the latter appears in $\ker\widetilde{\Psi}_{p}$) are zero. Thus
$\mathcal{K}=0,\qquad\text{and therefore}\qquad\dim\mathcal{K}=0.$ (3.2)
3.4 Summary of the topological data
Based on (3.2) we obtain the reduced rank
$r^{\sharp}\;:=\;r-\dim\mathcal{K}\;=\;1.$
All ingredients required for the application of the general geometric theorem are now in place: the extended parameter space $(M,g_{M})$ carries a metric connection $\nabla^{C}$ with totally skew torsion whose harmonic part represents a mixed cohomology class $\omega]$ of mixed tensor rank $r=1$, and the obstruction kernel $\mathcal{K}$ vanishes identically. In the next section, we apply the theorem to obtain a concrete lower bound on the off-diagonal holonomy of $\nabla^{C}$.
4 Geometric foundations for the PT theorem in Kaluza–Klein geometry
The physical metric (2.1) is not a Riemannian product, whereas Theorem 5.2 of PT [12] is stated for genuine product metrics. The following three results show that the theorem nevertheless applies to our Kaluza–Klein geometry, because the relevant topological and parallel-form data coincide with those of an auxiliary product metric.
4.1 Parallel forms in Kaluza–Klein metrics
Lemma 4.1 (Parallel forms in Kaluza–Klein metrics with flat fibres).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a Kaluza–Klein metric
$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$ (4.1)
where $g_{1}$ is a Riemannian metric on $M_{1}$, $\{\theta^{a}\}$ defines a flat left-invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$ (possibly with $dA^{a}\neq 0$).
Then the restriction map
$\Phi_{1}:\mathcal{P}_{k}(M_{1},g)\longrightarrow\mathcal{P}_{k}(M_{1},g_{1}),\qquad\eta\mapsto\eta|_{D_{1}},$
is a well-defined linear isomorphism.
Analogously, define $\mathcal{P}_{k}(M_{2},g)$ as the space of forms $\xi\in\Omega^{k}(M)$ that are invariant along $M_{1}$ (their coefficients depend only on the $M_{2}$-coordinates) and satisfy $\nabla^{LC}\xi=0$. Then
$\Phi_{2}:\mathcal{P}_{k}(M_{2},g)\longrightarrow\mathcal{P}_{k}(M_{2},h),\qquad\xi\mapsto\xi|_{D_{2}},$
is an isomorphism.
Proof.
Choose local coordinates $(x^{i},y^{\alpha})$ adapted to the orthogonal splitting $TM=D_{1}\oplus D_{2}$ induced by the product structure. In these coordinates the metric reads
$g=g_{ij}(x)\,dx^{i}dx^{j}+h_{\alpha\beta}(y)\bigl(dy^{\alpha}+A^{\alpha}_{i}(x)dx^{i}\bigr)\bigl(dy^{\beta}+A^{\beta}_{j}(x)dx^{j}\bigr).$ (4.2)
Let $\eta\in\mathcal{P}_{k}(M_{1},g)$. Because $\eta$ is parallel and tangent to $D_{1}$, in the coordinates above it can be written as
$\eta=\frac{1}{k!}\,\eta_{i_{1}\dots i_{k}}(x)\,dx^{i_{1}}\wedge\dots\wedge dx^{i_{k}}.$
We compute the Christoffel symbols of $g$. From (4.2) one obtains
$\Gamma^{i}_{jk}$ $=\Gamma^{(g_{1})i}_{jk},$
$\Gamma^{i}_{j\beta}$ $=-\frac{1}{2}\,g^{i\ell}\bigl(\partial_{j}g_{\ell\beta}+\partial_{\beta}g_{j\ell}-\partial_{\ell}g_{j\beta}\bigr),$
$\Gamma^{\alpha}_{jk}$ $=\frac{1}{2}\,h^{\alpha\gamma}\bigl(\partial_{j}g_{k\gamma}+\partial_{k}g_{j\gamma}-\partial_{\gamma}g_{jk}\bigr),$
$\Gamma^{\alpha}_{\beta\gamma}$ $=\Gamma^{(h)\alpha}_{\beta\gamma},$
where $\Gamma^{(g_{1})}$ and $\Gamma^{(h)}$ denote the Christoffel symbols of $g_{1}$ and $h$ respectively.
For any vector field $X=X^{i}\partial_{i}+X^{\alpha}\partial_{\alpha}$,
$(\nabla^{LC}_{X}\eta)_{i_{1}\dots i_{k}}=X^{j}\partial_{j}\eta_{i_{1}\dots i_{k}}-\sum_{r=1}^{k}\Gamma^{\ell}_{ji_{r}}X^{j}\eta_{i_{1}\dots i_{r-1}\ell i_{r+1}\dots i_{k}}-\Gamma^{\alpha}_{j\alpha}X^{j}\eta_{i_{1}\dots i_{k}}.$
The condition $\nabla^{LC}\eta=0$ therefore splits into two independent sets of equations. For $X=X^{i}\partial_{i}$ (tangent to $M_{1}$) we obtain
$\nabla^{(g_{1})}_{X}\eta=0,$
where $\nabla^{(g_{1})}$ is the Levi‑Civita connection of $g_{1}$. Hence $\eta|_{D_{1}}$ is $g_{1}$-parallel.
For $X=X^{\alpha}\partial_{\alpha}$ (tangent to $M_{2}$) we get
$\Gamma^{\alpha}_{j\alpha}(x)X^{j}\eta_{i_{1}\dots i_{k}}=0\qquad\text{for all }X^{j}.$
Why $\Gamma^{\alpha}_{j\alpha}=0$. Because the fibres are flat, the fibre metric $h_{\alpha\beta}$ is constant in the chosen coordinates; therefore $\partial_{j}h_{\alpha\beta}=0$. Moreover, from (4.2) we have $g_{j\alpha}=h_{\alpha\beta}A^{\beta}_{j}(x)$, which is independent of $y$. Consequently
$\partial_{\gamma}g_{j\alpha}=0,\qquad\partial_{\alpha}g_{j\gamma}=0.$
Substituting these facts into the expression for $\Gamma^{\alpha}_{j\alpha}$ gives
$\Gamma^{\alpha}_{j\alpha}=\frac{1}{2}h^{\alpha\gamma}\bigl(\partial_{j}h_{\alpha\gamma}+\partial_{\alpha}g_{j\gamma}-\partial_{\gamma}g_{j\alpha}\bigr)=0.$
Thus the condition coming from $X^{\alpha}\partial_{\alpha}$ is automatically satisfied.
Hence $\eta\in\mathcal{P}_{k}(M_{1},g)$ if and only if $\eta|_{D_{1}}\in\mathcal{P}_{k}(M_{1},g_{1})$. Consequently $\Phi_{1}$ is injective. Surjectivity follows because any $g_{1}$-parallel form $\widetilde{\eta}$ on $M_{1}$ can be lifted to a $g$-parallel form on $M$ by declaring it independent of the fibre coordinates; the computation above shows that the lifted form is indeed parallel with respect to $g$.
Therefore $\Phi_{1}$ is an isomorphism. The proof for $\Phi_{2}$ proceeds identically after exchanging the roles of base and fibre. ∎
Lemma 4.2 (Harmonic Künneth decomposition for flat‑fibre KK metrics).
Let $g$ be a Kaluza–Klein metric as in (4.1) with flat fibres. Then the natural wedge‑product map
$\Phi:\mathcal{H}^{p}(M_{1},g_{1})\otimes\mathcal{H}^{q}(M_{2},h)\longrightarrow\mathcal{H}^{p+q}(M,g)$
is an isomorphism onto the subspace of harmonic forms of pure bidegree $(p,q)$. In particular, every $g$-harmonic form $\eta\in\mathcal{H}^{p+q}(M,g)$ that belongs to the $(p,q)$-Künneth component admits a unique decomposition
$\eta=\sum_{i=1}^{r}\alpha_{i}\wedge\beta_{i},\qquad\alpha_{i}\in\mathcal{H}^{p}(M_{1},g_{1}),\;\beta_{i}\in\mathcal{H}^{q}(M_{2},h),$
with $r=\operatorname{rank}_{\mathbb{R}}([\eta]_{\mathrm{mixed}})$.
Proof.
Because the fibres are flat, the Hodge Laplacian of $g$ decouples as $\Delta_{g}=\Delta_{g_{1}}\otimes\mathrm{Id}+\mathrm{Id}\otimes\Delta_{h}$ on forms of pure bidegree. Hence a form of bidegree $(p,q)$ is $g$-harmonic iff its components are harmonic with respect to $g_{1}$ and $h$ respectively. The statement then follows from the classical Künneth theorem for de Rham cohomology combined with Hodge theory. ∎
4.2 Persistence of non‑parallelism under metric change
For the application of the PT theorem to an auxiliary product metric, we need to ensure that the non‑parallelism condition $\nabla^{LC}T\neq 0$ is preserved when we replace the original Kaluza–Klein metric by a nearby product metric.
Lemma 4.3 (Openness of the non‑parallelism condition).
Let $M$ be a smooth manifold and $T\in\Omega^{3}(M)$ a fixed 3‑form. Denote by $\mathcal{M}$ the space of Riemannian metrics on $M$ endowed with the $C^{1}$-topology. The set
$\mathcal{M}_{\neq 0}(T):=\bigl\{g\in\mathcal{M}\mid\nabla^{LC}(g)T\neq 0\bigr\}$
is open in $\mathcal{M}$. In particular, if $\nabla^{LC}(g)T\neq 0$ for some metric $g$, then there exists a $C^{1}$-neighbourhood of $g$ consisting of metrics for which the covariant derivative of $T$ is also non‑zero.
Proof.
The Levi‑Civita connection $\nabla^{LC}(g)$ depends smoothly (in fact, rationally) on the metric $g$ and its first derivatives. Consequently, the tensor field $\nabla^{LC}(g)T$ depends continuously on $g$ in the $C^{1}$-topology. The condition $\nabla^{LC}(g)T\neq 0$ means that at least one component of this tensor is non‑zero at some point; by continuity, the same component remains non‑zero for all metrics sufficiently close to $g$ in the $C^{1}$-sense. ∎
Remark 4.4.
In our setting, the original metric $g$ is the Kaluza–Klein metric (4.1) and the auxiliary product metric $g_{\mathrm{prod}}=g_{1}\oplus h$ can be chosen arbitrarily close to $g$ in the $C^{\infty}$ (hence also $C^{1}$) topology by taking the gauge potentials $A^{a}$ small, or by a direct approximation argument. Lemma 4.3 guarantees that if $\nabla^{LC}(g)T\neq 0$, then also $\nabla^{LC}(g_{\mathrm{prod}})T\neq 0$ for a generic choice of $g_{\mathrm{prod}}$ near $g$.
4.3 Harmonic representatives in Kaluza–Klein geometry
Lemma 4.5 (Invariance of the mixed tensor rank under Kaluza–Klein metrics).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold, and let $g$ be any Riemannian metric on $M$ (in particular, a Kaluza–Klein metric of the form (4.1) with flat fibres). For a cohomology class
$\omega]\in\bigl(H^{2}(M_{1})\otimes H^{1}(M_{2})\bigr)\oplus\bigl(H^{1}(M_{1})\otimes H^{2}(M_{2})\bigr)\subset H^{3}(M),$
its mixed tensor rank
$r:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$
depends only on the topological class $\omega]$; in particular, it is the same whether computed with respect to the metric $g$ or with respect to the genuine product metric $g_{\mathrm{prod}}=g_{1}\oplus h$.
Proof.
The mixed tensor rank is defined algebraically as the minimal number of simple tensors needed to represent the element $\omega]$ in the finite‑dimensional vector space
$\bigl(H^{2}(M_{1})\otimes H^{1}(M_{2})\bigr)\oplus\bigl(H^{1}(M_{1})\otimes H^{2}(M_{2})\bigr).$
This rank is independent of any auxiliary Riemannian metric, because the cohomology spaces $H^{p}(M_{i})$ themselves are topological invariants. Hence, for any two metrics $g$ and $g^{\prime}$ on $M$, the mixed tensor rank of $\omega]$ coincides. ∎
4.4 The PT bound for Kaluza–Klein metrics with flat fibres
With the previous preparatory lemmas, we can now state and prove the extension of the Pigazzini-Toda theorem to Kaluza-Klein geometries.
Theorem 4.6 (PT bound for Kaluza–Klein metrics with flat fibres).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold equipped with a Kaluza–Klein metric
$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$
where $\{\theta^{a}\}$ defines a flat left‑invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$. Let $\nabla^{C}=\nabla^{LC}(g)+\frac{1}{2}T$ be a metric connection on $(M,g)$ with totally skew‑symmetric torsion $T\in\Omega^{3}(M)$. Denote by $\omega_{h}$ the unique $g$-harmonic representative of $T]\in H^{3}(M;\mathbb{R})$ and assume
  1. 1.
    $\nabla^{LC}(g)T\neq 0$;
  2. 2.
    $\omega_{h}]$ has a non‑trivial mixed component in the Künneth decomposition;
  3. 3.
    the mixed tensor rank $r:=rank_{\mathbb{R}}\bigl([\omega_{h}]_{\mathrm{mixed}}\bigr)$ is finite.
Define the obstruction space
$cK\;:=\;\bigl(\pi_{2,1}([\omega_{h}])\cap(cP_{2}(M_{1},g)\otimes cP_{1}(M_{2},g))\bigr)\oplus\bigl(\pi_{1,2}([\omega_{h}])\cap(cP_{1}(M_{1},g)\otimes cP_{2}(M_{2},g))\bigr),$
and set $r^{\sharp}:=r-\dim cK$.
Then for every point $p\in M$,
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r^{\sharp}.$ (4.3)
Moreover, $r^{\sharp}$ depends only on the cohomology class $\omega_{h}]_{\mathrm{mixed}}$ and on the spaces $cP_{k}(M_{i},g)$; consequently, it is invariant under smooth deformations of $g$ that preserve those spaces.
Proof.
We proceed in several carefully justified stages.
Consider the genuine Riemannian product metric $g_{\mathrm{prod}}:=g_{1}\oplus h$ on $M=M_{1}\times M_{2}$. Define a new metric connection on $(M,g_{\mathrm{prod}})$ by $\widetilde{\nabla}^{C}:=\nabla^{LC}(g_{\mathrm{prod}})+\tfrac{1}{2}T$, where $T$ is the same 3‑form as in the statement. Since $\nabla^{LC}(g_{\mathrm{prod}})$ is torsion‑free and $T$ is totally skew, $\widetilde{\nabla}^{C}$ is metric with respect to $g_{\mathrm{prod}}$ and has torsion $T$.
Because the fibres of the Kaluza–Klein metric $g$ are flat, Lemma 4.1 provides canonical isomorphisms
$\Phi_{i}^{(k)}:cP_{k}(M_{i},g)\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}cP_{k}(M_{i},g_{\mathrm{prod}}),\qquad i=1,2,\;k=1,2,$
obtained by restricting a parallel form to the corresponding factor. Consequently, the obstruction spaces computed with $g$ and with $g_{\mathrm{prod}}$ have identical dimensions: $\dim cK=\dim\widetilde{cK}$, where $\widetilde{cK}$ is defined analogously to $cK$ but using the spaces $cP_{k}(M_{i},g_{\mathrm{prod}})$.
The mixed tensor rank $r=rank_{\mathbb{R}}\bigl([\omega_{h}]_{\mathrm{mixed}}\bigr)$ is a purely topological invariant of the cohomology class $\omega_{h}]$ (Lemma 4.5); therefore it is unchanged when we replace $g$ by $g_{\mathrm{prod}}$.
Let $\widetilde{\omega}_{h}$ be the unique $g_{\mathrm{prod}}$‑harmonic representative of the same class $\omega_{h}]$. By Hodge theory, $\widetilde{\omega}_{h}]=[\omega_{h}]$ and consequently $rank_{\mathbb{R}}\bigl([\widetilde{\omega}_{h}]_{\mathrm{mixed}}\bigr)=r$.
The triple $(M,g_{\mathrm{prod}},\widetilde{\nabla}^{C})$ satisfies all hypotheses of Theorem 5.2 in [12]. Indeed, $g_{\mathrm{prod}}$ is a genuine product metric; $\widetilde{\nabla}^{C}$ is a metric connection on $(M,g_{\mathrm{prod}})$ with totally skew torsion $T$; the condition $\nabla^{LC}(g)T\neq 0$ implies $\nabla^{LC}(g_{\mathrm{prod}})T\neq 0$ in the generic situation we consider—the difference $\nabla^{LC}(g)-\nabla^{LC}(g_{\mathrm{prod}})$ is a tensor field, and if $\nabla^{LC}(g)T$ were to vanish while $\nabla^{LC}(g_{\mathrm{prod}})T$ remained non‑zero, this would impose a special algebraic relation between the metrics and $T$ that we exclude as nongeneric; finally, the harmonic part $\widetilde{\omega}_{h}$ of $T$ with respect to $g_{\mathrm{prod}}$ represents a mixed cohomology class of tensor rank $r$.
Applying Theorem 5.2 we obtain
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\widetilde{\nabla}^{C})\geq r-\dim\widetilde{cK}=r^{\sharp}.$ (4.4)
Write the two connections as $\nabla^{C}=\nabla^{LC}(g)+K$ and $\widetilde{\nabla}^{C}=\nabla^{LC}(g_{\mathrm{prod}})+K$ with $K_{X}Y=\tfrac{1}{2}T(X,Y,\cdot)^{\sharp}$. Their curvature tensors are given by $R^{\nabla}=R^{\nabla^{LC}}+\nabla^{LC}K+K\circ K$. The off‑diagonal part with respect to the splitting $TM=D_{1}\oplus D_{2}$ arises from terms where the arguments $X,Y$ belong to different factors.
Because the Levi‑Civita curvature of a product metric is block‑diagonal, $R^{\nabla^{LC}(g_{\mathrm{prod}})}_{\mathrm{off}}=0$. For a general metric $g$, the difference $\Delta:=\nabla^{LC}(g)-\nabla^{LC}(g_{\mathrm{prod}})$ is a tensor field whose Christoffel symbols in adapted coordinates satisfy $\Delta^{\alpha}_{ij}=0$ and $\Delta^{i}_{j\alpha}=0$; the only non‑zero components mix one base and one fibre index. Consequently, when we compute the off‑diagonal component of $\nabla^{LC}(g)K$, the contribution of $\Delta$ takes the form $(\Delta_{X}K)_{Y}Z-(\Delta_{Y}K)_{X}Z$ with $X$ and $Y$ tangent to different factors. A direct inspection in coordinates shows that such terms are linear combinations of the components of $K$ itself, evaluated at the point $p$. Hence they do not increase the dimension of the space spanned by the off‑diagonal curvature operators beyond the dimension already produced by the $\nabla^{LC}(g_{\mathrm{prod}})K$ term together with the quadratic term $K\circ K$.
More formally, consider the linear map $\Psi_{p}:cV_{2,1}\to so(T_{p}M)$ defined in the proof of Theorem 5.2 of [12], which associates to each simple mixed tensor $\alpha\otimes\beta$ the off‑diagonal curvature operator generated by $\nabla^{LC}K$. Replacing $\nabla^{LC}(g_{\mathrm{prod}})$ by $\nabla^{LC}(g)$ changes $\Psi_{p}$ by an additive linear map whose image lies in the span of the operators obtained by evaluating $K$ itself on vectors from different factors. Because the quadratic term $K\circ K$ already accounts for all such algebraic combinations, the rank of the total off‑diagonal curvature cannot decrease when we replace $g_{\mathrm{prod}}$ by $g$; it may only increase. Therefore, for every $p\in M$,
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq\dim\mathrm{hol}^{\mathrm{off}}_{p}(\widetilde{\nabla}^{C}).$ (4.5)
Combining (4.4) with (4.5) yields precisely
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp},$
which is the desired estimate (4.3).
The integer $r^{\sharp}=r-\dim cK$ depends solely on the topological cohomology class $\omega_{h}]_{\mathrm{mixed}}$ (which determines $r$) and on the spaces $cP_{k}(M_{i},g)$ (which determine $\dim cK$). Any smooth deformation of $g$ that leaves these parallel‑form spaces unchanged therefore leaves $r^{\sharp}$ invariant. This completes the proof. ∎
4.5 Application to the SOC BEC model
Proposition 4.7 (Application to the SOC BEC model).
In the spin‑orbit‑coupled BEC model of Section 2, the extended parameter space is $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The metric (2.1) is precisely of the Kaluza–Klein type (4.1) with $M_{1}=T^{2}_{\mathrm{BZ}}$, $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, left‑invariant forms $\theta^{1}=d\phi_{+},\ \theta^{2}=d\phi_{-}$, and gauge potentials $A^{1}=A^{(+)},\ A^{2}=A^{(-)}$. The fibres are flat because $d\theta^{1}=d\theta^{2}=0$.
For this geometry:
  1. 1.
    $\mathcal{P}_{1}(M_{2},g)=\{0\}$ (Corollary 2.5).
  2. 2.
    $\mathcal{P}_{2}(M_{1},g)\cong\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})\cong\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}}$.
  3. 3.
    Consequently the obstruction kernel vanishes: $\mathcal{K}=0$.
If the Berry curvatures satisfy $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}})$, then the mixed tensor rank attains its maximal possible value for this geometry, $r=1$, and Theorem 4.6 gives
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
Proof.
The identification with the Kaluza–Klein framework is immediate from the definitions. Item 1 follows from Corollary 2.5, which shows that no non‑zero linear combination of $d\phi_{+}$ and $d\phi_{-}$ is parallel when the potentials $A^{(\pm)}$ are not constant. Item 2 is a direct consequence of Lemma 4.1 applied to the base factor. Since $\mathcal{P}_{1}(M_{2},g)=0$, the tensor products appearing in the definition of $\mathcal{K}$ are zero, hence $\mathcal{K}=0$.
For the last statement, recall that $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$. If both $F^{(+)}]$ and $F^{(-)}]$ are non‑zero, they are necessarily proportional in this one‑dimensional space; therefore the mixed class
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$
has tensor rank $r=1$ (it can be written as a single simple tensor $F^{(+)}]\otimes([d\phi_{+}]+\lambda[d\phi_{-}])$ for some $\lambda\in\mathbb{R}$). Substituting $r=1$ and $\dim\mathcal{K}=0$ into (4.3) yields the claimed lower bound. ∎
5 Topological Lower Bound and Its Physical Consequences
With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now apply the general theorem of [12] to the physical connection $\nabla^{C}$.
5.1 Statement of the bound
The PT theorem concerns a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on a product Riemannian manifold $M=M_{1}\times M_{2}$. If the harmonic part of $T$ represents a mixed cohomology class $\omega]\in H^{p}(M_{1})\otimes H^{q}(M_{2})$ and satisfies $\nabla^{LC}T\neq 0$, then the dimension of its off-diagonal holonomy subspace is bounded below by
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;:=\;r-\dim\mathcal{K},$ (5.1)
where $r=\operatorname{rank}_{\mathbb{R}}([\omega])$ is the mixed tensor rank of the class, and
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\pi_{2,1}([\omega])\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),$
with $\ker\widetilde{\Psi}_{p}$ defined analogously for the complementary bidegree.
5.2 Main theorem: topological lower bound for SOC BECs
Theorem 5.1 (Topological lower bound for off-diagonal holonomy in two-component SOC BECs).
Let $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be the extended parameter space equipped with the Kaluza–Klein metric
$g_{M}=\pi^{*}g_{\mathrm{BZ}}+\bigl(d\phi_{+}+\pi^{*}A^{(+)}\bigr)^{2}+\bigl(d\phi_{-}+\pi^{*}A^{(-)}\bigr)^{2},$
where $g_{\mathrm{BZ}}=dk_{x}^{2}+dk_{y}^{2}$ is the flat metric on the Brillouin torus, $A^{(\pm)}$ are the Berry connection one-forms on $T^{2}_{\mathrm{BZ}}$, and $\pi:M\to T^{2}_{\mathrm{BZ}}$ is the projection. Let $\nabla^{C}$ be the metric connection on $(M,g_{M})$ defined by
$\nabla^{C}_{X}Y=\nabla^{LC}_{X}Y+\frac{1}{2}\,T^{b}(X,Y,\cdot)^{\sharp},\qquad T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)},$
where $\nabla^{LC}$ is the Levi-Civita connection of $g_{M}$ and $F^{(\pm)}=dA^{(\pm)}$ are the Berry curvature two-forms.
Assume that the cohomology classes $F^{(+)}],[F^{(-)}]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{R})$ are both non-zero. Then:
  1. 1.
    The harmonic part $\omega_{h}$ of $T^{b}$ represents a mixed cohomology class
    $\omega]=[\omega_{h}]\in H^{3}(M;\mathbb{R})\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
    whose mixed tensor rank is $r:=\operatorname{rank}_{\mathbb{R}}([\omega])=1$ (the maximum possible value for this geometry).
  2. 2.
    The obstruction kernel $\mathcal{K}$ associated with the product splitting $M=T^{2}_{\mathrm{BZ}}\times(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ vanishes:
    $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}=0,$
    where $\Psi_{p},\widetilde{\Psi}_{p}$ are the linear maps defined in [12].
  3. 3.
    Consequently, the off-diagonal holonomy algebra of $\nabla^{C}$ satisfies the sharp lower bound
    $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;:=\;r-\dim\mathcal{K}\;=\;1.$
    In particular, there exists at least one linearly independent curvature operator $R^{C}(X,Y)$ that maps $T_{k}T^{2}_{\mathrm{BZ}}$ into $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$ and vice-versa.
Proof.
We prove the three statements in order.
1. Computation of the mixed tensor rank. By the Künneth theorem, every class in $H^{3}(M;\mathbb{R})$ decomposes uniquely as
$\eta]=[\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}],\qquad[\alpha_{\pm}]\in H^{2}(T^{2}_{\mathrm{BZ}}),\;[d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}}).$
The harmonic representative $\omega_{h}$ of $T^{b}]$ is cohomologous to $T^{b}$. A direct computation using the Hodge decomposition with respect to $g_{M}$ shows that
$\alpha_{+}]=[F^{(+)}],\qquad[\alpha_{-}]=[F^{(-)}].$
Since $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, any two non-zero elements in this space are linearly dependent. Therefore, if both $F^{(+)}]$ and $F^{(-)}]$ are non-zero, there exists $\lambda\in\mathbb{R}$ such that $F^{(-)}]=\lambda[F^{(+)}]$. Consequently,
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+\lambda[F^{(+)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
which is a single simple tensor. Hence the mixed tensor rank is $r=1$.
2. Vanishing of the obstruction kernel. The kernel $\mathcal{K}$ is defined as
$\mathcal{K}=\bigl(\pi_{2,1}([\omega])\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}))\bigr)\oplus\bigl(\pi_{1,2}([\omega])\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{-}}))\bigr),$
where $\pi_{2,1},\pi_{1,2}$ are the projections onto the respective Künneth components, and $\mathcal{P}_{k}(N)$ denotes the space of $g_{M}$-parallel $k$-forms on the factor $N$.
For the flat torus $T^{2}_{\mathrm{BZ}}$ we have
$\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=H^{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}},\qquad\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})=H^{1}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\langle dk_{x},dk_{y}\rangle.$
For the phase circles, Corollary 2.5 shows that
$\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},\,g_{M}|_{\mathrm{fibre}})=\{0\}$
when the Berry curvatures are non-flat. Because $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})=\{0\}$, the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{-}})$ are both zero. Hence $\mathcal{K}=0$ and $\dim\mathcal{K}=0$.
3. Application of the PT theorem. Theorem 4.6 (the extension of the PT theorem to Kaluza–Klein metrics with flat fibres) applies because:
  • $g_{M}$ is a Kaluza–Klein metric with flat fibres (the fibres are flat 2-tori);
  • $\nabla^{LC}T^{b}\neq 0$ unless $F^{(\pm)}$ are flat (which we exclude);
  • $\omega]$ has non-trivial mixed component;
  • The mixed tensor rank $r=1$ is finite.
Substituting $r=1$ and $\dim\mathcal{K}=0$ into the bound (5.1) yields
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
The geometric meaning of $\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})$ is precisely the algebra generated by those curvature operators $R^{C}(X,Y)$ that interchange the tangent spaces of the factors $T^{2}_{\mathrm{BZ}}$ and $S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The bound therefore guarantees the existence of at least one independent such operator. ∎
5.3 Physical interpretation: non-removable off-diagonal curvature
Inequality (5.1) has a direct geometric meaning. The holonomy algebra $\mathfrak{hol}(\nabla^{C})$ is generated by the curvature operators $R^{C}(X,Y)$. The “off-diagonal” part $\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})$ consists of those generators that map the tangent space of the Brillouin zone $T_{k}T^{2}_{\mathrm{BZ}}$ into the phase directions $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$ and vice-versa. The bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ therefore guarantees that there exists at least one linearly independent curvature operator that mixes momentum with the phase directions.
In physical terms, this means that the synthetic gauge curvature of the SOC BEC cannot be made completely block-diagonal by any smooth gauge transformation or deformation of the condensate wavefunction. One cannot “flatten” the Berry curvature along both phase directions simultaneously for all points in the Brillouin zone. The topology of the mixed class $\omega]$ forces a minimal geometric entanglement between the momentum and phase degrees of freedom.
5.4 Persistence of the bound when the total Chern number vanishes
A crucial feature of the bound is that it is insensitive to the value of the total first Chern number
$c_{1}^{\,\mathrm{tot}}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}\bigl(F^{(+)}+F^{(-)}\bigr).$
Corollary 5.2 (Bound independent of the total Chern number).
Let $c_{1}^{(+)}=\frac{1}{2\pi}\int_{T^{2}}F^{(+)}$ and $c_{1}^{(-)}=\frac{1}{2\pi}\int_{T^{2}}F^{(-)}$. If $c_{1}^{(+)}\neq 0$ and $c_{1}^{(-)}=-c_{1}^{(+)}$, then $c_{1}^{\,\mathrm{tot}}=0$. Nevertheless, provided $F^{(+)}]\neq 0$ (hence also $F^{(-)}]\neq 0$), the mixed tensor rank remains $r=1$ and Theorem 5.1 still gives
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
Thus the topological obstruction persists even when the net Berry flux vanishes.
Proof.
If $c_{1}^{(+)}\neq 0$ then $F^{(+)}]\neq 0$ in $H^{2}(T^{2})$. The condition $c_{1}^{(-)}=-c_{1}^{(+)}$ implies $F^{(-)}]=-[F^{(+)}]$ in cohomology (because integration gives an isomorphism $H^{2}(T^{2})\cong\mathbb{R}$). Hence both classes are non-zero and the mixed tensor rank attains its maximum $r=1$. The vanishing of $\mathcal{K}$ does not depend on the values of the Chern numbers but only on the non-existence of parallel one-forms on the fibres, which holds as long as $F^{(\pm)}\neq 0$. ∎
This demonstrates that $r^{\sharp}$ detects a local topological constraint that survives when the familiar global invariant vanishes. While the Chern number quantizes the net flux, it can be zero due to cancellation of local contributions; the mixed-rank bound, in contrast, certifies that the curvature cannot be simultaneously flattened along the two independent phase directions.
5.5 Invariance under metric deformations
The lower bound (5.1) depends only on the topological data $r$ and on the spaces of parallel forms $\mathcal{P}_{k}(M_{i})$. In our construction, these spaces are determined by the Kaluza–Klein metric $g_{M}$, which in turn is built from the physical Berry connections $A^{(\pm)}$. Any smooth deformation of the system that does not change the cohomology classes of $A^{(\pm)}$ and that preserves the vanishing of $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})$ (i.e., that does not make the Berry curvatures exactly flat) leaves the bound unchanged. This stability makes the result robust against typical experimental perturbations.
The inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ is the main result of this paper. It provides a cohomologically defined, locally non-trivial obstruction to the complete removal of Berry curvature in a two-component SOC BEC, valid even when the total Chern number is zero. In the following section, we illustrate how the same cohomological framework reproduces, in a particular Rashba–Dresselhaus texture, the familiar Chern-number bound on vortex count, while emphasizing that the holonomy bound operates beyond the Chern-number paradigm.
6 Physical Consequences and Experimental Signatures
The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$, derived in Section 5, imposes concrete physical constraints on the synthetic gauge structure of a two-component SOC BEC.
6.1 Geometric obstruction to flatness
The bound guarantees the existence of at least one independent off-diagonal curvature operator that mixes the momentum directions $T_{k}T^{2}_{\mathrm{BZ}}$ with the phase directions $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$. In terms of the synthetic gauge potential $\mathbf{A}_{\mathrm{syn}}$, this means that no smooth gauge transformation or continuous deformation of the condensate order parameter can simultaneously make all the Berry curvature components
$\frac{\partial A_{\phi_{+}}}{\partial k_{j}}-\frac{\partial A_{k_{j}}}{\partial\phi_{+}}\qquad\text{and}\qquad\frac{\partial A_{\phi_{-}}}{\partial k_{j}}-\frac{\partial A_{k_{j}}}{\partial\phi_{-}}$
vanish identically for all $j=x,y$ and for all points $(k,\phi_{+},\phi_{-})$ in the extended parameter space. The topology of the mixed cohomology class $\omega]$ therefore enforces a minimal geometric entanglement between the momentum and the two phase degrees of freedom: at least one independent coupling between them remains irreducible under any smooth gauge change.
6.2 Independence from the global Chern number
A key advantage of the bound is its insensitivity to the value of the total first Chern number
$c_{1}^{\,\mathrm{tot}}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}\bigl(F^{(+)}+F^{(-)}\bigr).$
One can easily construct configurations (e.g., by tuning the two Berry curvatures to have opposite but non-zero fluxes) where $c_{1}^{\,\mathrm{tot}}=0$ while the mixed tensor rank remains $r=1$. In such a situation, the usual global topological invariant would indicate no obstruction, yet the inequality $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ still holds, preventing the complete local flattening of the curvature. Thus $r^{\sharp}$ acts as a local topological invariant that can be non-trivial even when all traditional global invariants vanish.
7 Illustrative Examples: From Chern Numbers to the Mixed Rank
The general lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ derived in Section 5 applies to any generic two-component SOC BEC. In this section, we illustrate its working in two concrete situations: first in the well-studied Rashba–Dresselhaus texture, where it reproduces the familiar Chern-number bound on vortex count; second, in a deliberately tuned configuration where the total Chern number vanishes while the mixed-rank obstruction remains, highlighting the new phenomenon captured by $r^{\sharp}$.
7.1 Example 1: Rashba–Dresselhaus texture and the maximal rank $r=1$
Consider a two-dimensional two-level Hamiltonian of Rashba–Dresselhaus type
$\hat{H}_{\mathrm{RD}}(k)\;=\;\frac{1}{2m}\bigl(\mathbf{p}-\mathbf{A}_{\mathrm{syn}}(k)\bigr)^{2}\;+\;\mathbf{B}_{\mathrm{syn}}(k)\cdot\boldsymbol{\sigma},\qquad k=(k_{x},k_{y})\in T^{2}_{\mathrm{BZ}},$
with synthetic vector potential and synthetic Zeeman field of the schematic form
$\mathbf{A}_{\mathrm{syn}}(k)=\kappa(\sigma_{y},-\sigma_{x}),\qquad\mathbf{B}_{\mathrm{syn}}(k)=\Omega\,\mathbf{e}_{z},$
where $\kappa,\Omega$ are suitable coupling parameters. For generic parameters, the lower dressed band defines a spin texture
$\mathbf{n}:T^{2}_{\mathrm{BZ}}\longrightarrow S^{2},\qquad\|\mathbf{n}(k)\|=1,$
obtained by normalizing the Bloch vector of the occupied band.
The corresponding eigenline bundle $\mathbb{L}\to T^{2}_{\mathrm{BZ}}$ is the pullback of the tautological line bundle over $\mathbb{C}P^{1}\simeq S^{2}$ via $\mathbf{n}$. If $\deg(\mathbf{n})=d$, then
$c_{1}(\mathbb{L})\;=\;d\cdot c_{1}\bigl(\mathcal{O}(-1)\bigr),$
and hence $|c_{1}(\mathbb{L})|=|d|$. In particular, for the simplest Rashba–Dresselhaus texture in which $\mathbf{n}$ has degree $d=1$, one has a line bundle with non-vanishing first Chern number
$c_{1}(\mathbb{L})=\pm 1.$
(Up to sign, which is fixed by the chirality convention, only the fact that $c_{1}(\mathbb{L})\neq 0$ will matter in what follows.)
Let $F$ be the Berry curvature two-form of the lower band, representing the cohomology class $F]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{Z})$. As usual,
$c_{1}(\mathbb{L})\;=\;\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F\;=\;\pm 1.$
In the extended parameter space
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}},$
equipped with the Kaluza–Klein metric $g_{M}$ introduced in Section 2, the synthetic torsion three-form
$T^{\flat}=\operatorname{Tr}(\mathcal{F}\wedge\mathcal{A})$
has a $g_{M}$-harmonic representative $\omega_{h}$ whose de Rham class can be written, after a convenient gauge choice, as
$\omega]=[\omega_{h}]\;=\;[F]\otimes[d\phi_{+}]\;+\;[F]\otimes[d\phi_{-}]\;\in\;H^{2}(T^{2}_{\mathrm{BZ}})\otimes\bigl(H^{1}(S^{1}_{\phi_{+}})\oplus H^{1}(S^{1}_{\phi_{-}})\bigr).$
Here $F]\neq 0$ encodes the non-trivial first Chern class of the band, and $d\phi_{+}],[d\phi_{-}]$ are the fundamental classes of the phase circles.
Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the two terms $F]\otimes[d\phi_{+}]$ and $F]\otimes[d\phi_{-}]$ are not linearly independent; they share the same factor $[F]$. Consequently, the mixed class can be rewritten as a single simple tensor:
$\omega]=[F]\otimes\bigl([d\phi_{+}]+[d\phi_{-}]\bigr).$
Hence the mixed tensor rank of $\omega]$ is
$r=\operatorname{rank}_{\mathbb{R}}([\omega])=1.$
As shown in Section 3, the obstruction kernel $\mathcal{K}$ vanishes in our geometry, $\mathcal{K}=0$, so that the reduced rank is $r^{\sharp}=1$. Applying Theorem 5.1 we obtain the bound
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
In this Rashba–Dresselhaus example, the non-vanishing Chern number $c_{1}(\mathbb{L})$ already forces any section of $\mathbb{L}$ to develop vortices on $T^{2}_{\mathrm{BZ}}$; the standard argument shows that the total vortex number is at least $|c_{1}(\mathbb{L})|$. Thus, in this setting, the cohomological invariant $r^{\sharp}=1$ provides a complementary, algebraically refined constraint: it certifies that the curvature cannot be made block-diagonal, i.e., the momentum and phase degrees of freedom remain geometrically entangled. What is new in our framework is that the same mixed-rank invariant continues to impose a non-trivial holonomy lower bound even in situations where the total Chern number is tuned to zero, as we illustrate in Example 7.2 below.
7.2 Example 2: A configuration with vanishing total Chern number
The real advantage of the mixed-rank invariant becomes apparent when the total Chern number is zero. Consider a toy model where the two independent $U(1)$ sectors (associated with $\phi_{+}$ and $\phi_{-}$) carry Berry curvatures $F^{(+)}$ and $F^{(-)}$ that have opposite integrated fluxes:
$\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F^{(+)}=+1,\qquad\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F^{(-)}=-1.$
Then the total Chern number $c_{1}^{\,\mathrm{tot}}=\frac{1}{2\pi}\int(F^{(+)}+F^{(-)})=0$.
The synthetic torsion class is
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]\;+\;[F^{(-)}]\otimes[d\phi_{-}],$
where $F^{(+)}],[F^{(-)}]\in H^{2}(T^{2}_{\mathrm{BZ}})$ are both non-zero. Since $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, they are necessarily cohomologically proportional, i.e., $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$. Consequently, the mixed class can be written as a single simple tensor
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
and its mixed tensor rank is $r=1$. The obstruction kernel $\mathcal{K}$ still vanishes, as its vanishing depends only on the absence of parallel one-forms on the phase circles, not on the values of the Chern integrals. Hence $r^{\sharp}=1$, and the general bound gives
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1,$
exactly as in the previous example.
Now, however, the traditional Chern-number diagnostics would indicate no global obstruction ($c_{1}^{\,\mathrm{tot}}=0$), while the mixed-rank bound still forces an off-diagonal curvature operator. This means that, despite the vanishing net flux, the Berry curvature cannot be simultaneously flattened along the two phase directions. Such a configuration cannot be continuously deformed to a fully decoupled (product) gauge without altering the mixed cohomology class $\omega]$.
7.3 The role of the bound when $r\neq c_{1}^{\,\mathrm{tot}}$
The two examples above illustrate the distinct information carried by the global Chern number and by the mixed rank $r$ (or the reduced rank $r^{\sharp}$). The Chern number $c_{1}^{\,\mathrm{tot}}$ is a single integer obtained by integrating a curvature two-form over the whole Brillouin zone; it can vanish by cancellation of local contributions. The mixed tensor rank $r$, in contrast, counts how many independent tensorial couplings exist between the momentum torus and the two phase circles. Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the maximal possible value of $r$ in this geometry is $1$; this value is attained whenever both Berry curvature classes are non-zero, regardless of their integrals. It is a measure of the irreducible geometric entanglement between the factors of the product manifold.
The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq r^{\sharp}$ therefore constrains the algebraic structure of the curvature (the dimension of its off-diagonal part), a local property that is invisible to any integrated invariant. It is precisely this algebraic constraint that persists when the Chern number vanishes, as shown in Example 7.2.
7.4 Summary
The Rashba–Dresselhaus texture exemplifies how the mixed-cohomology framework reproduces known Chern-number bounds. The tuned model with $c_{1}^{\,\mathrm{tot}}=0$ demonstrates the framework’s true power: it reveals topological obstructions that remain hidden to conventional global invariants. The invariant $r^{\sharp}$ thus provides a finer topological classification of synthetic gauge fields, capable of detecting locally non-removable curvature even in the absence of net quantized flux.
8 Beyond the Chern Number: Local Rail-Tracks and Prospective Protocols
While $c_{1}$ counts the net vorticity (the integral of Berry curvature over the whole Brillouin torus), the lower bound on the off-diagonal holonomy dimension, $r^{\sharp}$, obstructs the local simultaneous annihilation of Berry curvature along independent directions in the product geometry. One may visualize the Berry curvature as being constrained to move along “rail-tracks” in the configuration space: even when the total topological charge can be tuned to zero, certain directional components of the curvature remain topologically locked and cannot be erased by any smooth gauge transformation.
In the two-component SOC BEC described by $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the bound $r^{\sharp}=1$ is non-degenerate because $\mathcal{K}=0$. Consequently, the curvature of any metric connection calibrated by the mixed class $\omega]$ cannot be made block-diagonal; at least one independent off-diagonal operator persists, mixing momentum directions with the phase directions $\phi_{+},\phi_{-}$.
This geometric prediction suggests concrete interferometric tests. Consider measuring the Berry phases $\Phi_{B}^{+}(\gamma)$ and $\Phi_{B}^{-}(\gamma)$ accumulated along a closed loop $\gamma\subset T^{2}_{\mathrm{BZ}}$ while separately addressing the global phase $\phi_{+}$ and the relative phase $\phi_{-}$. If the system were fully reducible, one could, in principle, choose a gauge that makes both $\Phi_{B}^{+}(\gamma)$ and $\Phi_{B}^{-}(\gamma)$ vanish simultaneously for every $\gamma$. The bound $r^{\sharp}=1$ forbids such a simultaneous gauging-away for all loops. In particular, for any pair of independent cycles $\gamma_{1},\gamma_{2}$ in $T^{2}_{\mathrm{BZ}}$, the two-component Berry-phase vector
$\bigl(\Phi_{B}^{+}(\gamma_{1}),\Phi_{B}^{-}(\gamma_{1}),\Phi_{B}^{+}(\gamma_{2}),\Phi_{B}^{-}(\gamma_{2})\bigr)$
cannot be continuously deformed to zero while preserving the mixed cohomology class $\omega]$.
Such measurements are within reach of present-day cold-atom experiments. For instance, the interferometric techniques used to map Berry curvature in Hofstadter bands [2] or to probe spin-texture topology in SOC BECs [9] can be adapted to separately track the two phase degrees of freedom $\phi_{+},\phi_{-}$ [5]. The required independent phase control can be implemented via the “painting” of arbitrary optical potentials [8], which provides precise spatiotemporal manipulation of the condensate’s phase profile. Observing that the two Berry-phase components cannot be simultaneously flattened would provide a direct experimental signature of the topological obstruction captured by $r^{\sharp}$.
Thus, the mixed-cohomology framework not only yields a rigorous mathematical lower bound but also points to a feasible protocol for detecting locally irremovable Berry curvature in ultracold quantum gases, going beyond the global Chern-number paradigm.
9 Conclusions
We have shown that a two-component spin–orbit-coupled BEC possesses a built-in geometric obstruction preventing the simultaneous flattening of Berry curvature along the two phase directions $\phi_{+}$ and $\phi_{-}$, even when all global Chern numbers vanish. The extended parameter space
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}},$
equipped with the natural Kaluza–Klein metric induced by the synthetic Berry connections, carries a metric connection $\nabla^{C}$ whose torsion encodes the synthetic gauge fields of the condensate. The harmonic part of this torsion defines a mixed cohomology class
$\omega]\in H^{2}(T^{2}_{\mathrm{BZ}})\otimes\bigl(H^{1}(S^{1}_{\phi_{+}})\oplus H^{1}(S^{1}_{\phi_{-}})\bigr),$
with mixed tensor rank $r=1$ for any generic SOC configuration in which the two Berry curvature classes are both non-zero. Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the two Berry curvature classes are necessarily cohomologically proportional, forcing the mixed class to be a single simple tensor. Moreover, the Kaluza–Klein geometry admits no parallel one-forms along the phase directions, so the obstruction kernel vanishes, $\mathcal{K}=0$, giving the reduced rank $r^{\sharp}=1$.
Applying the PT lower bound for connections with skew torsion yields the sharp inequality
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
This algebraic constraint on the curvature implies that the momentum degrees of freedom cannot be decoupled from the two-phase directions by any smooth gauge transformation. Importantly, the obstruction measured by $r^{\sharp}$ does not depend on the total Chern number: it persists even in regimes where global Berry flux cancels, a situation in which conventional topological invariants predict no restriction. The Rashba–Dresselhaus example illustrates compatibility with familiar Chern-number phenomena (e.g., $|c_{1}|=1$ in the simplest case), while also highlighting that the mixed-rank bound continues to operate beyond the Chern-number paradigm.
Experimentally, the bound suggests interferometric protocols capable of resolving the independent Berry phases associated with $\phi_{+}$ and $\phi_{-}$, thereby detecting the locally irremovable curvature. Conceptually, it identifies a new mechanism for topological robustness in synthetic gauge systems—one controlled not by global integrals of curvature but by the algebraic structure of its mixed components. We expect this framework to extend naturally to multicomponent cold-atom systems and synthetic-dimensional lattices where multiple phase or internal degrees of freedom couple to momentum-space geometry.
Acknowledgements
The authors gratefully acknowledge stimulating discussions with colleagues in geometry and cold-atom physics that helped shape the formulation of these results.

### v2

@@
 whose mixed tensor rank equals one.
-Using a general geometric bound for connections with skew torsion on product manifolds, we show that the obstruction kernel $\mathcal{K}$ vanishes, yielding the sharp inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$.
+Using a general geometric bound for metric connections with totally skew torsion on product manifolds, we show that the obstruction kernel $\mathcal{K}$ vanishes, yielding the sharp inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$.
 This forces at least one off-diagonal curvature operator, preventing complete gauging-away of Berry phases even when the total Chern number is zero.
@@
 Such a system possesses three universal, experimentally accessible degrees of freedom: the two crystal momenta forming a Brillouin-zone torus $T^{2}_{\mathrm{BZ}}$, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin components [9, 11].
-We therefore take as the extended parameter space the compact product manifold.
-$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}.$
+We therefore take as the extended parameter space the manifold
+$M=T_{BZ}^{2}\times S_{\phi_{+}}^{1}\times S_{\phi_{-}}^{1}.$
+We emphasize that treating $M$ as a global product manifold requires the underlying principal bundle to be trivial, which holds precisely in the physical regime of vanishing total Chern number ($c_{\text{tot}}=0$) that is central to our main result.
+In configurations with non-zero net topological charge, this product structure should be understood as a local description in a gauge patch, sufficient for deriving local curvature bounds.
 The first main task of the paper is to construct, in this space, a natural geometric structure to which the theorem of [12] applies.
@@
 Since the Berry curvatures $F^{(\pm)}$ are generically non-zero, we indeed have $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})=\{0\}$ in the physical geometry.
+By Lemma 4.5 (with $A^{(\pm)}$ non‑flat), the space of parallel 1‑forms on the phase circles vanishes: $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{M})=\{0\}$.
 Consequently the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(S^{1}_{\phi_{+}})\otimes\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ (the latter appears in $\ker\widetilde{\Psi}_{p}$) are zero.
@@
 4 Geometric foundations for the PT theorem in Kaluza–Klein geometry
-The physical metric (2.1) is not a Riemannian product, whereas Theorem 5.2 of PT [12] is stated for genuine product metrics.
-The following three results show that the theorem nevertheless applies to our Kaluza–Klein geometry, because the relevant topological and parallel-form data coincide with those of an auxiliary product metric.
+4.1 Deformation to a product metric and stability
+The physical Kaluza–Klein metric (2.1) is a warped structure that does not admit a global Riemannian product decomposition, whereas Theorem 5.2 in [12] is formally stated for genuine product metrics.
+To bridge this gap, we introduce a smooth deformation to a product metric and prove the stability of the holonomy bound.
+Definition 4.1 (Deformation family).
+Let $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ be the smooth one-parameter family of metrics on $M$ defined by:
+$g_{\varepsilon}=\pi^{*}g_{\mathrm{BZ}}+(d\phi_{+}+\varepsilon\pi^{*}A^{(+)})^{2}+(d\phi_{-}+\varepsilon\pi^{*}A^{(-)})^{2}.$
+For $\varepsilon=0$ we obtain the genuine Riemannian product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$, while for $\varepsilon=1$ we recover the physical Kaluza–Klein metric $g_{M}=g_{1}$.
+Lemma 4.2 (Stability of the obstruction data).
+For the family $g_{\varepsilon}$ in Definition 4.1, the following hold:
+  1. 1.
+    The spaces of parallel forms satisfy:
+    $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\varepsilon})=\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{0})=\mathbb{R}\operatorname{vol}_{\mathrm{BZ}}\quad\text{for all }\varepsilon\in[0,1].$
+  2. 2.
+    For $\varepsilon>0$, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{\varepsilon})=\{0\}$. For $\varepsilon=0$, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{0})=\mathbb{R}d\phi_{\pm}$.
+  3. 3.
+    The mixed tensor rank $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is constant for all $\varepsilon$.
+  4. 4.
+    The obstruction kernel $\mathcal{K}_{\varepsilon}$ satisfies $\dim\mathcal{K}_{\varepsilon}=0$ for all $\varepsilon\in[0,1]$.
+Proof.
+(1) follows because the restriction of $g_{\varepsilon}$ to $T^{2}_{\mathrm{BZ}}$ is $g_{\mathrm{BZ}}$, independent of $\varepsilon$.
+(2) For $\varepsilon>0$, Lemma 2.4 shows $d\phi_{\pm}$ are not parallel; the proof of Lemma 4.5 will show no other parallel 1-forms exist.
+For $\varepsilon=0$, $d\phi_{\pm}$ are clearly parallel.
+(3) is Lemma 4.6 (metric invariance).
+(4) For $\varepsilon=0$, the harmonic representative of $\omega]$ is
+$\omega_{h}=F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}.$
+The subspace $\mathcal{P}_{2}(T^{2})\otimes\mathcal{P}_{1}(S^{1}_{+}\times S^{1}_{-})$ consists of those harmonic 3‑forms that can be written as
+$\alpha\wedge\beta,$
+where $\alpha$ is a parallel 2‑form on $T^{2}$ (hence a constant multiple of $\text{vol}_{\text{BZ}}$) and $\beta$ is a parallel 1‑form on the fibres (hence a constant linear combination of $d\phi_{+}$ and $d\phi_{-}$).
+If $\omega_{h}$ belonged to this subspace, we would have
+$F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}=\alpha\wedge\beta=(\lambda\text{vol}_{\text{BZ}})\wedge(c_{+}d\phi_{+}+c_{-}d\phi_{-}),$
+which would imply $F^{(+)}=\lambda c_{+}\text{vol}_{\text{BZ}}$ and $F^{(-)}=\lambda c_{-}\text{vol}_{\text{BZ}}$ as differential forms, not just in cohomology.
+In a generic SOC texture, $F^{(+)}$ and $F^{(-)}$ are not parallel (they are not constant multiples of $\text{vol}_{\text{BZ}}$ pointwise), hence $\omega_{h}\notin\mathcal{P}_{2}\otimes\mathcal{P}_{1}$.
+Therefore $\mathcal{K}_{0}=0$.
+∎
+Lemma 4.3 (Lower semi-continuity of off-diagonal holonomy dimension).
+The function $F(\varepsilon):=\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower semi-continuous on $[0,1]$, where $\nabla^{C_{\varepsilon}}$ is the metric connection with fixed torsion $T$ constructed from $g_{\varepsilon}$.
+Proof.
+The curvature tensor $R^{C_{\varepsilon}}$ depends smoothly on $\varepsilon$. At any point $p\in M$, consider the evaluation map
+$\Phi_{\varepsilon}:\bigwedge^{2}T_{p}M\times T_{p}M\to\mathfrak{so}(T_{p}M),\quad\Phi_{\varepsilon}(X,Y,Z)=R^{C_{\varepsilon}}(X,Y)Z.$
+The off-diagonal holonomy algebra at $p$ is generated by the image of $\Phi_{\varepsilon}$ restricted to arguments that yield off-diagonal operators.
+The dimension of a finitely generated subspace is a lower semi-continuous function of the generating set (it can only jump upward at discontinuity points).
+Since $\Phi_{\varepsilon}$ varies continuously in $\varepsilon$, $F(\varepsilon)$ is lower semi-continuous.
+∎
+Proposition 4.4 (Persistence of the PT bound under deformation).
+If the Pigazzini–Toda theorem gives $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r^{\sharp}$ for the product metric $g_{0}$, then the same inequality holds for all $\varepsilon\in[0,1]$, and in particular for the physical metric $g_{M}=g_{1}$.
+Proof.
+By Lemma 4.2, the topological data $r^{\sharp}=r-\dim\mathcal{K}_{\varepsilon}$ is constant (=r=r) for all $\varepsilon$.
+The inequality holds at $\varepsilon=0$ by assumption.
+By Lemma 4.3, the set $\{\varepsilon:\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq r^{\sharp}\}$ is closed (by lower semi-continuity).
+Since it contains $\varepsilon=0$ and $[0,1]$ is connected, it must be all of $[0,1]$.
+∎
 4.2 Parallel forms in Kaluza–Klein metrics
-Lemma 4.1 (Parallel forms in Kaluza–Klein metrics with flat fibres).
+Lemma 4.5 (Parallel forms in Kaluza–Klein metrics).
 Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a Kaluza–Klein metric
-$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$ (4.1)
-where $g_{1}$ is a Riemannian metric on $M_{1}$, $\{\theta^{a}\}$ defines a flat left-invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$ (possibly with $dA^{a}\neq 0$).
-Then the restriction map
-$\Phi_{1}:\mathcal{P}_{k}(M_{1},g)\longrightarrow\mathcal{P}_{k}(M_{1},g_{1}),\qquad\eta\mapsto\eta|_{D_{1}},$
-is a well-defined linear isomorphism.
-Analogously, define $\mathcal{P}_{k}(M_{2},g)$ as the space of forms $\xi\in\Omega^{k}(M)$ that are invariant along $M_{1}$ (their coefficients depend only on the $M_{2}$-coordinates) and satisfy $\nabla^{LC}\xi=0$.
-Then
-$\Phi_{2}:\mathcal{P}_{k}(M_{2},g)\longrightarrow\mathcal{P}_{k}(M_{2},h),\qquad\xi\mapsto\xi|_{D_{2}},$
-is an isomorphism.
+$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$
+where $g_{1}$ is a Riemannian metric on $M_{1}$, $\{\theta^{a}\}$ defines a flat left-invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$.
+  1. 1.
+    If all connection forms $A^{a}$ are closed (i.e., $dA^{a}=0$), then every $h$-parallel form on the fibre is also $g$-parallel.
+  2. 2.
+    If at least one curvature $F^{a}:=dA^{a}$ is not identically zero, then the space of $g$-parallel 1-forms tangent to the fibres is trivial:
+    $\mathcal{P}_{1}(M_{2},g)=\{0\}.$
 Proof.
-Choose local coordinates $(x^{i},y^{\alpha})$ adapted to the orthogonal splitting $TM=D_{1}\oplus D_{2}$ induced by the product structure. In these coordinates the metric reads
-$g=g_{ij}(x)\,dx^{i}dx^{j}+h_{\alpha\beta}(y)\bigl(dy^{\alpha}+A^{\alpha}_{i}(x)dx^{i}\bigr)\bigl(dy^{\beta}+A^{\beta}_{j}(x)dx^{j}\bigr).$ (4.2)
-Let $\eta\in\mathcal{P}_{k}(M_{1},g)$. Because $\eta$ is parallel and tangent to $D_{1}$, in the coordinates above it can be written as
-$\eta=\frac{1}{k!}\,\eta_{i_{1}\dots i_{k}}(x)\,dx^{i_{1}}\wedge\dots\wedge dx^{i_{k}}.$
-We compute the Christoffel symbols of $g$. From (4.2) one obtains
-$\Gamma^{i}_{jk}$ $=\Gamma^{(g_{1})i}_{jk},$
-$\Gamma^{i}_{j\beta}$ $=-\frac{1}{2}\,g^{i\ell}\bigl(\partial_{j}g_{\ell\beta}+\partial_{\beta}g_{j\ell}-\partial_{\ell}g_{j\beta}\bigr),$
-$\Gamma^{\alpha}_{jk}$ $=\frac{1}{2}\,h^{\alpha\gamma}\bigl(\partial_{j}g_{k\gamma}+\partial_{k}g_{j\gamma}-\partial_{\gamma}g_{jk}\bigr),$
-$\Gamma^{\alpha}_{\beta\gamma}$ $=\Gamma^{(h)\alpha}_{\beta\gamma},$
-where $\Gamma^{(g_{1})}$ and $\Gamma^{(h)}$ denote the Christoffel symbols of $g_{1}$ and $h$ respectively.
-For any vector field $X=X^{i}\partial_{i}+X^{\alpha}\partial_{\alpha}$,
-$(\nabla^{LC}_{X}\eta)_{i_{1}\dots i_{k}}=X^{j}\partial_{j}\eta_{i_{1}\dots i_{k}}-\sum_{r=1}^{k}\Gamma^{\ell}_{ji_{r}}X^{j}\eta_{i_{1}\dots i_{r-1}\ell i_{r+1}\dots i_{k}}-\Gamma^{\alpha}_{j\alpha}X^{j}\eta_{i_{1}\dots i_{k}}.$
-The condition $\nabla^{LC}\eta=0$ therefore splits into two independent sets of equations. For $X=X^{i}\partial_{i}$ (tangent to $M_{1}$) we obtain
-$\nabla^{(g_{1})}_{X}\eta=0,$
-where $\nabla^{(g_{1})}$ is the Levi‑Civita connection of $g_{1}$. Hence $\eta|_{D_{1}}$ is $g_{1}$-parallel.
-For $X=X^{\alpha}\partial_{\alpha}$ (tangent to $M_{2}$) we get
-$\Gamma^{\alpha}_{j\alpha}(x)X^{j}\eta_{i_{1}\dots i_{k}}=0\qquad\text{for all }X^{j}.$
-Why $\Gamma^{\alpha}_{j\alpha}=0$.
-Because the fibres are flat, the fibre metric $h_{\alpha\beta}$ is constant in the chosen coordinates; therefore $\partial_{j}h_{\alpha\beta}=0$.
-Moreover, from (4.2) we have $g_{j\alpha}=h_{\alpha\beta}A^{\beta}_{j}(x)$, which is independent of $y$.
-Consequently
-$\partial_{\gamma}g_{j\alpha}=0,\qquad\partial_{\alpha}g_{j\gamma}=0.$
-Substituting these facts into the expression for $\Gamma^{\alpha}_{j\alpha}$ gives
-$\Gamma^{\alpha}_{j\alpha}=\frac{1}{2}h^{\alpha\gamma}\bigl(\partial_{j}h_{\alpha\gamma}+\partial_{\alpha}g_{j\gamma}-\partial_{\gamma}g_{j\alpha}\bigr)=0.$
-Thus the condition coming from $X^{\alpha}\partial_{\alpha}$ is automatically satisfied.
-Hence $\eta\in\mathcal{P}_{k}(M_{1},g)$ if and only if $\eta|_{D_{1}}\in\mathcal{P}_{k}(M_{1},g_{1})$.
-Consequently $\Phi_{1}$ is injective.
-Surjectivity follows because any $g_{1}$-parallel form $\widetilde{\eta}$ on $M_{1}$ can be lifted to a $g$-parallel form on $M$ by declaring it independent of the fibre coordinates; the computation above shows that the lifted form is indeed parallel with respect to $g$.
-Therefore $\Phi_{1}$ is an isomorphism. The proof for $\Phi_{2}$ proceeds identically after exchanging the roles of base and fibre. ∎
-Lemma 4.2 (Harmonic Künneth decomposition for flat‑fibre KK metrics).
-Let $g$ be a Kaluza–Klein metric as in (4.1) with flat fibres. Then the natural wedge‑product map
-$\Phi:\mathcal{H}^{p}(M_{1},g_{1})\otimes\mathcal{H}^{q}(M_{2},h)\longrightarrow\mathcal{H}^{p+q}(M,g)$
-is an isomorphism onto the subspace of harmonic forms of pure bidegree $(p,q)$.
-In particular, every $g$-harmonic form $\eta\in\mathcal{H}^{p+q}(M,g)$ that belongs to the $(p,q)$-Künneth component admits a unique decomposition
-$\eta=\sum_{i=1}^{r}\alpha_{i}\wedge\beta_{i},\qquad\alpha_{i}\in\mathcal{H}^{p}(M_{1},g_{1}),\;\beta_{i}\in\mathcal{H}^{q}(M_{2},h),$
-with $r=\operatorname{rank}_{\mathbb{R}}([\eta]_{\mathrm{mixed}})$.
+Write the metric locally as $g=g_{ij}(x)dx^{i}dx^{j}+\sum_{a=1}^{m}(\Theta^{a})^{2}$, where $\Theta^{a}=\theta^{a}+A^{a}_{i}(x)dx^{i}$ and $\{\theta^{a}\}$ is the orthonormal coframe on $M_{2}$ induced by the flat metric $h$.
+In our SOC BEC case, $m=2$, $\theta^{1}=d\phi_{+}$, $\theta^{2}=d\phi_{-}$.
+Let $\xi=\sum_{a=1}^{m}f_{a}(x,y)\Theta^{a}$ be a 1‑form tangent to the fibres.
+The Levi‑Civita connection of $g$ satisfies the following relation for a horizontal vector field $X=X^{i}\partial_{i}$ and the vertical coordinate field $V_{a}$ dual to $\theta^{a}$:
+$\nabla^{LC}_{X}V_{a}=\tfrac{1}{2}F^{a}(X)^{\#},$
+where $F^{a}(X)^{\#}$ is the vertical vector field obtained by raising the index of the 1‑form $F^{a}(X,\cdot)$ with the fibre metric $h$.
+Formula (i) follows from the Koszul formula and uses only that the fibres are flat.
+Now suppose $\xi$ is parallel, $\nabla^{LC}\xi=0$. Evaluating on the pair $(X,V_{a})$ gives
+$0=(\nabla^{LC}_{X}\xi)(V_{a})=X\bigl(\xi(V_{a})\bigr)-\xi(\nabla^{LC}_{X}V_{a})=X(f_{a})-\tfrac{1}{2}\sum_{b=1}^{m}f_{b}\,h\!\bigl(F^{a}(X)^{\#},V_{b}\bigr).$
+Because the functions $f_{a}$ are constant along the fibres (they depend only on the base coordinates $x$ due to the left‑invariance of the metric $h$), equation (ii) reduces to
+$\tfrac{1}{2}\sum_{b=1}^{m}f_{b}(x)\,F^{b}(X,V_{a})=0\qquad\text{for every horizontal }X\text{ and every }a.$
+If all $F^{b}=0$, condition (iii) is vacuous and any choice of constant $f_{b}$ yields a parallel form, proving statement (i).
+If, however, some $F^{b_{0}}$ is not identically zero, we can choose a horizontal vector field $X$ and an index $a$ such that $F^{b_{0}}(X,V_{a})\neq 0$ at some point; then (iii) forces $f_{b_{0}}=0$.
+Repeating the argument for every independent curvature forces all coefficients $f_{b}$ to vanish, hence $\xi=0$.
+This proves statement (ii) for 1‑forms.
+The argument for higher‑degree forms is analogous, using that a parallel $k$-form on the fibre must have constant coefficients with respect to the left‑invariant coframe, and the same curvature coupling appears in the covariant derivative.
+∎
+Lemma 4.6 (Harmonic representative and Künneth component).
+Let $g$ be a Kaluza–Klein metric as in Definition 4.1 on $M=M_{1}\times M_{2}$.
+Let $\omega]\in H^{k}(M;\mathbb{R})$ be a cohomology class and $\omega_{h}$ its unique $g$-harmonic representative.
+Then $\omega_{h}$ respects the Künneth decomposition in the sense that its projection onto the $(p,q)$-bidegree component corresponds to the harmonic representative of the mixed class $\omega]_{p,q}\in H^{p}(M_{1})\otimes H^{q}(M_{2})$.
+In particular, $\omega_{h}$ admits a unique decomposition
+$\omega_{h}=\sum_{i=1}^{r}\alpha_{i}\wedge\beta_{i}+\gamma,$
+where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$, the forms $\alpha_{i},\beta_{i}$ are the lifted harmonic representatives of the factor classes, and $\gamma$ contains terms of different bidegrees.
 Proof.
-Because the fibres are flat, the Hodge Laplacian of $g$ decouples as $\Delta_{g}=\Delta_{g_{1}}\otimes\mathrm{Id}+\mathrm{Id}\otimes\Delta_{h}$ on forms of pure bidegree.
-Hence a form of bidegree $(p,q)$ is $g$-harmonic iff its components are harmonic with respect to $g_{1}$ and $h$ respectively.
-The statement then follows from the classical Künneth theorem for de Rham cohomology combined with Hodge theory.
+By Hodge theory on compact Riemannian manifolds, every de Rham cohomology class $\omega]$ possesses a unique harmonic representative $\omega_{h}$ with respect to the metric $g$.
+The Künneth theorem provides a canonical isomorphism $H^{k}(M)\cong\bigoplus_{p+q=k}H^{p}(M_{1})\otimes H^{q}(M_{2})$.
+While the Hodge Laplacian $\Delta_{g}$ of a Kaluza-Klein metric does not generally commute with the factor-wise Laplacians $\Delta_{g_{1}}$ and $\Delta_{h}$ due to the presence of the connection forms $A^{a}$, the existence of the harmonic representative for each mixed component is guaranteed by the ellipticity of $\Delta_{g}$.
+The tensor rank $r$ is a purely algebraic property of the class $\omega]$ in the tensor product of the cohomology groups $H^{*}(M_{i})$ and is therefore independent of the specific metric $g$.
+The decomposition follows by projecting the global harmonic form $\omega_{h}$ onto the subspace of forms of pure bidegree, which remains well-defined under the orthogonal splitting $TM=D_{1}\oplus D_{2}$ induced by the Kaluza-Klein structure.
 ∎
 4.3 Persistence of non‑parallelism under metric change
-For the application of the PT theorem to an auxiliary product metric, we need to ensure that the non‑parallelism condition $\nabla^{LC}T\neq 0$ is preserved when we replace the original Kaluza–Klein metric by a nearby product metric.
+To ensure that the lower bound on the off-diagonal holonomy is a robust feature of the model rather than a consequence of a specific metric choice, we must examine the stability of the non-parallelism condition $\nabla^{LC}T\neq 0$.
+The following result establishes that this property is open in the space of Riemannian metrics, guaranteeing that the topological obstruction remains effective under generic deformations of the Kaluza–Klein geometry.
 Lemma 4.7 (Openness of the non‑parallelism condition).
@@
 Proof.
-The Levi‑Civita connection $\nabla^{LC}(g)$ depends smoothly (in fact, rationally) on the metric $g$ and its first derivatives.
-Consequently, the tensor field $\nabla^{LC}(g)T$ depends continuously on $g$ in the $C^{1}$-topology.
-The condition $\nabla^{LC}(g)T\neq 0$ means that at least one component of this tensor is non‑zero at some point; by continuity, the same component remains non‑zero for all metrics sufficiently close to $g$ in the $C^{1}$-sense.
+The Levi‑Civita connection $\nabla^{LC}(g)$ depends smoothly on the components of the metric $g$ and its first derivatives through the Christoffel symbols.
+Consequently, the map $g\mapsto\nabla^{LC}(g)T$ is continuous with respect to the $C^{1}$-topology on $\mathcal{M}$.
+Since the condition $\nabla^{LC}(g)T\neq 0$ defines the complement of the zero set of a continuous map, $\mathcal{M}_{\neq 0}(T)$ is an open subset of $\mathcal{M}$.
 ∎
 Remark 4.8.
-In our setting, the original metric $g$ is the Kaluza–Klein metric (4.1) and the auxiliary product metric $g_{\mathrm{prod}}=g_{1}\oplus h$ can be chosen arbitrarily close to $g$ in the $C^{\infty}$ (hence also $C^{1}$) topology by taking the gauge potentials $A^{a}$ small, or by a direct approximation argument.
-Lemma 4.3 guarantees that if $\nabla^{LC}(g)T\neq 0$, then also $\nabla^{LC}(g_{\mathrm{prod}})T\neq 0$ for a generic choice of $g_{\mathrm{prod}}$ near $g$.
+In the context of spin-orbit coupled BECs, Lemma 4.7 ensures the structural stability of our topological bound.
+Since the Kaluza–Klein metric $g$ satisfies $\nabla^{LC}(g)T\neq 0$ due to the non-triviality of the synthetic gauge curvature, any sufficiently small perturbation of the metric—arising, for instance, from experimental fluctuations in the trapping potentials or the SOC strength—will preserve the non-parallelism of the torsion.
+This confirms that the lower bound on the off-diagonal holonomy is not a fine-tuned property but a robust feature of the SOC BEC geometry.
 4.4 Harmonic representatives in Kaluza–Klein geometry
-Lemma 4.5 (Invariance of the mixed tensor rank under Kaluza–Klein metrics).
-Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold, and let $g$ be any Riemannian metric on $M$ (in particular, a Kaluza–Klein metric of the form (4.1) with flat fibres).
-For a cohomology class
-$\omega]\in\bigl(H^{2}(M_{1})\otimes H^{1}(M_{2})\bigr)\oplus\bigl(H^{1}(M_{1})\otimes H^{2}(M_{2})\bigr)\subset H^{3}(M),$
-its mixed tensor rank
-$r:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$
-depends only on the topological class $\omega]$; in particular, it is the same whether computed with respect to the metric $g$ or with respect to the genuine product metric $g_{\mathrm{prod}}=g_{1}\oplus h$.
+Lemma 4.9 (Metric invariance of the mixed tensor rank).
+Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold.
+For any cohomology class $\omega]\in H^{k}(M;\mathbb{R})$, let $\omega]_{\mathrm{mixed}}$ be its projection onto the mixed components of the Künneth decomposition
+$\bigoplus_{p+q=k,p,q\geq 1}H^{p}(M_{1};\mathbb{R})\otimes H^{q}(M_{2};\mathbb{R}).$
+The mixed tensor rank $r:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is a topological invariant of the class $\omega]$.
+In particular, $r$ is independent of the choice of the Riemannian metric $g$ on $M$.
 Proof.
-The mixed tensor rank is defined algebraically as the minimal number of simple tensors needed to represent the element $\omega]$ in the finite‑dimensional vector space
-$\bigl(H^{2}(M_{1})\otimes H^{1}(M_{2})\bigr)\oplus\bigl(H^{1}(M_{1})\otimes H^{2}(M_{2})\bigr).$
-This rank is independent of any auxiliary Riemannian metric, because the cohomology spaces $H^{p}(M_{i})$ themselves are topological invariants.
-Hence, for any two metrics $g$ and $g^{\prime}$ on $M$, the mixed tensor rank of $\omega]$ coincides.
+The mixed tensor rank is defined algebraically as the minimal number of simple tensors (elementary tensors of the form $\alpha\otimes\beta$) required to represent $\omega]_{\mathrm{mixed}}$ within the finite-dimensional vector space $\bigoplus H^{p}(M_{1})\otimes H^{q}(M_{2})$.
+Since the de Rham cohomology groups $H^{p}(M_{i};\mathbb{R})$ are topological invariants associated with the factor manifolds $M_{i}$, the algebraic structure of their tensor product is fixed once the manifold topology is given.
+Because the projection of a cohomology class onto its Künneth components is canonical and independent of any Riemannian structure, the rank $r$ remains invariant under any smooth deformation of the metric $g$ on $M$.
+This ensures that the value of $r$ computed for a Kaluza–Klein metric coincides with the value assigned to the underlying topological class.
 ∎
@@
 Define the obstruction space
-$cK\;:=\;\bigl(\pi_{2,1}([\omega_{h}])\cap(cP_{2}(M_{1},g)\otimes cP_{1}(M_{2},g))\bigr)\oplus\bigl(\pi_{1,2}([\omega_{h}])\cap(cP_{1}(M_{1},g)\otimes cP_{2}(M_{2},g))\bigr),$
-and set $r^{\sharp}:=r-\dim cK$.
+$\mathcal{K}\;:=\;\bigl(\pi_{2,1}([\omega_{h}])\cap(\mathcal{P}_{2}(M_{1},g)\otimes\mathcal{P}_{1}(M_{2},g))\bigr)\oplus\bigl(\pi_{1,2}([\omega_{h}])\cap(\mathcal{P}_{1}(M_{1},g)\otimes\mathcal{P}_{2}(M_{2},g))\bigr),$
+and set $r^{\sharp}:=r-\dim\mathcal{K}$.
 Then for every point $p\in M$,
-$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r^{\sharp}.$ (4.3)
-Moreover, $r^{\sharp}$ depends only on the cohomology class $\omega_{h}]_{\mathrm{mixed}}$ and on the spaces $cP_{k}(M_{i},g)$; consequently, it is invariant under smooth deformations of $g$ that preserve those spaces.
+$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r^{\sharp}.$ (4.0)
+Moreover, $r^{\sharp}$ depends only on the cohomology class $\omega_{h}]_{\mathrm{mixed}}$ and on the spaces $\mathcal{P}_{k}(M_{i},g)$; consequently, it is invariant under smooth deformations of $g$ that preserve those spaces.
 Proof.
-We proceed in several carefully justified stages.
-Consider the genuine Riemannian product metric $g_{\mathrm{prod}}:=g_{1}\oplus h$ on $M=M_{1}\times M_{2}$.
-Define a new metric connection on $(M,g_{\mathrm{prod}})$ by $\widetilde{\nabla}^{C}:=\nabla^{LC}(g_{\mathrm{prod}})+\tfrac{1}{2}T$, where $T$ is the same 3‑form as in the statement.
-Since $\nabla^{LC}(g_{\mathrm{prod}})$ is torsion‑free and $T$ is totally skew, $\widetilde{\nabla}^{C}$ is metric with respect to $g_{\mathrm{prod}}$ and has torsion $T$.
-Because the fibres of the Kaluza–Klein metric $g$ are flat, Lemma 4.1 provides canonical isomorphisms
-$\Phi_{i}^{(k)}:cP_{k}(M_{i},g)\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}cP_{k}(M_{i},g_{\mathrm{prod}}),\qquad i=1,2,\;k=1,2,$
-obtained by restricting a parallel form to the corresponding factor.
-Consequently, the obstruction spaces computed with $g$ and with $g_{\mathrm{prod}}$ have identical dimensions: $\dim cK=\dim\widetilde{cK}$, where $\widetilde{cK}$ is defined analogously to $cK$ but using the spaces $cP_{k}(M_{i},g_{\mathrm{prod}})$.
-The mixed tensor rank $r=rank_{\mathbb{R}}\bigl([\omega_{h}]_{\mathrm{mixed}}\bigr)$ is a purely topological invariant of the cohomology class $\omega_{h}]$ (Lemma 4.5); therefore it is unchanged when we replace $g$ by $g_{\mathrm{prod}}$.
-Let $\widetilde{\omega}_{h}$ be the unique $g_{\mathrm{prod}}$‑harmonic representative of the same class $\omega_{h}]$.
-By Hodge theory, $\widetilde{\omega}_{h}]=[\omega_{h}]$ and consequently $rank_{\mathbb{R}}\bigl([\widetilde{\omega}_{h}]_{\mathrm{mixed}}\bigr)=r$.
-The triple $(M,g_{\mathrm{prod}},\widetilde{\nabla}^{C})$ satisfies all hypotheses of Theorem 5.2 in [12].
-Indeed, $g_{\mathrm{prod}}$ is a genuine product metric; $\widetilde{\nabla}^{C}$ is a metric connection on $(M,g_{\mathrm{prod}})$ with totally skew torsion $T$; the condition $\nabla^{LC}(g)T\neq 0$ implies $\nabla^{LC}(g_{\mathrm{prod}})T\neq 0$ in the generic situation we consider—the difference $\nabla^{LC}(g)-\nabla^{LC}(g_{\mathrm{prod}})$ is a tensor field, and if $\nabla^{LC}(g)T$ were to vanish while $\nabla^{LC}(g_{\mathrm{prod}})T$ remained non‑zero, this would impose a special algebraic relation between the metrics and $T$ that we exclude as nongeneric; finally, the harmonic part $\widetilde{\omega}_{h}$ of $T$ with respect to $g_{\mathrm{prod}}$ represents a mixed cohomology class of tensor rank $r$.
-Applying Theorem 5.2 we obtain
-$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\widetilde{\nabla}^{C})\geq r-\dim\widetilde{cK}=r^{\sharp}.$ (4.4)
-Write the two connections as $\nabla^{C}=\nabla^{LC}(g)+K$ and $\widetilde{\nabla}^{C}=\nabla^{LC}(g_{\mathrm{prod}})+K$ with $K_{X}Y=\tfrac{1}{2}T(X,Y,\cdot)^{\sharp}$.
-Their curvature tensors are given by $R^{\nabla}=R^{\nabla^{LC}}+\nabla^{LC}K+K\circ K$.
-The off‑diagonal part with respect to the splitting $TM=D_{1}\oplus D_{2}$ arises from terms where the arguments $X,Y$ belong to different factors.
-Because the Levi‑Civita curvature of a product metric is block‑diagonal, $R^{\nabla^{LC}(g_{\mathrm{prod}})}_{\mathrm{off}}=0$.
-For a general metric $g$, the difference $\Delta:=\nabla^{LC}(g)-\nabla^{LC}(g_{\mathrm{prod}})$ is a tensor field whose Christoffel symbols in adapted coordinates satisfy $\Delta^{\alpha}_{ij}=0$ and $\Delta^{i}_{j\alpha}=0$; the only non‑zero components mix one base and one fibre index.
-Consequently, when we compute the off‑diagonal component of $\nabla^{LC}(g)K$, the contribution of $\Delta$ takes the form $(\Delta_{X}K)_{Y}Z-(\Delta_{Y}K)_{X}Z$ with $X$ and $Y$ tangent to different factors.
-A direct inspection in coordinates shows that such terms are linear combinations of the components of $K$ itself, evaluated at the point $p$.
-Hence they do not increase the dimension of the space spanned by the off‑diagonal curvature operators beyond the dimension already produced by the $\nabla^{LC}(g_{\mathrm{prod}})K$ term together with the quadratic term $K\circ K$.
-More formally, consider the linear map $\Psi_{p}:cV_{2,1}\to so(T_{p}M)$ defined in the proof of Theorem 5.2 of [12], which associates to each simple mixed tensor $\alpha\otimes\beta$ the off‑diagonal curvature operator generated by $\nabla^{LC}K$.
-Replacing $\nabla^{LC}(g_{\mathrm{prod}})$ by $\nabla^{LC}(g)$ changes $\Psi_{p}$ by an additive linear map whose image lies in the span of the operators obtained by evaluating $K$ itself on vectors from different factors.
-Because the quadratic term $K\circ K$ already accounts for all such algebraic combinations, the rank of the total off‑diagonal curvature cannot decrease when we replace $g_{\mathrm{prod}}$ by $g$; it may only increase.
-Therefore, for every $p\in M$,
-$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq\dim\mathrm{hol}^{\mathrm{off}}_{p}(\widetilde{\nabla}^{C}).$ (4.5)
-Combining (4.4) with (4.5) yields precisely
-$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp},$
-which is the desired estimate (4.3).
-The integer $r^{\sharp}=r-\dim cK$ depends solely on the topological cohomology class $\omega_{h}]_{\mathrm{mixed}}$ (which determines $r$) and on the spaces $cP_{k}(M_{i},g)$ (which determine $\dim cK$).
-Any smooth deformation of $g$ that leaves these parallel‑form spaces unchanged therefore leaves $r^{\sharp}$ invariant.
-This completes the proof.
+The result follows from a direct application of the topological bound established in [12] to the Kaluza–Klein geometry $(M,g,\nabla^{C})$.
+We verify that the structural requirements are met and compute the resulting invariant.
+First, we observe that the metric connection $\nabla^{C}$ satisfies the requisite non-degeneracy conditions.
+The torsion form $T$, constructed from the synthetic gauge fields, is totally skew-symmetric.
+Moreover, the condition $\nabla^{LC}(g)T\neq 0$ is satisfied whenever the Berry curvature $F=dA\neq 0$, which holds by assumption in the generic SOC regime.
+The mixed tensor rank $r=\text{rank}_{\mathbb{R}}([\omega_{h}]_{\mathrm{mixed}})$ is a cohomological invariant of the class $[T]$; it is independent of the Riemannian metric used to define the harmonic representative, ensuring that $r$ is well-defined and finite.
+Crucially, the dimension of the obstruction kernel $\mathcal{K}$ is determined by the intersection of the mixed cohomology components with the spaces of parallel forms.
+By Lemma 4.5, the non-vanishing Berry curvature $F=dA\neq 0$ implies that the space of fibre-parallel 1-forms vanishes:
+$\mathcal{P}_{1}(M_{2},g)=\{0\}.$
+Consequently, any tensor product involving $\mathcal{P}_{1}(M_{2},g)$ is zero, and the obstruction space $\mathcal{K}$ is identically zero in the generic SOC BEC regime where synthetic gauge fields are present.
+Thus, we have $\dim\mathcal{K}=0$, and the general theorem of [12] then states that the dimension of the off-diagonal holonomy is bounded by $r-\dim\mathcal{K}$, yielding
+$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r.$
+The invariance of this bound follows from the stability of the topological rank $r$ and the robustness of the vanishing condition $\mathcal{P}_{1}(M_{2},g)=\{0\}$ under metric deformations that preserve the non-integrability of the horizontal distribution.
 ∎
@@
 Proposition 4.11 (Application to the SOC BEC model).
-In the spin‑orbit‑coupled BEC model of Section 2, the extended parameter space is $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$.
-The metric (2.1) is precisely of the Kaluza–Klein type (4.1) with $M_{1}=T^{2}_{\mathrm{BZ}}$, $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, left‑invariant forms $\theta^{1}=d\phi_{+},\ \theta^{2}=d\phi_{-}$, and gauge potentials $A^{1}=A^{(+)},\ A^{2}=A^{(-)}$.
-The fibres are flat because $d\theta^{1}=d\theta^{2}=0$.
-For this geometry:
+In the spin‑orbit‑coupled BEC model described in Section 2, the extended parameter space is given by the product manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$.
+The metric (2.1) is of the Kaluza–Klein type $g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2}$ with the base $M_{1}=T^{2}_{\mathrm{BZ}}$ and the fibre $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$.
+The left‑invariant forms are $\theta^{1}=d\phi_{+}$ and $\theta^{2}=d\phi_{-}$, with corresponding gauge potentials $A^{1}=A^{(+)}$ and $A^{2}=A^{(-)}$.
+The fibres are flat since $d\theta^{a}=0$.
+For this specific geometry, the following properties hold:
   1. 1.
-    $\mathcal{P}_{1}(M_{2},g)=\{0\}$ (Corollary 2.5).
+    $\mathcal{P}_{1}(M_{2},g)=\{0\}$, due to the non-triviality of the synthetic gauge potentials $A^{(\pm)}$.
   2. 2.
@@
   3. 3.
-    Consequently the obstruction kernel vanishes: $\mathcal{K}=0$.
-If the Berry curvatures satisfy $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}})$, then the mixed tensor rank attains its maximal possible value for this geometry, $r=1$, and Theorem 4.6 gives
+    The obstruction space vanishes identically: $\mathcal{K}=0$.
+Furthermore, if the Berry curvatures satisfy $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}})$, then the mixed tensor rank is $r=1$, and Theorem 4.10 implies:
 $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
 Proof.
-The identification with the Kaluza–Klein framework is immediate from the definitions.
-Item 1 follows from Corollary 2.5, which shows that no non‑zero linear combination of $d\phi_{+}$ and $d\phi_{-}$ is parallel when the potentials $A^{(\pm)}$ are not constant.
-Item 2 is a direct consequence of Lemma 4.1 applied to the base factor.
-Since $\mathcal{P}_{1}(M_{2},g)=0$, the tensor products appearing in the definition of $\mathcal{K}$ are zero, hence $\mathcal{K}=0$.
-For the last statement, recall that $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$.
-If both $F^{(+)}]$ and $F^{(-)}]$ are non‑zero, they are necessarily proportional in this one‑dimensional space; therefore the mixed class
-$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$
-has tensor rank $r=1$ (it can be written as a single simple tensor $F^{(+)}]\otimes([d\phi_{+}]+\lambda[d\phi_{-}])$ for some $\lambda\in\mathbb{R}$).
-Substituting $r=1$ and $\dim\mathcal{K}=0$ into (4.3) yields the claimed lower bound.
-∎
+The identification of the BEC metric with the Kaluza–Klein framework follows directly from the construction in Section 2.
+The first item is a direct consequence of Lemma 4.5; as established in the physical model, the synthetic gauge potentials $A^{(\pm)}$ possess non-vanishing Berry curvature, which acts as an obstruction to the existence of parallel 1-forms along the fibres.
+On the base manifold $M_{1}=T^{2}_{\mathrm{BZ}}$, the restriction map onto $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ is an isomorphism, identifying the space of parallel 2-forms with the span of the Riemannian volume form.
+Since $\mathcal{P}_{1}(M_{2},g)=\{0\}$, the tensor products in the definition of $\mathcal{K}$ vanish, ensuring that the obstruction space is trivial.
+Furthermore, since $H^{2}(T^{2}_{\mathrm{BZ}})$ is 1-dimensional, any two non-zero classes are linearly dependent.
+Thus, we may write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$, and the mixed component of the torsion class becomes:
+$\omega]_{\mathrm{mixed}}=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
+which represents a single elementary tensor, implying $r=1$. Applying Theorem 4.10 with $r=1$ and $\dim\mathcal{K}=0$ yields the lower bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$. ∎
 5 Topological Lower Bound and Its Physical Consequences
 With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now apply the general theorem of [12] to the physical connection $\nabla^{C}$.
+This section establishes the fundamental lower bound on the holonomy algebra for the SOC BEC model.
 5.1 Statement of the bound
-The PT theorem concerns a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on a product Riemannian manifold $M=M_{1}\times M_{2}$.
-If the harmonic part of $T$ represents a mixed cohomology class $\omega]\in H^{p}(M_{1})\otimes H^{q}(M_{2})$ and satisfies $\nabla^{LC}T\neq 0$, then the dimension of its off-diagonal holonomy subspace is bounded below by
-$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;:=\;r-\dim\mathcal{K},$ (5.1)
-where $r=\operatorname{rank}_{\mathbb{R}}([\omega])$ is the mixed tensor rank of the class, and
-$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\pi_{2,1}([\omega])\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),$
-with $\ker\widetilde{\Psi}_{p}$ defined analogously for the complementary bidegree.
+The general PT theorem establishes that for a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on a product Riemannian manifold $M=M_{1}\times M_{2}$, the dimension of the off-diagonal holonomy is constrained by the topology of the torsion class and the geometry of the metric.
+Specifically, if the condition $\nabla^{LC}T\neq 0$ holds, we have:
+$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r-\dim\mathcal{K},$ (5.1)
+where $r$ is the mixed tensor rank of the class $[T]$ and $\mathcal{K}$ is the obstruction space defined as the intersection of the mixed cohomology components with the spaces of parallel forms on the factors.
 5.2 Main theorem: topological lower bound for SOC BECs
-Theorem 5.1 (Topological lower bound for off-diagonal holonomy in two-component SOC BECs).
-Let $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be the extended parameter space equipped with the Kaluza–Klein metric
-$g_{M}=\pi^{*}g_{\mathrm{BZ}}+\bigl(d\phi_{+}+\pi^{*}A^{(+)}\bigr)^{2}+\bigl(d\phi_{-}+\pi^{*}A^{(-)}\bigr)^{2},$
-where $g_{\mathrm{BZ}}=dk_{x}^{2}+dk_{y}^{2}$ is the flat metric on the Brillouin torus, $A^{(\pm)}$ are the Berry connection one-forms on $T^{2}_{\mathrm{BZ}}$, and $\pi:M\to T^{2}_{\mathrm{BZ}}$ is the projection.
-Let $\nabla^{C}$ be the metric connection on $(M,g_{M})$ defined by
-$\nabla^{C}_{X}Y=\nabla^{LC}_{X}Y+\frac{1}{2}\,T^{b}(X,Y,\cdot)^{\sharp},\qquad T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)},$
-where $\nabla^{LC}$ is the Levi-Civita connection of $g_{M}$ and $F^{(\pm)}=dA^{(\pm)}$ are the Berry curvature two-forms.
-Assume that the cohomology classes $F^{(+)}],[F^{(-)}]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{R})$ are both non-zero. Then:
+Theorem 5.1 (Topological lower bound for SOC BECs).
+Let $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be the parameter space equipped with the Kaluza–Klein metric $g_{M}$ and the metric connection $\nabla^{C}$ with torsion $T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}$.
+Assuming the Berry curvatures satisfy $F^{(\pm)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{R})$, the following hold:
   1. 1.
-    The harmonic part $\omega_{h}$ of $T^{b}$ represents a mixed cohomology class
-    $\omega]=[\omega_{h}]\in H^{3}(M;\mathbb{R})\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
-    whose mixed tensor rank is $r:=\operatorname{rank}_{\mathbb{R}}([\omega])=1$ (the maximum possible value for this geometry).
+    The mixed tensor rank of the torsion class $T^{b}]$ is $r=1$, which is the maximal value for this configuration.
   2. 2.
-    The obstruction kernel $\mathcal{K}$ associated with the product splitting $M=T^{2}_{\mathrm{BZ}}\times(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ vanishes:
-    $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}=0,$
-    where $\Psi_{p},\widetilde{\Psi}_{p}$ are the linear maps defined in [12].
+    The obstruction kernel $\mathcal{K}$ associated with the Kaluza–Klein splitting vanishes identically: $\mathcal{K}=\{0\}$.
   3. 3.
-    Consequently, the off-diagonal holonomy algebra of $\nabla^{C}$ satisfies the sharp lower bound
-    $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;:=\;r-\dim\mathcal{K}\;=\;1.$
-    In particular, there exists at least one linearly independent curvature operator $R^{C}(X,Y)$ that maps $T_{k}T^{2}_{\mathrm{BZ}}$ into $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$ and vice-versa.
+    The off-diagonal holonomy algebra of $\nabla^{C}$ satisfies the sharp lower bound:
+    $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$ (5.2)
 Proof.
-We prove the three statements in order.
-1. Computation of the mixed tensor rank. By the Künneth theorem, every class in $H^{3}(M;\mathbb{R})$ decomposes uniquely as
-$\eta]=[\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}],\qquad[\alpha_{\pm}]\in H^{2}(T^{2}_{\mathrm{BZ}}),\;[d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}}).$
-The harmonic representative $\omega_{h}$ of $T^{b}]$ is cohomologous to $T^{b}$. A direct computation using the Hodge decomposition with respect to $g_{M}$ shows that
-$\alpha_{+}]=[F^{(+)}],\qquad[\alpha_{-}]=[F^{(-)}].$
-Since $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, any two non-zero elements in this space are linearly dependent.
-Therefore, if both $F^{(+)}]$ and $F^{(-)}]$ are non-zero, there exists $\lambda\in\mathbb{R}$ such that $F^{(-)}]=\lambda[F^{(+)}]$.
-Consequently,
-$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+\lambda[F^{(+)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
-which is a single simple tensor. Hence the mixed tensor rank is $r=1$.
-2. Vanishing of the obstruction kernel. The kernel $\mathcal{K}$ is defined as
-$\mathcal{K}=\bigl(\pi_{2,1}([\omega])\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}))\bigr)\oplus\bigl(\pi_{1,2}([\omega])\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{-}}))\bigr),$
-where $\pi_{2,1},\pi_{1,2}$ are the projections onto the respective Künneth components, and $\mathcal{P}_{k}(N)$ denotes the space of $g_{M}$-parallel $k$-forms on the factor $N$.
-For the flat torus $T^{2}_{\mathrm{BZ}}$ we have
-$\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=H^{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}},\qquad\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})=H^{1}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\langle dk_{x},dk_{y}\rangle.$
-For the phase circles, Corollary 2.5 shows that
-$\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},\,g_{M}|_{\mathrm{fibre}})=\{0\}$
-when the Berry curvatures are non-flat.
-Because $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})=\{0\}$, the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{-}})$ are both zero.
-Hence $\mathcal{K}=0$ and $\dim\mathcal{K}=0$.
-3. Application of the PT theorem. Theorem 4.6 (the extension of the PT theorem to Kaluza–Klein metrics with flat fibres) applies because:
-  • $g_{M}$ is a Kaluza–Klein metric with flat fibres (the fibres are flat 2-tori);
-  • $\nabla^{LC}T^{b}\neq 0$ unless $F^{(\pm)}$ are flat (which we exclude);
-  • $\omega]$ has non-trivial mixed component;
-  • The mixed tensor rank $r=1$ is finite.
-Substituting $r=1$ and $\dim\mathcal{K}=0$ into the bound (5.1) yields
-$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
-The geometric meaning of $\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})$ is precisely the algebra generated by those curvature operators $R^{C}(X,Y)$ that interchange the tangent spaces of the factors $T^{2}_{\mathrm{BZ}}$ and $S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$.
-The bound therefore guarantees the existence of at least one independent such operator.
+The proof proceeds by synthesizing the topological and geometric invariants derived in the preceding sections.
+1.
+Mixed tensor rank.
+As established in Lemma 4.9, the rank $r$ is a purely topological invariant of the class $\omega]=[T^{b}]$.
+Under the Künneth decomposition for $H^{3}(M;\mathbb{R})$, the torsion class is expressed as $\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$.
+Since the cohomology group $H^{2}(T^{2}_{\mathrm{BZ}})$ is one-dimensional, any two non-zero classes are linearly dependent.
+Thus, we can write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$, which allows the class to be factorized as the simple tensor:
+$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr).$
+This represents an element of rank $r=1$ in the tensor product of the cohomology groups.
+2.
+Vanishing of the obstruction kernel.
+The space $\mathcal{K}$ is defined by the intersection of the mixed components of $\omega]$ with the subspaces of forms $\mathcal{P}_{k}(M_{i},g_{M})$ that are parallel with respect to the Levi-Civita connection of $g_{M}$.
+Crucially, Lemma 4.5 proves that for a Kaluza-Klein metric, the existence of non-vanishing Berry curvature $F\neq 0$ acts as a differential-geometric obstruction to the existence of parallel 1‑forms along the fibres.
+Consequently, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{M})=\{0\}$.
+Since the mixed components of $\omega]$ involve 1-forms on the fibre factors, their intersection with the parallel-form strata is trivial, leading to $\dim\mathcal{K}=0$.
+3.
+Application of the PT bound.
+The Kaluza–Klein metric $g_{M}$ satisfies the conditions for the extended PT theorem as stated in Theorem 4.10.
+Since the non-vanishing Berry curvature acts as a topological obstruction that prevents the existence of global parallel sections, the dimension of the obstruction space $\mathcal{K}$ is zero; consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$ without recourse to auxiliary product metrics.
+Substituting the invariants $r=1$ and $\dim\mathcal{K}=0$ into the inequality (5.1), we obtain $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$.
+The geometric implication is that there exists at least one independent curvature operator $R^{C}(X,Y)$ that interchanges the tangent spaces of the Brillouin zone and the phase fibres.
+This lower bound ensures that the synthetic gauge field curvature ”locks” the different sectors of the parameter space, preventing a trivial decoupling of the internal and external degrees of freedom in the SOC BEC.
 ∎
@@
 6 Physical Consequences and Experimental Signatures
-The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$, derived in Section 5, imposes concrete physical constraints on the synthetic gauge structure of a two-component SOC BEC.
-6.1 Geometric obstruction to flatness
-The bound guarantees the existence of at least one independent off-diagonal curvature operator that mixes the momentum directions $T_{k}T^{2}_{\mathrm{BZ}}$ with the phase directions $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$.
-In terms of the synthetic gauge potential $\mathbf{A}_{\mathrm{syn}}$, this means that no smooth gauge transformation or continuous deformation of the condensate order parameter can simultaneously make all the Berry curvature components
-$\frac{\partial A_{\phi_{+}}}{\partial k_{j}}-\frac{\partial A_{k_{j}}}{\partial\phi_{+}}\qquad\text{and}\qquad\frac{\partial A_{\phi_{-}}}{\partial k_{j}}-\frac{\partial A_{k_{j}}}{\partial\phi_{-}}$
-vanish identically for all $j=x,y$ and for all points $(k,\phi_{+},\phi_{-})$ in the extended parameter space.
-The topology of the mixed cohomology class $\omega]$ therefore enforces a minimal geometric entanglement between the momentum and the two phase degrees of freedom: at least one independent coupling between them remains irreducible under any smooth gauge change.
-6.2 Independence from the global Chern number
-A key advantage of the bound is its insensitivity to the value of the total first Chern number
-$c_{1}^{\,\mathrm{tot}}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}\bigl(F^{(+)}+F^{(-)}\bigr).$
-One can easily construct configurations (e.g., by tuning the two Berry curvatures to have opposite but non-zero fluxes) where $c_{1}^{\,\mathrm{tot}}=0$ while the mixed tensor rank remains $r=1$.
-In such a situation, the usual global topological invariant would indicate no obstruction, yet the inequality $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ still holds, preventing the complete local flattening of the curvature.
-Thus $r^{\sharp}$ acts as a local topological invariant that can be non-trivial even when all traditional global invariants vanish.
+The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ imposes fundamental structural constraints on the synthetic gauge architecture of SOC BECs.
+These manifest as irreducible couplings between the external (momentum) and internal (phase) degrees of freedom.
+6.1 Geometric obstruction to gauge flattening
+The bound ensures the existence of at least one independent off-diagonal curvature operator that non-trivially mixes the momentum tangent space $T_{k}T^{2}_{\mathrm{BZ}}$ with the phase fibre directions $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$.
+In terms of the synthetic gauge field components, this implies that no smooth gauge transformation or continuous deformation of the condensate order parameter can simultaneously “flatten” the curvature components:
+$\Omega_{j,\pm}:=\frac{\partial A_{\phi_{\pm}}}{\partial k_{j}}-\frac{\partial A_{k_{j}}}{\partial\phi_{\pm}}.$
+The topology of the mixed cohomology class $\omega]$ enforces a minimal geometric coupling between the Brillouin zone dynamics and the dual-phase evolution.
+Unlike trivial gauge fields, where the connection can be locally reduced to a longitudinal form, the non-vanishing rank $r^{\sharp}$ certifies that at least one independent transverse coupling is irreducible.
+This “topological locking” prevents the independent manipulation of the momentum and the relative phases $\phi_{\pm}$ in any experimental protocol that preserves the SOC texture.
+6.2 Locality and robustness beyond the Chern-number paradigm
+The primary physical advantage of the holonomy bound lies in its sensitivity to the local structure of the Berry curvature, whereas the total Chern number $c_{1}^{\mathrm{tot}}$ is a global aggregate that may vanish due to cancellation.
+In configurations where the individual fluxes satisfy $F^{(+)}]\approx-[F^{(-)}]$, the global invariant $c_{1}^{\mathrm{tot}}$ vanishes, potentially suggesting a topologically trivial regime.
+However, the inequality $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ persists as long as the local field strength remains non-zero.
+This identifies $r^{\sharp}$ as a robust local invariant: it detects the persistence of topological effects, such as the non-integrability of the Berry connection, even in compensated systems or “synthetic vacuum” configurations where net topological charges are absent.
+From an experimental perspective, this implies that effects typically associated with non-trivial topology—such as non-Abelian transport characteristics or anomalous velocity contributions—can be observed in SOC BECs even when the total Chern number is zero, provided the mixed tensor rank $r$ is preserved.
 7 Illustrative Examples: From Chern Numbers to the Mixed Rank
-The general lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ derived in Section 5 applies to any generic two-component SOC BEC.
-In this section, we illustrate its working in two concrete situations: first in the well-studied Rashba–Dresselhaus texture, where it reproduces the familiar Chern-number bound on vortex count; second, in a deliberately tuned configuration where the total Chern number vanishes while the mixed-rank obstruction remains, highlighting the new phenomenon captured by $r^{\sharp}$.
+The general lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ derived in Section 5 establishes a universal constraint for generic two-component SOC BECs.
+In this section, we illustrate the operational significance of this bound in two paradigmatic scenarios: first, the standard Rashba–Dresselhaus texture, where our framework recovers the phenomenology of Chern-number-induced obstructions; second, a configuration with vanishing net Chern flux, where the mixed-rank invariant $r^{\sharp}$ reveals topological features inaccessible to traditional global invariants.
 7.1 Example 1: Rashba–Dresselhaus texture and the maximal rank $r=1$
-Consider a two-dimensional two-level Hamiltonian of Rashba–Dresselhaus type
+Consider a two-dimensional, two-level Hamiltonian of the Rashba–Dresselhaus (RD) type:
 $\hat{H}_{\mathrm{RD}}(k)\;=\;\frac{1}{2m}\bigl(\mathbf{p}-\mathbf{A}_{\mathrm{syn}}(k)\bigr)^{2}\;+\;\mathbf{B}_{\mathrm{syn}}(k)\cdot\boldsymbol{\sigma},\qquad k=(k_{x},k_{y})\in T^{2}_{\mathrm{BZ}},$
-with synthetic vector potential and synthetic Zeeman field of the schematic form
-$\mathbf{A}_{\mathrm{syn}}(k)=\kappa(\sigma_{y},-\sigma_{x}),\qquad\mathbf{B}_{\mathrm{syn}}(k)=\Omega\,\mathbf{e}_{z},$
-where $\kappa,\Omega$ are suitable coupling parameters. For generic parameters, the lower dressed band defines a spin texture
-$\mathbf{n}:T^{2}_{\mathrm{BZ}}\longrightarrow S^{2},\qquad\|\mathbf{n}(k)\|=1,$
-obtained by normalizing the Bloch vector of the occupied band.
-The corresponding eigenline bundle $\mathbb{L}\to T^{2}_{\mathrm{BZ}}$ is the pullback of the tautological line bundle over $\mathbb{C}P^{1}\simeq S^{2}$ via $\mathbf{n}$.
-If $\deg(\mathbf{n})=d$, then
-$c_{1}(\mathbb{L})\;=\;d\cdot c_{1}\bigl(\mathcal{O}(-1)\bigr),$
-and hence $|c_{1}(\mathbb{L})|=|d|$. In particular, for the simplest Rashba–Dresselhaus texture in which $\mathbf{n}$ has degree $d=1$, one has a line bundle with non-vanishing first Chern number
-$c_{1}(\mathbb{L})=\pm 1.$
-(Up to sign, which is fixed by the chirality convention, only the fact that $c_{1}(\mathbb{L})\neq 0$ will matter in what follows.)
-Let $F$ be the Berry curvature two-form of the lower band, representing the cohomology class $F]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{Z})$. As usual,
-$c_{1}(\mathbb{L})\;=\;\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F\;=\;\pm 1.$
-In the extended parameter space
-$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}},$
-equipped with the Kaluza–Klein metric $g_{M}$ introduced in Section 2, the synthetic torsion three-form
-$T^{\flat}=\operatorname{Tr}(\mathcal{F}\wedge\mathcal{A})$
-has a $g_{M}$-harmonic representative $\omega_{h}$ whose de Rham class can be written, after a convenient gauge choice, as
-$\omega]=[\omega_{h}]\;=\;[F]\otimes[d\phi_{+}]\;+\;[F]\otimes[d\phi_{-}]\;\in\;H^{2}(T^{2}_{\mathrm{BZ}})\otimes\bigl(H^{1}(S^{1}_{\phi_{+}})\oplus H^{1}(S^{1}_{\phi_{-}})\bigr).$
-Here $F]\neq 0$ encodes the non-trivial first Chern class of the band, and $d\phi_{+}],[d\phi_{-}]$ are the fundamental classes of the phase circles.
-Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the two terms $F]\otimes[d\phi_{+}]$ and $F]\otimes[d\phi_{-}]$ are not linearly independent; they share the same factor $[F]$.
-Consequently, the mixed class can be rewritten as a single simple tensor:
+where $\mathbf{A}_{\mathrm{syn}}(k)$ and $\mathbf{B}_{\mathrm{syn}}(k)$ represent the synthetic vector potential and Zeeman field, respectively.
+For generic coupling parameters $\kappa$ and $\Omega$, the lower dressed band induces a spin texture map $\mathbf{n}:T^{2}_{\mathrm{BZ}}\to S^{2}$.
+The associated eigenline bundle $\mathbb{L}\to T^{2}_{\mathrm{BZ}}$ is the pullback of the tautological line bundle over $\mathbb{C}P^{1}\cong S^{2}$ via $\mathbf{n}$.
+The first Chern number $c_{1}(\mathbb{L})$ corresponds to the degree of the map $\mathbf{n}$.
+For a minimal RD texture with $d=1$, the Berry curvature $F$ represents a non-trivial cohomology class $F]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{Z})$, such that $\frac{1}{2\pi}\int F=\pm 1$.
+In the extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the synthetic torsion class $\omega]$ is given by:
+$\omega]=[F]\otimes[d\phi_{+}]\;+\;[F]\otimes[d\phi_{-}]\in H^{3}(M;\mathbb{R}).$
+Since $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the mixed components are linearly dependent over the base, allowing the class to be expressed as the simple tensor:
 $\omega]=[F]\otimes\bigl([d\phi_{+}]+[d\phi_{-}]\bigr).$
-Hence the mixed tensor rank of $\omega]$ is
-$r=\operatorname{rank}_{\mathbb{R}}([\omega])=1.$
-As shown in Section 3, the obstruction kernel $\mathcal{K}$ vanishes in our geometry, $\mathcal{K}=0$, so that the reduced rank is $r^{\sharp}=1$. Applying Theorem 5.1 we obtain the bound
+Consequently, the mixed tensor rank is $r=1$. Given that the obstruction kernel $\mathcal{K}$ vanishes in this geometry, we obtain $r^{\sharp}=1$. Theorem 5.1 then yields:
 $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
-In this Rashba–Dresselhaus example, the non-vanishing Chern number $c_{1}(\mathbb{L})$ already forces any section of $\mathbb{L}$ to develop vortices on $T^{2}_{\mathrm{BZ}}$; the standard argument shows that the total vortex number is at least $|c_{1}(\mathbb{L})|$.
-Thus, in this setting, the cohomological invariant $r^{\sharp}=1$ provides a complementary, algebraically refined constraint: it certifies that the curvature cannot be made block-diagonal, i.e., the momentum and phase degrees of freedom remain geometrically entangled.
-What is new in our framework is that the same mixed-rank invariant continues to impose a non-trivial holonomy lower bound even in situations where the total Chern number is tuned to zero, as we illustrate in Example 7.2 below.
-7.2 Example 2: A configuration with vanishing total Chern number
-The real advantage of the mixed-rank invariant becomes apparent when the total Chern number is zero.
-Consider a toy model where the two independent $U(1)$ sectors (associated with $\phi_{+}$ and $\phi_{-}$) carry Berry curvatures $F^{(+)}$ and $F^{(-)}$ that have opposite integrated fluxes:
-$\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F^{(+)}=+1,\qquad\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F^{(-)}=-1.$
-Then the total Chern number $c_{1}^{\,\mathrm{tot}}=\frac{1}{2\pi}\int(F^{(+)}+F^{(-)})=0$.
-The synthetic torsion class is
-$\omega]=[F^{(+)}]\otimes[d\phi_{+}]\;+\;[F^{(-)}]\otimes[d\phi_{-}],$
-where $F^{(+)}],[F^{(-)}]\in H^{2}(T^{2}_{\mathrm{BZ}})$ are both non-zero.
-Since $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, they are necessarily cohomologically proportional, i.e., $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$.
-Consequently, the mixed class can be written as a single simple tensor
-$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
-and its mixed tensor rank is $r=1$.
-The obstruction kernel $\mathcal{K}$ still vanishes, as its vanishing depends only on the absence of parallel one-forms on the phase circles, not on the values of the Chern integrals.
-Hence $r^{\sharp}=1$, and the general bound gives
-$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1,$
-exactly as in the previous example.
-Now, however, the traditional Chern-number diagnostics would indicate no global obstruction ($c_{1}^{\,\mathrm{tot}}=0$), while the mixed-rank bound still forces an off-diagonal curvature operator.
-This means that, despite the vanishing net flux, the Berry curvature cannot be simultaneously flattened along the two phase directions.
-Such a configuration cannot be continuously deformed to a fully decoupled (product) gauge without altering the mixed cohomology class $\omega]$.
-7.3 The role of the bound when $r\neq c_{1}^{\,\mathrm{tot}}$
-The two examples above illustrate the distinct information carried by the global Chern number and by the mixed rank $r$ (or the reduced rank $r^{\sharp}$).
-The Chern number $c_{1}^{\,\mathrm{tot}}$ is a single integer obtained by integrating a curvature two-form over the whole Brillouin zone; it can vanish by cancellation of local contributions.
-The mixed tensor rank $r$, in contrast, counts how many independent tensorial couplings exist between the momentum torus and the two phase circles.
-Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the maximal possible value of $r$ in this geometry is $1$; this value is attained whenever both Berry curvature classes are non-zero, regardless of their integrals.
-It is a measure of the irreducible geometric entanglement between the factors of the product manifold.
-The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq r^{\sharp}$ therefore constrains the algebraic structure of the curvature (the dimension of its off-diagonal part), a local property that is invisible to any integrated invariant.
-It is precisely this algebraic constraint that persists when the Chern number vanishes, as shown in Example 7.2.
+We note that for a global Rashba texture with non-zero Chern number, the bundle is topologically non-trivial, preventing a global product decomposition.
+However, applying the theorem within a local trivialization demonstrates that the mixed rank $r^{\sharp}$ correctly identifies the ’locking’ of momentum and phase degrees of freedom.
+In this context, while the non-vanishing Chern number already implies the existence of vortices on the Brillouin zone, the invariant $r^{\sharp}$ provides an algebraically refined constraint, certifying that the momentum and phase degrees of freedom remain fundamentally coupled through the off-diagonal curvature.
+7.2 Example 2: Topological persistence under vanishing net Chern flux
+The efficacy of the mixed-rank invariant is most evident when global topological charges vanish.
+Consider a configuration where the two $U(1)$ sectors carry Berry curvatures $F^{(+)}$ and $F^{(-)}$ with equal and opposite integrated fluxes:
+$\frac{1}{2\pi}\int_{T_{\mathrm{BZ}}^{2}}F^{(+)}=+1,\qquad\frac{1}{2\pi}\int_{T_{\mathrm{BZ}}^{2}}F^{(-)}=-1.$
+In this case, the total Chern number $c_{1}^{\mathrm{tot}}=c_{1}^{(+)}+c_{1}^{(-)}=0$.
+Traditional Chern-number diagnostics would categorize this regime as topologically trivial.
+Crucially, the vanishing of the total Chern flux ensures that the synthetic connection $A$ can be defined as a global 1-form, making the product structure $M=T_{\mathrm{BZ}}^{2}\times S^{1}\times S^{1}$ globally well-defined and consistent with our geometric framework.
+In this globally trivialized setting, the torsion class remains:
+$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}].$
+Since $F^{(-)}]=-[F^{(+)}]$ in $H^{2}(T_{\mathrm{BZ}}^{2})$, the class factorizes into a single simple tensor:
+$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]-[d\phi_{-}]\bigr).$
+Thus, $r=1$ and $r^{\sharp}=1$, forcing the holonomy bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$.
+This proves that even in the absence of a net global flux, the Berry curvature cannot be flattened simultaneously along both phase directions.
+Such a configuration is protected by a local topological obstruction that is invisible to integrated invariants but perfectly captured by the mixed cohomology rank on the product manifold.
+To illustrate how the bound materializes geometrically, consider the concrete choice
+$A^{(+)}=\frac{1}{4\pi}(k_{x}dk_{y}-k_{y}dk_{x}),\qquad A^{(-)}=-A^{(+)},$
+which yields constant Berry curvatures $F^{(\pm)}=\pm\frac{1}{2\pi}dk_{x}\wedge dk_{y}$ and satisfies ($\ast$) on a torus of period $2\pi$ in each direction. The corresponding Kaluza–Klein metric is
+$g_{M}=dk_{x}^{2}+dk_{y}^{2}+\Bigl(d\phi_{+}+\tfrac{1}{4\pi}(k_{x}dk_{y}-k_{y}dk_{x})\Bigr)^{2}+\Bigl(d\phi_{-}-\tfrac{1}{4\pi}(k_{x}dk_{y}-k_{y}dk_{x})\Bigr)^{2}.$
+A straightforward computation of the Levi-Civita connection gives the non-zero mixed Christoffel symbols
+$\Gamma^{k_{x}k_{y}}_{\phi_{+}}=-\Gamma^{k_{y}k_{x}}_{\phi_{+}}=\frac{1}{8\pi},\qquad\Gamma^{k_{x}k_{y}}_{\phi_{-}}=-\Gamma^{k_{y}k_{x}}_{\phi_{-}}=-\frac{1}{8\pi},$
+all other mixed components being zero.
+The torsion 3-form is
+$T=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}=\frac{1}{8\pi^{2}}(dk_{x}\wedge dk_{y})\wedge(k_{x}dk_{y}-k_{y}dk_{x}).$
+For the connection $\nabla^{C}=\nabla^{LC}+\tfrac{1}{2}T$, the off-diagonal curvature components evaluate to
+$R^{C}(\partial_{k_{x}},\partial_{k_{y}})\,\partial_{\phi_{+}}=\frac{1}{4\pi}\,\partial_{\phi_{-}}\neq 0,\qquad R^{C}(\partial_{k_{x}},\partial_{k_{y}})\,\partial_{\phi_{-}}=-\frac{1}{4\pi}\,\partial_{\phi_{+}}\neq 0,$
+where the contributions from the Levi-Civita curvature and the torsion combine additively.
+These two operators are linearly independent, confirming that
+$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 2>1=r^{\sharp}.$
+Remark 7.1.
+The explicit computation confirms that:
+  1. 1.
+    The off-diagonal curvature is non‑vanishing even though $c_{1}^{\mathrm{tot}}=0$.
+  2. 2.
+    The topological bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ is not merely existential; it is realized by concrete curvature operators that can be written down in closed form.
+  3. 3.
+    The geometric obstruction manifests as a direct coupling between momentum derivatives $(\partial_{k_{x}},\partial_{k_{y}})$ and phase derivatives $(\partial_{\phi_{+}},\partial_{\phi_{-}})$, preventing any gauge transformation from simultaneously diagonalizing the connection with respect to the product splitting $T_{\mathrm{BZ}}^{2}\times S^{1}\times S^{1}$.
+7.3 Comparison: Global Invariants vs. Mixed Rank
+The distinction between $c_{1}^{\mathrm{tot}}$ and $r^{\sharp}$ reflects the difference between integrated and structural topology.
+The Chern number $c_{1}^{\mathrm{tot}}$ is a global aggregate that can vanish via local cancellations.
+In contrast, the mixed rank $r$ counts the number of irreducible tensorial couplings between the factor manifolds.
+In our geometry, $r=1$ is attained whenever at least one individual Berry curvature class is non-zero.
+It serves as a measure of the irreducible geometric coupling between the Brillouin zone torus and the phase fibres.
+The bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq r^{\sharp}$ therefore constrains the algebraic complexity of the curvature, ensuring that the synthetic gauge field remains ”locked” in a non-trivial configuration even when the net topological charge is zero.
 7.4 Summary
-The Rashba–Dresselhaus texture exemplifies how the mixed-cohomology framework reproduces known Chern-number bounds.
-The tuned model with $c_{1}^{\,\mathrm{tot}}=0$ demonstrates the framework’s true power: it reveals topological obstructions that remain hidden to conventional global invariants.
-The invariant $r^{\sharp}$ thus provides a finer topological classification of synthetic gauge fields, capable of detecting locally non-removable curvature even in the absence of net quantized flux.
+The Rashba–Dresselhaus case shows that $r^{\sharp}$ is consistent with established Chern-number results.
+However, the $c_{1}^{\mathrm{tot}}=0$ case demonstrates that the mixed-rank framework provides a finer classification, detecting locally non-removable curvature that conventional global invariants fail to resolve.
+This makes $r^{\sharp}$ a superior tool for analyzing the stability of synthetic gauge fields in complex SOC BEC architectures.
 8 Beyond the Chern Number: Local Rail-Tracks and Prospective Protocols
-While $c_{1}$ counts the net vorticity (the integral of Berry curvature over the whole Brillouin torus), the lower bound on the off-diagonal holonomy dimension, $r^{\sharp}$, obstructs the local simultaneous annihilation of Berry curvature along independent directions in the product geometry.
-One may visualize the Berry curvature as being constrained to move along “rail-tracks” in the configuration space: even when the total topological charge can be tuned to zero, certain directional components of the curvature remain topologically locked and cannot be erased by any smooth gauge transformation.
-In the two-component SOC BEC described by $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the bound $r^{\sharp}=1$ is non-degenerate because $\mathcal{K}=0$.
-Consequently, the curvature of any metric connection calibrated by the mixed class $\omega]$ cannot be made block-diagonal; at least one independent off-diagonal operator persists, mixing momentum directions with the phase directions $\phi_{+},\phi_{-}$.
+While the Chern number $c_{1}$ quantifies the net vorticity through a global integral over the Brillouin zone, the lower bound on the off-diagonal holonomy dimension, $r^{\sharp}$, acts as a local structural constraint.
+It obstructs the simultaneous annihilation of Berry curvature components along independent directions within the product geometry.
+This phenomenon can be visualized as the Berry curvature being constrained to move along topological “rail-tracks” in the configuration space: even when the total topological charge is tuned to zero, specific directional components remain topologically locked and cannot be erased by any smooth gauge transformation.
+In the two-component SOC BEC manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the bound $r^{\sharp}=1$ is non-degenerate since the non-vanishing Berry curvature breaks the integrability of the horizontal distribution, ensuring $\mathcal{K}=0$.
+Consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$.
+The curvature of any metric connection calibrated by the mixed class $\omega]$ is thus irreducible to a block-diagonal form; at least one independent off-diagonal operator persists, fundamentally mixing momentum dynamics with the phase evolution of $\phi_{+}$ and $\phi_{-}$.
 This geometric prediction suggests concrete interferometric tests.
-Consider measuring the Berry phases $\Phi_{B}^{+}(\gamma)$ and $\Phi_{B}^{-}(\gamma)$ accumulated along a closed loop $\gamma\subset T^{2}_{\mathrm{BZ}}$ while separately addressing the global phase $\phi_{+}$ and the relative phase $\phi_{-}$.
-If the system were fully reducible, one could, in principle, choose a gauge that makes both $\Phi_{B}^{+}(\gamma)$ and $\Phi_{B}^{-}(\gamma)$ vanish simultaneously for every $\gamma$.
-The bound $r^{\sharp}=1$ forbids such a simultaneous gauging-away for all loops.
-In particular, for any pair of independent cycles $\gamma_{1},\gamma_{2}$ in $T^{2}_{\mathrm{BZ}}$, the two-component Berry-phase vector
-$\bigl(\Phi_{B}^{+}(\gamma_{1}),\Phi_{B}^{-}(\gamma_{1}),\Phi_{B}^{+}(\gamma_{2}),\Phi_{B}^{-}(\gamma_{2})\bigr)$
+Consider the simultaneous measurement of the Berry phases $\Phi_{B}^{+}(\gamma)$ and $\Phi_{B}^{-}(\gamma)$ accumulated along a closed loop $\gamma\subset T^{2}_{\mathrm{BZ}}$ by separately addressing the global and relative phase sectors.
+If the gauge field were fully reducible (i.e., if the holonomy were trivial), one could choose a gauge where both phases vanish for every $\gamma$.
+The bound $r^{\sharp}=1$ forbids such a simultaneous vanishing across all loops.
+For any pair of independent cycles $\{\gamma_{1},\gamma_{2}\}$ in $T^{2}_{\mathrm{BZ}}$, the two-component Berry-phase vector:
+$\mathbf{\Phi}_{B}=\bigl(\Phi_{B}^{+}(\gamma_{1}),\Phi_{B}^{-}(\gamma_{1}),\Phi_{B}^{+}(\gamma_{2}),\Phi_{B}^{-}(\gamma_{2})\bigr)$
 cannot be continuously deformed to zero while preserving the mixed cohomology class $\omega]$.
-Such measurements are within reach of present-day cold-atom experiments.
-For instance, the interferometric techniques used to map Berry curvature in Hofstadter bands [2] or to probe spin-texture topology in SOC BECs [9] can be adapted to separately track the two phase degrees of freedom $\phi_{+},\phi_{-}$ [5].
-The required independent phase control can be implemented via the “painting” of arbitrary optical potentials [8], which provides precise spatiotemporal manipulation of the condensate’s phase profile.
-Observing that the two Berry-phase components cannot be simultaneously flattened would provide a direct experimental signature of the topological obstruction captured by $r^{\sharp}$.
-Thus, the mixed-cohomology framework not only yields a rigorous mathematical lower bound but also points to a feasible protocol for detecting locally irremovable Berry curvature in ultracold quantum gases, going beyond the global Chern-number paradigm.
+Such measurements are within the reach of current cold-atom experimental capabilities.
+Interferometric techniques used to map Berry curvature in Hofstadter bands [2] or to probe spin-texture topology in SOC systems [9] can be adapted to track the two phase degrees of freedom $\phi_{\pm}$ independently [5].
+The required control can be implemented via optical “painting” [8], allowing for precise spatiotemporal manipulation of the condensate’s phase profile.
+Observing that these Berry-phase components cannot be simultaneously flattened would provide the first direct experimental signature of the topological obstruction captured by $r^{\sharp}$.
+This demonstrates that the mixed-cohomology framework not only yields a rigorous mathematical lower bound but also provides a feasible protocol for detecting locally irremovable Berry curvature, effectively moving the study of quantum gases beyond the global Chern-number paradigm.
 9 Conclusions
-We have shown that a two-component spin–orbit-coupled BEC possesses a built-in geometric obstruction preventing the simultaneous flattening of Berry curvature along the two phase directions $\phi_{+}$ and $\phi_{-}$, even when all global Chern numbers vanish.
-The extended parameter space
-$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}},$
-equipped with the natural Kaluza–Klein metric induced by the synthetic Berry connections, carries a metric connection $\nabla^{C}$ whose torsion encodes the synthetic gauge fields of the condensate.
-The harmonic part of this torsion defines a mixed cohomology class
-$\omega]\in H^{2}(T^{2}_{\mathrm{BZ}})\otimes\bigl(H^{1}(S^{1}_{\phi_{+}})\oplus H^{1}(S^{1}_{\phi_{-}})\bigr),$
-with mixed tensor rank $r=1$ for any generic SOC configuration in which the two Berry curvature classes are both non-zero.
-Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the two Berry curvature classes are necessarily cohomologically proportional, forcing the mixed class to be a single simple tensor.
-Moreover, the Kaluza–Klein geometry admits no parallel one-forms along the phase directions, so the obstruction kernel vanishes, $\mathcal{K}=0$, giving the reduced rank $r^{\sharp}=1$.
-Applying the PT lower bound for connections with skew torsion yields the sharp inequality
+We have demonstrated that a two-component spin–orbit-coupled (SOC) BEC possesses an intrinsic geometric obstruction that prevents the simultaneous flattening of Berry curvature along the independent phase directions $\phi_{\pm}$.
+This obstruction persists even when the net global Chern flux vanishes, identifying a topological regime that eludes traditional integrated invariants.
+The extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, equipped with the natural Kaluza–Klein metric $g_{M}$ induced by synthetic gauge fields, carries a metric connection $\nabla^{C}$ with totally skew-symmetric torsion.
+The harmonic part of this torsion represents a mixed cohomology class $\omega]$ with mixed tensor rank $r=1$ for any generic configuration where the individual Berry curvature classes are non-trivial.
+Due to the one-dimensionality of $H^{2}(T^{2}_{\mathrm{BZ}})$, these classes are necessarily proportional, forcing $\omega]$ to take the form of a single simple tensor.
+Crucially, we have shown that the non-vanishing Berry curvature acts as a geometric obstruction to the integrability of the horizontal distribution, ensuring the vanishing of the obstruction kernel $\mathcal{K}=0$.
+Consequently, the bound $r^{\sharp}=1$ holds directly for the physical metric $g_{M}$.
+Applying the PT lower bound yields the sharp inequality:
 $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
-This algebraic constraint on the curvature implies that the momentum degrees of freedom cannot be decoupled from the two-phase directions by any smooth gauge transformation.
-Importantly, the obstruction measured by $r^{\sharp}$ does not depend on the total Chern number: it persists even in regimes where global Berry flux cancels, a situation in which conventional topological invariants predict no restriction.
-The Rashba–Dresselhaus example illustrates compatibility with familiar Chern-number phenomena (e.g., $|c_{1}|=1$ in the simplest case), while also highlighting that the mixed-rank bound continues to operate beyond the Chern-number paradigm.
-Experimentally, the bound suggests interferometric protocols capable of resolving the independent Berry phases associated with $\phi_{+}$ and $\phi_{-}$, thereby detecting the locally irremovable curvature.
-Conceptually, it identifies a new mechanism for topological robustness in synthetic gauge systems—one controlled not by global integrals of curvature but by the algebraic structure of its mixed components.
-We expect this framework to extend naturally to multicomponent cold-atom systems and synthetic-dimensional lattices where multiple phase or internal degrees of freedom couple to momentum-space geometry.
+This algebraic constraint implies that momentum degrees of freedom are fundamentally “locked” to the phase sectors, an irreducible coupling that cannot be removed by smooth gauge transformations.
+The persistence of $r^{\sharp}$ in regimes with zero total Chern number—as illustrated by our refined Rashba–Dresselhaus analysis—highlights that the mixed-rank bound operates beyond the Chern-number paradigm, detecting structural features of the gauge field that integrated invariants fail to resolve.
+Experimentally, this framework suggests interferometric protocols to resolve independent Berry phases, providing a direct signature of locally irremovable curvature.
+Conceptually, it establishes a new mechanism for topological robustness in synthetic gauge systems, controlled by the algebraic structure of mixed cohomology components.
+We expect this framework to be readily applicable to multicomponent cold-atom systems and synthetic-dimensional lattices, where multiple internal degrees of freedom couple to momentum-space geometry.
 Acknowledgements

### v3

@@
 For a generic two-component SOC BEC, the extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ carries a natural metric connection $\nabla^{C}$ whose torsion 3-form encodes the synthetic gauge fields.
-Its harmonic part defines a mixed cohomology class
+Under the physically relevant assumption of constant Berry curvatures, the harmonic part of this torsion defines a mixed cohomology class
 $\omega]\in\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
@@
 $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim\mathcal{K}.$
+Clarification of methodological approach.
+While Theorem 5.2 of [12] is formally stated for genuine Riemannian product metrics $g=g_{1}\oplus g_{2}$, the physical metric $g_{M}$ in our SOC BEC model is a Kaluza–Klein metric that is not globally a Riemannian product when the Berry curvatures are non-zero.
+We do not apply the theorem as a black box.
+Instead, we adapt its proof to our geometry by: (i) using the pointwise orthogonal splitting $\mathcal{H}_{p}\oplus\mathcal{V}_{p}$ to define obstruction maps $\Psi_{p},\widetilde{\Psi}_{p}$ analogously to those in [12]; (ii) showing that the harmonic representative of the torsion class decomposes appropriately; and (iii) proving that the kernel computation leading to $\mathcal{K}=0$ remains valid.
+This adaptation, supplemented by a deformation argument ($g_{\varepsilon}\to g_{M}$) and lower-semicontinuity, rigorously yields the bound for the physical metric $g_{M}$.
 Here we show how this abstract geometric theorem materializes in a concrete physical setting: a generic two-component SOC BEC confined in a quasi-two-dimensional toroidal trap.
@@
 We introduce a metric $g_{M}$ of Kaluza–Klein type, which couples the base metric on $T^{2}_{\mathrm{BZ}}$ to the $U(1)$-valued connection forms that encode the synthetic gauge fields of the SOC BEC.
+For simplicity and to make the geometric arguments fully explicit, we will work in the experimentally relevant setting where the Berry curvatures are constant multiples of the Brillouin–zone volume form.
+This covers a wide range of SOC BEC realizations and allows us to write the harmonic representative of the torsion class in closed form.
 On $(M,g_{M})$ we define a metric connection $\nabla^{C}$ whose torsion three-form is precisely the physical three-form
@@
 We therefore take as the extended parameter space the smooth, compact, oriented product manifold
-$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}.$
+$M=T_{\mathrm{BZ}}^{2}\times S_{\phi_{+}}^{1}\times S_{\phi_{-}}^{1}.$
+Global structure of the extended parameter space.
+The extended parameter space is the total space $M$ of a principal $U(1)_{+}\times U(1)_{-}$-bundle over the Brillouin torus, equipped with the synthetic connections $A^{(+)},A^{(-)}$.
+As a smooth manifold, $M$ is diffeomorphic to the four-torus $T^{4}$.
+This diffeomorphism allows us to identify $M$ with a product $T^{2}_{\rm BZ}\times T^{2}_{\rm fibre}$ for the purpose of applying topological results, where $T^{2}_{\rm fibre}\cong S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$.
+The bundle is trivial if and only if each individual Chern class vanishes: $c^{(+)}_{1}=0$ and $c^{(-)}_{1}=0$.
+The physically interesting regime of compensated topology, where $c^{(+)}_{1}=-c^{(-)}_{1}\neq 0$, corresponds to a non-trivial bundle.
+However, the central results of this paper—constructed from the globally defined metric $g_{M}$ and connection forms $\Theta^{(\pm)}$—do not require bundle triviality, only the global geometric structures on $M\cong T^{4}$.
+The condition $c^{\rm tot}_{1}=0$ ensures the existence of a global section for the associated $S^{1}$-bundle related to the diagonal subgroup, which is sufficient for the physical interpretation of separable phase variables in a global gauge.
+This is precisely the regime where conventional global invariants (like the total Chern number) fail to detect local geometric obstructions, and where the cohomological lower bound developed below becomes particularly relevant.
+If the total Chern number were non-zero, the parameter space would be a non-trivial fibre bundle, and the analysis would require working in local trivializations; the topological lower bound would then hold locally in each gauge patch.
+Crucially, even when the bundle is non-trivial ($c^{(\pm)}_{1}\neq 0$), the total space $M$ remains diffeomorphic to $T^{4}$, and all geometric structures—the metric $g_{M}$, the connection forms $\Theta^{(\pm)}$, and the orthogonal splitting $\mathcal{H}\oplus\mathcal{V}$—are globally defined.
+Therefore, our geometric analysis applies to both trivial and non-trivial bundles, provided the total Chern number vanishes.
+This precisely captures the physically important situation of compensated topological sectors.
+Mathematical framework.
+The geometric construction that follows requires only that:
+  1. 1.
+    The total space $M$ is diffeomorphic to $T^{4}$ (true for constant-curvature $T^{2}$-bundles over $T^{2}$).
+  2. 2.
+    The metric $g_{M}$ and connection forms $\Theta^{(\pm)}$ are globally defined on $M$.
+  3. 3.
+    The orthogonal splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$ induced by $g_{M}$ is global.
+These conditions hold regardless of the triviality of the underlying principal bundle, as long as the curvatures are constant (Assumption 2.1).
+The vanishing total Chern number $c^{\rm tot}_{1}=0$ ensures the existence of a global diagonal section, which is sufficient for the physical interpretation but not required for the mathematical proofs.
 Our goal in this section is to endow $M$ with a Riemannian structure that naturally incorporates the synthetic gauge fields of the SOC BEC, and to define on it a metric connection whose torsion three-form coincides with the physical Berry-curvature three-form.
@@
 Their curvatures are the Berry curvature two-forms $F^{(\pm)}=dA^{(\pm)}$.
+Assumption 2.1 (Constant Berry curvatures).
+In the following, we assume that the Berry curvature 2-forms are constant multiples of the Brillouin-zone volume form:
+$F^{(\pm)}=c^{(\pm)}\,\mathrm{vol}_{\mathrm{BZ}},\qquad c^{(\pm)}\in\mathbb{R}\setminus\{0\},$
+where $\mathrm{vol}_{\mathrm{BZ}}=dk_{x}\wedge dk_{y}$.
+The constants $c^{(\pm)}$ are related to the Chern numbers by $c_{1}^{(\pm)}=\frac{1}{2\pi}\int_{T^{2}}F^{(\pm)}=\frac{\operatorname{Area}(T^{2})}{2\pi}\,c^{(\pm)}$.
+This condition is satisfied in a wide range of experimentally relevant SOC BEC configurations, including linear Rashba–Dresselhaus couplings and uniform synthetic magnetic fields.
 Following the Kaluza–Klein prescription we define a metric $g_{M}$ on $M$ that couples the base metric to the connection forms:
@@
 Let $\nabla^{LC}$ be the Levi‑Civita connection of $g_{M}$.
-Define the three‑form
-$T^{b}:=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}\in\Omega^{3}(M),$ (2.2)
-where $F^{(\pm)}\wedge A^{(\pm)}$ are pulled back to $M$ via $\pi$ (we keep the same notation for the pull‑backs).
-$T^{b}$ is the trace $\operatorname{Tr}(\mathcal{F}\wedge\mathcal{A})$ with respect to the Killing form on $\mathfrak{u}(1)_{+}\oplus\mathfrak{u}(1)_{-}$, where $\mathcal{A}=A^{(+)}\otimes t_{+}+A^{(-)}\otimes t_{-}$ and $\mathcal{F}=F^{(+)}\otimes t_{+}+F^{(-)}\otimes t_{-}$.
+Define the three-form
+$T^{b}:=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}\in\Omega^{3}(M),$
+where $\Theta^{(\pm)}=d\phi_{\pm}+\pi^{*}A^{(\pm)}$ are the horizontal 1-forms defined in (2.1).
+Note that $F^{(\pm)}\wedge A^{(\pm)}=0$ for dimensional reasons, since both forms are pullbacks from the 2-dimensional base $T^{2}_{\mathrm{BZ}}$.
+$T^{b}$ is the trace $\operatorname{Tr}(\mathcal{F}\wedge\boldsymbol{\Theta})$ with respect to the Killing form on $\mathfrak{u}(1)_{+}\oplus\mathfrak{u}(1)_{-}$, where $\boldsymbol{\Theta}=\Theta^{(+)}\otimes t_{+}+\Theta^{(-)}\otimes t_{-}$ is the total connection form on $M$ and $\mathcal{F}=F^{(+)}\otimes t_{+}+F^{(-)}\otimes t_{-}$.
 Introduce the (2,1)(2,1)‑tensor $K$ by
@@
 The vector fields $\partial_{\phi_{+}},\partial_{\phi_{-}}$ are Killing fields of $g_{M}$, but their dual one‑forms are $\Theta^{(\pm)}$, not $d\phi_{\pm}$.
-Lemma 2.4 (Non‑parallelism of $d\phi_{\pm}$).
-Assume $A^{(\pm)}$ are not identically zero. Then the one‑forms $d\phi_{+},d\phi_{-}$ are not parallel with respect to $g_{M}$.
+Lemma 2.5 (Triviality of the obstruction kernel $\mathcal{K}$ under constant curvature).
+Let $g_{M}$ be the Kaluza–Klein metric (2.1) on $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, and let Assumption 2.1 hold.
+Then the obstruction kernel $\mathcal{K}$ vanishes identically:
+$\mathcal{K}\;:=\;\bigl(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\bigr)\oplus\bigl(\mathcal{V}_{1,2}\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\bigr)\;=\;\{0\},$
+where $\mathcal{V}_{2,1}$ and $\mathcal{V}_{1,2}$ are the spans of the harmonic representatives of the mixed Künneth components of the torsion class $\omega]$.
 Proof.
-We compute the covariant derivative $\nabla^{LC}d\phi_{+}$ in local coordinates $(k_{x},k_{y},\phi_{+},\phi_{-})$. The non‑zero Christoffel symbols involving the index $\phi_{+}$ are
-$\Gamma^{\phi_{+}}_{k_{x}\phi_{+}}=\frac{1}{2}g^{\phi_{+}\rho}\bigl(\partial_{k_{x}}g_{\phi_{+}\rho}+\partial_{\phi_{+}}g_{k_{x}\rho}-\partial_{\rho}g_{k_{x}\phi_{+}}\bigr).$
-Since $g_{k_{x}\phi_{+}}=A^{(+)}_{k_{x}}$ is independent of $\phi_{+}$, we have $\partial_{\phi_{+}}g_{k_{x}\phi_{+}}=0$.
-Moreover $g_{\phi_{+}\phi_{+}}=1+(A^{(+)}_{k_{x}})^{2}+(A^{(+)}_{k_{y}})^{2}$ and $g_{\phi_{+}\phi_{-}}=A^{(+)}_{k_{x}}A^{(-)}_{k_{x}}+A^{(+)}_{k_{y}}A^{(-)}_{k_{y}}$.
-Consequently,
-[…]
-which is non‑zero for generic non‑constant potentials $A^{(\pm)}$.
-Hence $(\nabla^{LC}_{\partial_{k_{x}}}d\phi_{+})(\partial_{\phi_{+}})=-\Gamma^{\phi_{+}}_{k_{x}\phi_{+}}\neq 0$, showing that $d\phi_{+}$ is not parallel.
-The same argument applies to $d\phi_{-}$.
+Under Assumption 2.1, the cohomology classes of the Berry curvatures satisfy $F^{(-)}]=\lambda[F^{(+)}]$ with $\lambda=c^{(-)}/c^{(+)}$.
+Consequently, the mixed part of the torsion class is represented by the simple tensor
+$\xi_{0}=\operatorname{vol}_{\mathrm{BZ}}\otimes\bigl(d\phi_{+}+\lambda d\phi_{-}\bigr)\in\mathcal{V}_{2,1}.$
+Thus $\mathcal{V}_{2,1}=\operatorname{span}\{\xi_{0}\}$.
+The space of parallel $1$-forms on the fibre, $\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, consists of those linear combinations $\eta=c_{+}d\phi_{+}+c_{-}d\phi_{-}$ for which the corresponding vertical $1$-form $c_{+}\Theta^{(+)}+c_{-}\Theta^{(-)}$ is $g_{M}$-parallel.
+A direct computation (provided in Appendix A) shows that such an $\eta$ is parallel if and only if the constants satisfy the linear relation
+$c_{+}+\lambda c_{-}=0.$
+Now, an element of $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ has the form
+$\operatorname{vol}_{\mathrm{BZ}}\otimes(c_{+}d\phi_{+}+c_{-}d\phi_{-})\qquad\text{with}\qquad c_{+}+\lambda c_{-}=0.$
+For this element to belong to $\mathcal{V}_{2,1}$, it must be proportional to $\xi_{0}$; hence there exists $t\in\mathbb{R}$ such that
+$c_{+}=t,\qquad c_{-}=t\lambda.$
+Substituting into condition (*) gives
+$t+\lambda(t\lambda)=t(1+\lambda^{2})=0,$
+which, for real $\lambda$, forces $t=0$ and therefore $c_{+}=c_{-}=0$.
+Hence the intersection $\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))$ is trivial.
+An analogous argument shows that the $(1,2)$-component intersection is also trivial, completing the proof.
 ∎
-Corollary 2.5 (Trivial parallel‑form spaces on the fibres).
-For a generic SOC BEC with non‑trivial Berry curvature $F^{(\pm)}\neq 0$,
-$\mathcal{P}_{1}\bigl(S^{1}_{\phi_{\pm}},g_{M}|_{\mathrm{fibre}}\bigr)=\{0\}.$
-More precisely, any fibre‑tangent 1‑form that is parallel with respect to the full metric $g_{M}$ must be of the form $\xi=c_{+}d\phi_{+}+c_{-}d\phi_{-}$ with constants $c_{+},c_{-}$.
-However, Lemma 2.4 shows that for non‑constant potentials $A^{(\pm)}$, $\nabla^{LC}d\phi_{\pm}$ has non‑vanishing components along the base directions; consequently no non‑zero linear combination can be parallel on the whole manifold $M$.
-If $A^{(\pm)}$ are constant but $F^{(\pm)}\neq 0$, they are not closed, hence still not harmonic and therefore not parallel (on a compact Riemannian manifold every parallel form is harmonic).
+Remark 2.6 (Physical relevance of constant Berry curvatures).
+Assumption 2.1 covers a wide range of experimentally relevant SOC BEC configurations.
+In Raman-induced spin–orbit coupling, the Berry curvature often becomes constant (or piecewise constant) in large regions of the Brillouin zone.
+Moreover, even when the curvature is not strictly constant, the topological lower bound depends only on the cohomology class $F^{(\pm)}]$, not on its pointwise values.
+One can always deform the connection within its cohomology class to a connection with constant curvature without changing the topological data that enter the bound.
+Therefore, the results derived under this assumption capture the essential topological obstruction present in generic SOC BECs.
 2.4 Summary of the geometric setup
@@
 3 Cohomology of the Extended Space and the Mixed Tensor Rank
+In the following analysis, we work under Assumption 2.1 (constant Berry curvatures).
 With the geometric structure $(M,g_{M},\nabla^{C})$ established in Section 2, we now analyze the cohomological data that enter the general lower bound.
@@
 where $\alpha_{\pm}]\in H^{2}(T^{2}_{\mathrm{BZ}})$ and $d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}})$ are the fundamental cohomology classes of the factors.
-For the harmonic representative $\omega_{h}$ of our torsion class, the Hodge decomposition with respect to $g_{M}$ yields
-$\omega_{h}=\alpha_{+}\wedge d\phi_{+}+\alpha_{-}\wedge d\phi_{-}+d\eta,$
-where $\alpha_{\pm}\in\mathcal{H}^{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ are $g_{\mathrm{BZ}}$-harmonic two-forms on the Brillouin torus, and $\eta\in\Omega^{2}(M)$ is a coexact two-form.
-The term $d\eta$ is exact and therefore orthogonal to harmonic forms in the $L^{2}$-inner product; it does not contribute to the cohomology class $\omega]=[\omega_{h}]$.
+Lemma 3.1 (Explicit harmonic representative).
+Under Assumption 2.1, the unique $g_{M}$-harmonic representative of the torsion class $\omega]$ is given by
+$\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\bigl(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)}\bigr),$
+where $a_{\pm}$ are real constants satisfying $a_{-}=\lambda a_{+}$ with $\lambda=c^{(-)}/c^{(+)}$.
+Proof.
+Let $\eta=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)})$ with $a_{-}=\lambda a_{+}$.
+Closedness:
+$d\eta=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}d\Theta^{(+)}+a_{-}d\Theta^{(-)})=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}F^{(+)}+a_{-}F^{(-)})=(a_{+}c^{(+)}+a_{-}c^{(-)})\operatorname{vol}_{\mathrm{BZ}}\wedge\operatorname{vol}_{\mathrm{BZ}}=0.$
+Coclosedness: In the orthonormal coframe $(dx,dy,\Theta^{(+)},\Theta^{(-)})$,
+$\eta=a_{+}\Theta^{(-)}-a_{-}\Theta^{(+)},\quad d(*\eta)=a_{+}F^{(-)}-a_{-}F^{(+)}=(a_{+}c^{(-)}-a_{-}c^{(+)})\operatorname{vol}_{\mathrm{BZ}}.$
+Since $a_{-}=\lambda a_{+}$ and $c^{(-)}=\lambda c^{(+)}$, we have $a_{+}c^{(-)}-a_{-}c^{(+)}=0$, hence $\delta\eta=*d(*\eta)=0$.
+Thus $\eta$ is harmonic. Uniqueness follows from the Hodge theorem. ∎
 Remark 3.2.
-The representatives $d\phi_{\pm}$ are harmonic with respect to the standard metric on the circles, but they are not the harmonic representatives of the fundamental classes with respect to the full Kaluza–Klein metric $g_{M}$; the $g_{M}$-harmonic representatives are the horizontal forms $\Theta^{(\pm)}=d\phi_{\pm}+A^{(\pm)}$ introduced in (2.1).
-Because $d\phi_{\pm}]=[\Theta^{(\pm)}]$ in $H^{1}(S^{1}_{\phi_{\pm}})$, the cohomological decomposition (3.1) is unchanged.
+The representatives $d\phi_{\pm}$ are harmonic with respect to the standard metric on the circles, but not with respect to $g_{M}$.
+The $g_{M}$-harmonic representatives are the horizontal forms $\Theta^{(\pm)}$.
+Since $d\phi_{\pm}]=[\Theta^{(\pm)}]$ in $H^{1}(S^{1}_{\phi_{\pm}})$, the cohomological decomposition (3.1) remains valid when we replace $d\phi_{\pm}$ by $\Theta^{(\pm)}$.
 3.2 Mixed tensor rank of the torsion class
@@
 Therefore, any two non-zero classes in this space are necessarily cohomologically proportional.
-In a generic two-component SOC BEC, both Berry curvature classes $F^{(+)}]$ and $F^{(-)}]$ are non-zero; consequently, there exists a non-zero constant $\lambda\in\mathbb{R}$ such that
-$F^{(-)}]=\lambda[F^{(+)}]\quad\text{in }H^{2}(T^{2}_{\mathrm{BZ}}).$
+Under Assumption 2.1, we have
+$F^{(-)}]=\lambda[F^{(+)}]\quad\text{in }H^{2}(T^{2}_{\mathrm{BZ}}),\qquad\lambda=c^{(-)}/c^{(+)}.$
 Substituting this relation into (3.1) gives
@@
 In our geometry, we have $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}}$ (the space of parallel two-forms on the flat torus).
-The crucial observation, already made in Section 2.3, is that for the Kaluza–Klein metric $g_{M}$ the spaces of parallel one-forms on the phase circles are trivial:
-$\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},\,g_{M}|_{\text{\rm fibre}})=\{0\},$
-because any $g_{M}$-parallel one-form tangent to a fibre would have to be a constant multiple of $\Theta^{(\pm)}=d\phi_{\pm}+A^{(\pm)}$, and $\Theta^{(\pm)}$ is not closed (hence not of the form $d\phi_{\pm}$) unless $F^{(\pm)}=0$.
-Since the Berry curvatures $F^{(\pm)}$ are generically non-zero, we indeed have $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})=\{0\}$ in the physical geometry.
-By Lemma 4.5 (with $A^{(\pm)}$ non‑flat), the space of parallel 1‑forms on the phase circles vanishes: $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{M})=\{0\}$.
-Consequently the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(S^{1}_{\phi_{+}})\otimes\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ (the latter appears in $\ker\widetilde{\Psi}_{p}$) are zero.
-Thus
+The obstruction kernel $\mathcal{K}$ is shown to vanish by Lemma 2.5 under Assumption 2.1.
+The proof of Lemma 2.5 demonstrates that no non-zero element of the form $\operatorname{vol}_{\mathrm{BZ}}\otimes(c_{+}d\phi_{+}+c_{-}d\phi_{-})$ with $c_{+}+\lambda c_{-}=0$ can belong to $\mathcal{V}_{2,1}=\operatorname{span}\{\operatorname{vol}_{\mathrm{BZ}}\otimes(d\phi_{+}+\lambda d\phi_{-})\}$.
+Thus we have
 $\mathcal{K}=0,\qquad\text{and therefore}\qquad\dim\mathcal{K}=0.$ (3.2)
@@
 4 Geometric foundations for the PT theorem in Kaluza–Klein geometry
+The general topological lower bound of Pigazzini and Toda [12] is stated for genuine Riemannian product manifolds.
+Our physical metric (2.1), however, is a Kaluza–Klein metric that does not globally factor as a product.
+We now explain why, despite this apparent mismatch, the bound applies to our specific geometry.
+The argument relies on three observations that hold for the particular Kaluza–Klein structure arising from a two–component SOC BEC.
+Cohomology class and its harmonic representative
+The torsion 3–form $T^{b}=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}$ need not be closed, but its de Rham class $\omega]=[T^{b}]\in H^{3}(M;\mathbb{R})$ is well defined.
+(Recall that $\Theta^{(\pm)}=d\phi_{\pm}+\pi^{*}A^{(\pm)}$; note that the pullback $\pi^{*}\bigl(F^{(\pm)}\wedge A^{(\pm)}\bigr)$ vanishes because on the 2-dimensional base $T^{2}_{\mathrm{BZ}}$ we have $F^{(\pm)}\wedge A^{(\pm)}=0$.
+Therefore, $T^{b}$ is cohomologous to $F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}$.) Under the Künneth decomposition
+$H^{3}(M)\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
+the class $\omega]$ can be written as
+$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}].$
+Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the two Berry–curvature classes are necessarily proportional; hence there exists $\lambda\in\mathbb{R}$ such that $F^{(-)}]=\lambda[F^{(+)}]$.
+Consequently,
+$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
+which is a single simple tensor. The mixed tensor rank is therefore $r=1$, a purely topological invariant.
+Let $\omega_{h}$ be the unique $g_{M}$-harmonic representative of $\omega]$. Under Assumption 2.1, Lemma 3.1 gives an explicit form:
+$\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\bigl(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)}\bigr),$
+where $a_{\pm}$ are constants with $a_{-}=\lambda a_{+}$ and $\lambda=c^{(-)}/c^{(+)}$.
+This factorised form is the exact analogue, for our Kaluza–Klein metric, of the factorised harmonic decomposition used in the product-metric case.
+Spaces of parallel forms
+The bound involves the obstruction kernel
+$\mathcal{K}=\bigl(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}))\bigr)\oplus\bigl(\mathcal{V}_{1,2}\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{-}}))\bigr).$
+For the flat base we have $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\operatorname{vol}_{\mathrm{BZ}}$. By Lemma 2.5, the obstruction kernel vanishes: $\mathcal{K}=0$.
+Pointwise construction of the obstruction maps
+Even though the metric is not a product, at each point $p$ we have an orthogonal splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$.
+The non-integrability of $\mathcal{H}$ (measured by the Berry curvatures $F^{(\pm)}$) does not affect pointwise evaluations.
+Consider the (2,1)–component.
+Using the explicit harmonic representative from Lemma 3.1, we set
+$\mathcal{V}_{2,1}:=\operatorname{span}\{\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\}\subset H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}).$
+At a point $p=(k,\phi_{+},\phi_{-})$ choose any vectors $X,Y\in\mathcal{H}_{p}$ with $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$ and a vertical vector $Z\in\mathcal{V}_{p}$ such that $(a_{+}d\phi_{+}+a_{-}d\phi_{-})(Z)=1$.
+Define
+$\Psi_{p}\bigl(\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\bigr):=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z),$
+where $(\cdot)_{\mathrm{off}}$ denotes the part mapping $\mathcal{H}$ to $\mathcal{V}$ or vice versa.
+The map is well-defined because any other choice $(X^{\prime},Y^{\prime},Z^{\prime})$ satisfying the same conditions gives the same operator up to a constant factor, due to the linearity of the curvature in its arguments.
+A direct computation (provided in Appendix B) shows that for $\xi=\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\in\mathcal{V}_{2,1}$, we have:
+$\Psi_{p}(\xi)=\frac{1}{2}\left(a_{+}F^{(+)}(\pi(p))+a_{-}F^{(-)}(\pi(p))\right)J,$
+where $J:T_{p}M\to T_{p}M$ is a fixed off-diagonal operator (independent of $a_{\pm}$) that exchanges horizontal and vertical directions. The calculation uses:
+  1. 1.
+    $\nabla^{LC}(\operatorname{vol}_{\mathrm{BZ}}\wedge\Theta^{(\pm)})=\operatorname{vol}_{\mathrm{BZ}}\wedge\pi^{*}F^{(\pm)}$ (since $\operatorname{vol}_{\mathrm{BZ}}$ is parallel on the flat base);
+  2. 2.
+    the definition of $\Psi_{p}$ in terms of $(\nabla^{LC}T)_{\mathrm{off}}$;
+  3. 3.
+    the choice of vectors $X,Y\in\mathcal{H}_{p}$ with $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$ and $Z\in\mathcal{V}_{p}$ with $(a_{+}d\phi_{+}+a_{-}d\phi_{-})(Z)=1$.
+Therefore, $\Psi_{p}(\xi)=0$ if and only if $a_{+}F^{(+)}(\pi(p))+a_{-}F^{(-)}(\pi(p))=0$.
+Under Assumption 2.1, $F^{(-)}=\lambda F^{(+)}$ with $\lambda=c^{(-)}/c^{(+)}$.
+Moreover, since $\xi\in\mathcal{V}_{2,1}$, the coefficients satisfy $a_{-}=\lambda a_{+}$ (because $\mathcal{V}_{2,1}$ is spanned by $\operatorname{vol}_{\mathrm{BZ}}\otimes(d\phi_{+}+\lambda d\phi_{-})$).
+Substituting gives
+$a_{+}F^{(+)}+a_{-}F^{(-)}=a_{+}F^{(+)}+\lambda a_{+}(\lambda F^{(+)})=a_{+}(1+\lambda^{2})F^{(+)}.$
+Since $F^{(+)}\neq 0$, this vanishes precisely when $a_{+}=0$, which also implies $a_{-}=0$.
+Hence $\ker\Psi_{p}=\{0\}$ and $\operatorname{im}\Psi_{p}$ is one-dimensional.
+The same argument applied to the (1,2)–component gives a map $\widetilde{\Psi}_{p}$ with trivial kernel.
+Because the images of $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ consist of off–diagonal curvature operators, the Ambrose–Singer theorem forces
+$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;\dim\operatorname{im}\Psi_{p}+\dim\operatorname{im}\widetilde{\Psi}_{p}=1+0=1.$
+Thus the topological lower bound holds directly for the physical Kaluza–Klein metric $g_{M}$.
+The deformation family introduced below will serve only to connect this result to the product–metric case and to make the semicontinuity argument transparent.
+Note 1: The reasoning above does not require the metric to be a global Riemannian product.
+It uses only the pointwise orthogonal splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$, the explicit description of the harmonic representative (Lemma 3.1, which relies on the flatness of the base and the fibres under Assumption 2.1), and the vanishing of $\mathcal{P}_{1}$ on the fibres (Lemma 2.5).
+All these ingredients are present in the SOC BEC geometry.
+Deformation to a product metric and stability
+To provide a transparent link with the original product–metric setting, and to control the behaviour of the holonomy dimension via lower–semicontinuity, we now introduce a smooth deformation that connects the physical Kaluza–Klein metric to a genuine product metric.
+Note 2: [On the wedge decomposition in Kaluza–Klein geometry] In the product case, the harmonic representative decomposes as a sum of wedge products of harmonic forms on the factors.
+In our Kaluza–Klein geometry, the analogous decomposition (Lemma 3.1) involves the horizontal forms $\Theta^{(\pm)}$, which are harmonic with respect to the full metric $g_{M}$ but are not pullbacks from the fibres alone.
+However, they satisfy:
+$\Delta_{g_{M}}\Theta^{(\pm)}=0\quad\text{and}\quad[\Theta^{(\pm)}]=[d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}}),$
+which is sufficient for the cohomological arguments.
+The key point is that $\operatorname{vol}_{\mathrm{BZ}}\wedge\Theta^{(\pm)}$ represents the mixed class $F^{(\pm)}]\otimes[d\phi_{\pm}]$ in the Künneth decomposition, regardless of the non-product nature of the metric.
 4.1 Deformation to a product metric and stability
@@
   2. 2.
-    For $\varepsilon>0$, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{\varepsilon})=\{0\}$. For $\varepsilon=0$, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{0})=\mathbb{R}d\phi_{\pm}$.
+    The mixed tensor rank $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is constant for all $\varepsilon$.
   3. 3.
-    The mixed tensor rank $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is constant for all $\varepsilon$.
-  4. 4.
-    The obstruction kernel $\mathcal{K}_{\varepsilon}$ satisfies $\dim\mathcal{K}_{\varepsilon}=0$ for all $\varepsilon\in[0,1]$.
+    The obstruction kernel $\mathcal{K}_{\varepsilon}$ satisfies:
+    • (a)
+      $\dim\mathcal{K}_{0}=1$.
+    • (b)
+      $\dim\mathcal{K}_{\varepsilon}=0$ for every $\varepsilon>0$.
+    Consequently, the corrected rank $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}$ equals 0 for $\varepsilon=0$ and equals $1$ for every $\varepsilon>0$.
 Proof.
 (1) follows because the restriction of $g_{\varepsilon}$ to $T^{2}_{\mathrm{BZ}}$ is $g_{\mathrm{BZ}}$, independent of $\varepsilon$.
-(2) For $\varepsilon>0$, Lemma 2.4 shows $d\phi_{\pm}$ are not parallel; the proof of Lemma 4.5 will show no other parallel 1-forms exist.
-For $\varepsilon=0$, $d\phi_{\pm}$ are clearly parallel.
-(3) is Lemma 4.6 (metric invariance).
-(4) For $\varepsilon=0$, the harmonic representative of $\omega]$ is
-$\omega_{h}=F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}.$
-The subspace $\mathcal{P}_{2}(T^{2})\otimes\mathcal{P}_{1}(S^{1}_{+}\times S^{1}_{-})$ consists of those harmonic 3‑forms that can be written as
-$\alpha\wedge\beta,$
-where $\alpha$ is a parallel 2‑form on $T^{2}$ (hence a constant multiple of $\text{vol}_{\text{BZ}}$) and $\beta$ is a parallel 1‑form on the fibres (hence a constant linear combination of $d\phi_{+}$ and $d\phi_{-}$).
-If $\omega_{h}$ belonged to this subspace, we would have
-$F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}=\alpha\wedge\beta=(\lambda\text{vol}_{\text{BZ}})\wedge(c_{+}d\phi_{+}+c_{-}d\phi_{-}),$
-which would imply $F^{(+)}=\lambda c_{+}\text{vol}_{\text{BZ}}$ and $F^{(-)}=\lambda c_{-}\text{vol}_{\text{BZ}}$ as differential forms, not just in cohomology.
-In a generic SOC texture, $F^{(+)}$ and $F^{(-)}$ are not parallel (they are not constant multiples of $\text{vol}_{\text{BZ}}$ pointwise), hence $\omega_{h}\notin\mathcal{P}_{2}\otimes\mathcal{P}_{1}$.
-Therefore $\mathcal{K}_{0}=0$.
-∎
+(2) The mixed tensor rank is a topological invariant (Lemma 4.7) and therefore constant under metric deformations.
+(3a) For the product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$, the Brillouin torus is flat.
+On a flat Riemannian manifold every harmonic form is parallel; thus $\mathcal{H}^{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})=\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$.
+The harmonic representative of the torsion class $\omega]$ with respect to $g_{0}$ is
+$\omega_{h}^{(0)}=\alpha_{+}\wedge d\phi_{+}+\alpha_{-}\wedge d\phi_{-},$
+where $\alpha_{\pm}$ are the harmonic representatives of the Berry‑curvature classes $F^{(\pm)}]$.
+Since $\alpha_{\pm}$ are parallel and $d\phi_{\pm}$ are parallel, $\omega_{h}^{(0)}$ belongs to $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{+}\times S^{1}_{-})$.
+Hence $\mathcal{K}_{0}=\mathcal{V}_{2,1}$ and $\dim\mathcal{K}_{0}=1$.
+(3b) For $\varepsilon>0$ the metric $g_{\varepsilon}$ is a Kaluza–Klein metric satisfying Assumption 2.1 (with curvatures scaled by $\varepsilon$).
+By Lemma 2.5, the obstruction kernel vanishes: $\mathcal{K}_{\varepsilon}=0$.
+Thus $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}$ equals 0 for $\varepsilon=0$ and equals $1$ for every $\varepsilon>0$. ∎
 Lemma 4.3 (Lower semi-continuity of off-diagonal holonomy dimension).
@@
 Proof.
-By Lemma 4.2, the topological data $r^{\sharp}=r-\dim\mathcal{K}_{\varepsilon}$ is constant (=r=r) for all $\varepsilon$.
-The inequality holds at $\varepsilon=0$ by assumption.
-By Lemma 4.3, the set $\{\varepsilon:\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq r^{\sharp}\}$ is closed (by lower semi-continuity).
-Since it contains $\varepsilon=0$ and $[0,1]$ is connected, it must be all of $[0,1]$.
-∎
-4.2 Parallel forms in Kaluza–Klein metrics
-Lemma 4.5 (Parallel forms in Kaluza–Klein metrics).
-Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a Kaluza–Klein metric
-$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$
-where $g_{1}$ is a Riemannian metric on $M_{1}$, $\{\theta^{a}\}$ defines a flat left-invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$.
-  1. 1.
-    If all connection forms $A^{a}$ are closed (i.e., $dA^{a}=0$), then every $h$-parallel form on the fibre is also $g$-parallel.
-  2. 2.
-    If at least one curvature $F^{a}:=dA^{a}$ is not identically zero, then the space of $g$-parallel 1-forms tangent to the fibres is trivial:
-    $\mathcal{P}_{1}(M_{2},g)=\{0\}.$
-Proof.
-Write the metric locally as $g=g_{ij}(x)dx^{i}dx^{j}+\sum_{a=1}^{m}(\Theta^{a})^{2}$, where $\Theta^{a}=\theta^{a}+A^{a}_{i}(x)dx^{i}$ and $\{\theta^{a}\}$ is the orthonormal coframe on $M_{2}$ induced by the flat metric $h$.
-In our SOC BEC case, $m=2$, $\theta^{1}=d\phi_{+}$, $\theta^{2}=d\phi_{-}$.
-Let $\xi=\sum_{a=1}^{m}f_{a}(x,y)\Theta^{a}$ be a 1‑form tangent to the fibres.
-The Levi‑Civita connection of $g$ satisfies the following relation for a horizontal vector field $X=X^{i}\partial_{i}$ and the vertical coordinate field $V_{a}$ dual to $\theta^{a}$:
-$\nabla^{LC}_{X}V_{a}=\tfrac{1}{2}F^{a}(X)^{\#},$
-where $F^{a}(X)^{\#}$ is the vertical vector field obtained by raising the index of the 1‑form $F^{a}(X,\cdot)$ with the fibre metric $h$.
-Formula (i) follows from the Koszul formula and uses only that the fibres are flat.
-Now suppose $\xi$ is parallel, $\nabla^{LC}\xi=0$. Evaluating on the pair $(X,V_{a})$ gives
-$0=(\nabla^{LC}_{X}\xi)(V_{a})=X\bigl(\xi(V_{a})\bigr)-\xi(\nabla^{LC}_{X}V_{a})=X(f_{a})-\tfrac{1}{2}\sum_{b=1}^{m}f_{b}\,h\!\bigl(F^{a}(X)^{\#},V_{b}\bigr).$
-Because the functions $f_{a}$ are constant along the fibres (they depend only on the base coordinates $x$ due to the left‑invariance of the metric $h$), equation (ii) reduces to
-$\tfrac{1}{2}\sum_{b=1}^{m}f_{b}(x)\,F^{b}(X,V_{a})=0\qquad\text{for every horizontal }X\text{ and every }a.$
-If all $F^{b}=0$, condition (iii) is vacuous and any choice of constant $f_{b}$ yields a parallel form, proving statement (i).
-If, however, some $F^{b_{0}}$ is not identically zero, we can choose a horizontal vector field $X$ and an index $a$ such that $F^{b_{0}}(X,V_{a})\neq 0$ at some point; then (iii) forces $f_{b_{0}}=0$.
-Repeating the argument for every independent curvature forces all coefficients $f_{b}$ to vanish, hence $\xi=0$.
-This proves statement (ii) for 1‑forms.
-The argument for higher‑degree forms is analogous, using that a parallel $k$-form on the fibre must have constant coefficients with respect to the left‑invariant coframe, and the same curvature coupling appears in the covariant derivative.
-∎
-Lemma 4.6 (Harmonic representative and Künneth component).
-Let $g$ be a Kaluza–Klein metric as in Definition 4.1 on $M=M_{1}\times M_{2}$.
-Let $\omega]\in H^{k}(M;\mathbb{R})$ be a cohomology class and $\omega_{h}$ its unique $g$-harmonic representative.
-Then $\omega_{h}$ respects the Künneth decomposition in the sense that its projection onto the $(p,q)$-bidegree component corresponds to the harmonic representative of the mixed class $\omega]_{p,q}\in H^{p}(M_{1})\otimes H^{q}(M_{2})$.
-In particular, $\omega_{h}$ admits a unique decomposition
-$\omega_{h}=\sum_{i=1}^{r}\alpha_{i}\wedge\beta_{i}+\gamma,$
-where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$, the forms $\alpha_{i},\beta_{i}$ are the lifted harmonic representatives of the factor classes, and $\gamma$ contains terms of different bidegrees.
-Proof.
-By Hodge theory on compact Riemannian manifolds, every de Rham cohomology class $\omega]$ possesses a unique harmonic representative $\omega_{h}$ with respect to the metric $g$.
-The Künneth theorem provides a canonical isomorphism $H^{k}(M)\cong\bigoplus_{p+q=k}H^{p}(M_{1})\otimes H^{q}(M_{2})$.
-While the Hodge Laplacian $\Delta_{g}$ of a Kaluza-Klein metric does not generally commute with the factor-wise Laplacians $\Delta_{g_{1}}$ and $\Delta_{h}$ due to the presence of the connection forms $A^{a}$, the existence of the harmonic representative for each mixed component is guaranteed by the ellipticity of $\Delta_{g}$.
-The tensor rank $r$ is a purely algebraic property of the class $\omega]$ in the tensor product of the cohomology groups $H^{*}(M_{i})$ and is therefore independent of the specific metric $g$.
-The decomposition follows by projecting the global harmonic form $\omega_{h}$ onto the subspace of forms of pure bidegree, which remains well-defined under the orthogonal splitting $TM=D_{1}\oplus D_{2}$ induced by the Kaluza-Klein structure.
+By Lemma 4.2, the corrected rank $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}$ is constant on $(0,1]$ and equals $1$.
+For every $\varepsilon>0$, Theorem 4.8 (the extension of the PT theorem to Kaluza–Klein metrics) therefore yields the non‑trivial lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1$.
+The function $\varepsilon\mapsto\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower‑semicontinuous by Lemma 4.3; hence the set
+$S=\{\varepsilon\in[0,1]:\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1\}$
+is closed.
+Since $S$ contains the interval $(0,1]$, its closure contains the endpoints $\varepsilon=0$ and $\varepsilon=1$.
+Because $S$ is closed, we must have $0\in S$ and $1\in S$.
+In particular, $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{1}})\geq 1$, which is the desired bound for the physical metric.
 ∎
@@
 With the previous preparatory lemmas, we can now state and prove the extension of the Pigazzini-Toda theorem to Kaluza-Klein geometries.
-Theorem 4.10 (PT bound for Kaluza–Klein metrics with flat fibres).
-Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold equipped with a Kaluza–Klein metric
-$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$
-where $\{\theta^{a}\}$ defines a flat left‑invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$.
-Let $\nabla^{C}=\nabla^{LC}(g)+\frac{1}{2}T$ be a metric connection on $(M,g)$ with totally skew‑symmetric torsion $T\in\Omega^{3}(M)$.
-Denote by $\omega_{h}$ the unique $g$-harmonic representative of $T]\in H^{3}(M;\mathbb{R})$ and assume
-  1. 1.
-    $\nabla^{LC}(g)T\neq 0$;
-  2. 2.
-    $\omega_{h}]$ has a non‑trivial mixed component in the Künneth decomposition;
-  3. 3.
-    the mixed tensor rank $r:=rank_{\mathbb{R}}\bigl([\omega_{h}]_{\mathrm{mixed}}\bigr)$ is finite.
-Define the obstruction space
-$\mathcal{K}\;:=\;\bigl(\pi_{2,1}([\omega_{h}])\cap(\mathcal{P}_{2}(M_{1},g)\otimes\mathcal{P}_{1}(M_{2},g))\bigr)\oplus\bigl(\pi_{1,2}([\omega_{h}])\cap(\mathcal{P}_{1}(M_{1},g)\otimes\mathcal{P}_{2}(M_{2},g))\bigr),$
-and set $r^{\sharp}:=r-\dim\mathcal{K}$.
-Then for every point $p\in M$,
-$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r^{\sharp}.$ (4.0)
-Moreover, $r^{\sharp}$ depends only on the cohomology class $\omega_{h}]_{\mathrm{mixed}}$ and on the spaces $\mathcal{P}_{k}(M_{i},g)$; consequently, it is invariant under smooth deformations of $g$ that preserve those spaces.
-Proof.
-The result follows from a direct application of the topological bound established in [12] to the Kaluza–Klein geometry $(M,g,\nabla^{C})$.
-We verify that the structural requirements are met and compute the resulting invariant.
-First, we observe that the metric connection $\nabla^{C}$ satisfies the requisite non-degeneracy conditions.
-The torsion form $T$, constructed from the synthetic gauge fields, is totally skew-symmetric.
-Moreover, the condition $\nabla^{LC}(g)T\neq 0$ is satisfied whenever the Berry curvature $F=dA\neq 0$, which holds by assumption in the generic SOC regime.
-The mixed tensor rank $r=\text{rank}_{\mathbb{R}}([\omega_{h}]_{\mathrm{mixed}})$ is a cohomological invariant of the class $[T]$; it is independent of the Riemannian metric used to define the harmonic representative, ensuring that $r$ is well-defined and finite.
-Crucially, the dimension of the obstruction kernel $\mathcal{K}$ is determined by the intersection of the mixed cohomology components with the spaces of parallel forms.
-By Lemma 4.5, the non-vanishing Berry curvature $F=dA\neq 0$ implies that the space of fibre-parallel 1-forms vanishes:
-$\mathcal{P}_{1}(M_{2},g)=\{0\}.$
-Consequently, any tensor product involving $\mathcal{P}_{1}(M_{2},g)$ is zero, and the obstruction space $\mathcal{K}$ is identically zero in the generic SOC BEC regime where synthetic gauge fields are present.
-Thus, we have $\dim\mathcal{K}=0$, and the general theorem of [12] then states that the dimension of the off-diagonal holonomy is bounded by $r-\dim\mathcal{K}$, yielding
-$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r.$
-The invariance of this bound follows from the stability of the topological rank $r$ and the robustness of the vanishing condition $\mathcal{P}_{1}(M_{2},g)=\{0\}$ under metric deformations that preserve the non-integrability of the horizontal distribution.
-∎
+Theorem 4.8 (PT bound for the SOC BEC geometry).
+Let $M=T_{\mathrm{BZ}}^{2}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be equipped with the Kaluza-Klein metric $g_{M}$ as in (2.1), and let $\nabla^{C}=\nabla^{LC}+\frac{1}{2}T$ be the metric connection with torsion $T=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}$.
+Assume that $\nabla^{LC}T\neq 0$ and that the Berry curvature classes satisfy $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}})$.
+Then the off-diagonal holonomy of $\nabla^{C}$ satisfies
+$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq 1\quad\text{for every }p\in M.$
+Proof for the SOC BEC geometry.
+We prove the bound via deformation and lower semi-continuity, without requiring global integrability of the horizontal distribution.
+Consider the family $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ from Definition 4.1.
+For each $\varepsilon>0$, the orthogonal splitting $T_{p}M=\mathcal{H}^{(\varepsilon)}_{p}\oplus\mathcal{V}_{p}$ (horizontal/vertical) is pointwise well-defined.
+Although for $\varepsilon>0$ this splitting is not integrable unless $F^{(\pm)}=0$, the original Pigazzini–Toda theorem [12] uses only the pointwise orthogonal decomposition of the tangent space, not its integrability.
+Indeed, the proof of Theorem 5.2 in [12] relies solely on algebraic computations with curvature operators and on the properties of harmonic forms, which remain valid for any orthogonal splitting.
+For $\varepsilon>0$, Lemma 2.5 implies that the obstruction kernel vanishes: $\mathcal{K}_{\varepsilon}=0$.
+The mixed tensor rank $r=1$ is topological (Lemma 4.7), so $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}=1$.
+Applying Theorem 5.2 of [12] to the splitting $\mathcal{H}^{(\varepsilon)}\oplus\mathcal{V}$ yields
+$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq r^{\sharp}_{\varepsilon}=1\quad\forall\varepsilon>0,$
+where $\nabla^{C_{\varepsilon}}$ is the connection built from $g_{\varepsilon}$ with the same torsion form $T$.
+This application is justified because Theorem 5.2 of [12] requires only a pointwise orthogonal decomposition $T_{p}M=V_{1}\oplus V_{2}$ and the existence of the harmonic representative $\omega_{h}$ with the stated properties.
+The non-integrability of $\mathcal{H}^{(\varepsilon)}$ (when $\varepsilon>0$) does not affect the proof, which is based on pointwise algebraic computations with curvature operators.
+Define $F(\varepsilon):=\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ for $\varepsilon\in[0,1]$.
+Then $F$ is integer-valued (a dimension) and, by Lemma 4.3, lower semi-continuous on $[0,1]$.
+We now prove $F(1)\geq 1$.
+For an integer-valued lower semi-continuous function, the set $U:=\{\varepsilon\in[0,1]\mid F(\varepsilon)\geq 1\}$ is open in $[0,1]$: if $\varepsilon_{0}\in U$, lower semi-continuity gives $\delta>0$ such that $F(\varepsilon)>F(\varepsilon_{0})-\frac{1}{2}$ for all $\varepsilon$ in $(\varepsilon_{0}-\delta,\varepsilon_{0}+\delta)\cap[0,1]$.
+Since $F(\varepsilon_{0})\geq 1$ and $F$ is integer, this implies $F(\varepsilon)\geq 1$ on that neighbourhood, hence the neighbourhood is contained in $U$.
+We have shown $F(\varepsilon)\geq 1$ for all $\varepsilon>0$; thus $(0,1]\subseteq U$.
+Because $U$ is open and contains $(0,1]$, any open neighbourhood of $1$ in $[0,1]$ intersects $(0,1)\subset U$; openness of $U$ then forces $1\in U$.
+Therefore $F(1)\geq 1$, i.e., $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{1}})\geq 1$.
+Since $\nabla^{C_{1}}=\nabla^{C}$ and $g_{1}=g_{M}$, the bound holds for the physical metric. ∎
 4.5 Application to the SOC BEC model
@@
   1. 1.
-    $\mathcal{P}_{1}(M_{2},g)=\{0\}$, due to the non-triviality of the synthetic gauge potentials $A^{(\pm)}$.
+    The obstruction kernel vanishes: $\mathcal{K}=0$, as shown by Lemma 2.5 under Assumption 2.1.
   2. 2.
@@
 The identification of the BEC metric with the Kaluza–Klein framework follows directly from the construction in Section 2.
-The first item is a direct consequence of Lemma 4.5; as established in the physical model, the synthetic gauge potentials $A^{(\pm)}$ possess non-vanishing Berry curvature, which acts as an obstruction to the existence of parallel 1-forms along the fibres.
+The first item follows directly from Lemma 2.5; under Assumption 2.1 the non‑vanishing Berry curvature forces $\mathcal{K}=0$.
 On the base manifold $M_{1}=T^{2}_{\mathrm{BZ}}$, the restriction map onto $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ is an isomorphism, identifying the space of parallel 2-forms with the span of the Riemannian volume form.
@@
 5 Topological Lower Bound and Its Physical Consequences
-With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now apply the general theorem of [12] to the physical connection $\nabla^{C}$.
+With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now apply the extended Pigazzini-Toda theorem (Theorem 4.8) to the physical connection $\nabla^{C}$.
 This section establishes the fundamental lower bound on the holonomy algebra for the SOC BEC model.
 5.1 Statement of the bound
-The general PT theorem establishes that for a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on a product Riemannian manifold $M=M_{1}\times M_{2}$, the dimension of the off-diagonal holonomy is constrained by the topology of the torsion class and the geometry of the metric.
-Specifically, if the condition $\nabla^{LC}T\neq 0$ holds, we have:
+The extended PT theorem (Theorem 4.8) establishes that for a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on the Kaluza-Klein manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the dimension of the off-diagonal holonomy is constrained by the topology of the torsion class and the geometry of the metric.
+Specifically, under Assumption 2.1 and with $\nabla^{LC}T\neq 0$, we have:
 $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r-\dim\mathcal{K},$ (5.1)
@@
 Theorem 5.1 (Topological lower bound for SOC BECs).
-Let $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be the parameter space equipped with the Kaluza–Klein metric $g_{M}$ and the metric connection $\nabla^{C}$ with torsion $T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}$.
-Assuming the Berry curvatures satisfy $F^{(\pm)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{R})$, the following hold:
+Let $M=T_{\mathrm{BZ}}^{2}\times S_{\phi_{+}}^{1}\times S_{\phi_{-}}^{1}$ be the parameter space equipped with the Kaluza-Klein metric $g_{M}$ (2.1) and the metric connection $\nabla^{C}$ with torsion $T^{b}=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}$.
+Under Assumption 2.1 (constant Berry curvatures), the following hold:
   1. 1.
@@
 Proof.
-The proof proceeds by synthesizing the topological and geometric invariants derived in the preceding sections.
+The proof proceeds by synthesizing the topological and geometric invariants derived in the preceding sections and applying the deformation argument developed in Section 4.
 1.
@@
 Under the Künneth decomposition for $H^{3}(M;\mathbb{R})$, the torsion class is expressed as $\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$.
-Since the cohomology group $H^{2}(T^{2}_{\mathrm{BZ}})$ is one-dimensional, any two non-zero classes are linearly dependent.
-Thus, we can write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$, which allows the class to be factorized as the simple tensor:
+Since the cohomology group $H^{2}(T^{2}_{\mathrm{BZ}})$ is one-dimensional and both $F^{(\pm)}]$ are non-zero under Assumption 2.1, we can write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda=c^{(-)}/c^{(+)}\in\mathbb{R}$, which allows the class to be factorized as the simple tensor:
 $\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr).$
@@
 The space $\mathcal{K}$ is defined by the intersection of the mixed components of $\omega]$ with the subspaces of forms $\mathcal{P}_{k}(M_{i},g_{M})$ that are parallel with respect to the Levi-Civita connection of $g_{M}$.
-Crucially, Lemma 4.5 proves that for a Kaluza-Klein metric, the existence of non-vanishing Berry curvature $F\neq 0$ acts as a differential-geometric obstruction to the existence of parallel 1‑forms along the fibres.
-Consequently, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{M})=\{0\}$.
-Since the mixed components of $\omega]$ involve 1-forms on the fibre factors, their intersection with the parallel-form strata is trivial, leading to $\dim\mathcal{K}=0$.
+Under Assumption 2.1, Lemma 2.5 proves that the obstruction kernel vanishes: $\mathcal{K}=\{0\}$.
 3.
-Application of the PT bound.
-The Kaluza–Klein metric $g_{M}$ satisfies the conditions for the extended PT theorem as stated in Theorem 4.10.
-Since the non-vanishing Berry curvature acts as a topological obstruction that prevents the existence of global parallel sections, the dimension of the obstruction space $\mathcal{K}$ is zero; consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$ without recourse to auxiliary product metrics.
-Substituting the invariants $r=1$ and $\dim\mathcal{K}=0$ into the inequality (5.1), we obtain $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$.
+Application via deformation and lower semi-continuity.
+Consider the deformation family $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ from Definition 4.1, which interpolates between the product metric $g_{0}$ and the physical Kaluza-Klein metric $g_{M}=g_{1}$.
+For each $\varepsilon>0$, the metric $g_{\varepsilon}$ satisfies Assumption 2.1 with curvatures scaled by $\varepsilon$.
+By Lemma 4.2, we have $\mathcal{K}_{\varepsilon}=0$ and $r^{\sharp}_{\varepsilon}=1$ for all $\varepsilon>0$.
+Applying Theorem 4.8 (the extension of the PT theorem to Kaluza-Klein metrics with flat fibres) to each $g_{\varepsilon}$ with $\varepsilon>0$, we obtain
+$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1\quad\forall\varepsilon>0,$
+where $\nabla^{C_{\varepsilon}}$ is the connection built from $g_{\varepsilon}$ with the same torsion form $T$.
+By Lemma 4.3, the function $\varepsilon\mapsto\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower semi-continuous on $[0,1]$. Therefore,
+$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{1}})\geq\liminf_{\varepsilon\to 1}\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1.$
+Since $\nabla^{C_{1}}=\nabla^{C}$ and $g_{1}=g_{M}$, this establishes the bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ for the physical metric.
 The geometric implication is that there exists at least one independent curvature operator $R^{C}(X,Y)$ that interchanges the tangent spaces of the Brillouin zone and the phase fibres.
@@
 If $c_{1}^{(+)}\neq 0$ and $c_{1}^{(-)}=-c_{1}^{(+)}$, then $c_{1}^{\,\mathrm{tot}}=0$.
-Nevertheless, provided $F^{(+)}]\neq 0$ (hence also $F^{(-)}]\neq 0$), the mixed tensor rank remains $r=1$ and Theorem 5.1 still gives
+Nevertheless, under Assumption 2.1 with $c^{(\pm)}\neq 0$, the mixed tensor rank remains $r=1$ and Theorem 5.1 still gives
 $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
@@
 Hence both classes are non-zero and the mixed tensor rank attains its maximum $r=1$.
-The vanishing of $\mathcal{K}$ does not depend on the values of the Chern numbers but only on the non-existence of parallel one-forms on the fibres, which holds as long as $F^{(\pm)}\neq 0$.
+The vanishing of $\mathcal{K}$ does not depend on the values of the Chern numbers but only on the non-existence of parallel one-forms on the fibres, which holds as long as $F^{(\pm)}\neq 0$ (guaranteed by Assumption 2.1).
 ∎
+This corollary is consistent with the geometric setup of Section 2: the global product structure of $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ already implies $c_{1}^{\mathrm{tot}}=0$.
+The example $c_{1}^{(+)}=-c_{1}^{(-)}$ is therefore not a special case but a generic realization of the trivial-bundle regime under constant curvatures with $\lambda=-1$.
 This demonstrates that $r^{\sharp}$ detects a local topological constraint that survives when the familiar global invariant vanishes.
 While the Chern number quantizes the net flux, it can be zero due to cancellation of local contributions; the mixed-rank bound, in contrast, certifies that the curvature cannot be simultaneously flattened along the two independent phase directions.
+The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ thus captures a strictly local topological obstruction that remains active even when global charges cancel.
 5.5 Invariance under metric deformations
@@
 Any smooth deformation of the system that does not change the cohomology classes of $A^{(\pm)}$ and that preserves the vanishing of $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})$ (i.e., that does not make the Berry curvatures exactly flat) leaves the bound unchanged.
-This stability makes the result robust against typical experimental perturbations.
+This stability is ensured by Lemma 4.5, which guarantees that the non-parallelism condition $\nabla^{LC}T\neq 0$ is open in the space of metrics.
+This makes the result robust against typical experimental perturbations.
 The inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ is the main result of this paper.
@@
 The primary physical advantage of the holonomy bound lies in its sensitivity to the local structure of the Berry curvature, whereas the total Chern number $c_{1}^{\mathrm{tot}}$ is a global aggregate that may vanish due to cancellation.
+In the constant-curvature setting of Assumption 2.1, the local field strength is captured by the constants $c^{(\pm)}$.
 In configurations where the individual fluxes satisfy $F^{(+)}]\approx-[F^{(-)}]$, the global invariant $c_{1}^{\mathrm{tot}}$ vanishes, potentially suggesting a topologically trivial regime.
-However, the inequality $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ persists as long as the local field strength remains non-zero.
+However, the inequality $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ persists as long as both constants $c^{(+)}$ and $c^{(-)}$ are non-zero, regardless of their sum.
 This identifies $r^{\sharp}$ as a robust local invariant: it detects the persistence of topological effects, such as the non-integrability of the Berry connection, even in compensated systems or “synthetic vacuum” configurations where net topological charges are absent.
@@
 $\omega]=[F]\otimes\bigl([d\phi_{+}]+[d\phi_{-}]\bigr).$
-Consequently, the mixed tensor rank is $r=1$. Given that the obstruction kernel $\mathcal{K}$ vanishes in this geometry, we obtain $r^{\sharp}=1$. Theorem 5.1 then yields:
+Consequently, the mixed tensor rank is $r=1$.
+Given that the obstruction kernel $\mathcal{K}$ vanishes in this geometry (as shown in Lemma 2.5 with $\lambda=1$), we obtain $r^{\sharp}=1$.
+Theorem 5.1 then yields:
 $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
-We note that for a global Rashba texture with non-zero Chern number, the bundle is topologically non-trivial, preventing a global product decomposition.
-However, applying the theorem within a local trivialization demonstrates that the mixed rank $r^{\sharp}$ correctly identifies the ’locking’ of momentum and phase degrees of freedom.
-In this context, while the non-vanishing Chern number already implies the existence of vortices on the Brillouin zone, the invariant $r^{\sharp}$ provides an algebraically refined constraint, certifying that the momentum and phase degrees of freedom remain fundamentally coupled through the off-diagonal curvature.
+Remark 7.1.
+For a global Rashba texture with non-zero Chern number, the underlying principal bundle is topologically non-trivial, preventing a global product decomposition.
+However, as noted in Section 2, our analysis can be applied within local trivializations.
+The mixed rank $r^{\sharp}$ correctly identifies the ’locking’ of momentum and phase degrees of freedom even in such local patches.
+While the non-vanishing Chern number already implies the existence of vortices on the Brillouin zone, the invariant $r^{\sharp}$ provides an algebraically refined constraint, certifying that the momentum and phase degrees of freedom remain fundamentally coupled through the off-diagonal curvature.
 7.2 Example 2: Topological persistence under vanishing net Chern flux
@@
 Such a configuration is protected by a local topological obstruction that is invisible to integrated invariants but perfectly captured by the mixed cohomology rank on the product manifold.
-To illustrate how the bound materializes geometrically, consider the concrete choice
+To illustrate how the bound materializes geometrically, consider the concrete choice (satisfying Assumption 2.1)
 $A^{(+)}=\frac{1}{4\pi}(k_{x}dk_{y}-k_{y}dk_{x}),\qquad A^{(-)}=-A^{(+)},$
@@
 The torsion 3-form is
-$T=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}=\frac{1}{8\pi^{2}}(dk_{x}\wedge dk_{y})\wedge(k_{x}dk_{y}-k_{y}dk_{x}).$
+$T=F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}=\frac{1}{2\pi}\,dk_{x}\wedge dk_{y}\wedge(d\phi_{+}-d\phi_{-}).$
 For the connection $\nabla^{C}=\nabla^{LC}+\tfrac{1}{2}T$, the off-diagonal curvature components evaluate to
@@
     The geometric obstruction manifests as a direct coupling between momentum derivatives $(\partial_{k_{x}},\partial_{k_{y}})$ and phase derivatives $(\partial_{\phi_{+}},\partial_{\phi_{-}})$, preventing any gauge transformation from simultaneously diagonalizing the connection with respect to the product splitting $T_{\mathrm{BZ}}^{2}\times S^{1}\times S^{1}$.
+The fact that $\dim\mathfrak{hol}^{\mathrm{off}}=2$ in this specific example shows that the bound $\geq 1$ is sharp in the sense that it cannot be improved without additional assumptions, yet particular geometries may yield a larger off-diagonal holonomy algebra.
 7.3 Comparison: Global Invariants vs. Mixed Rank
@@
 This phenomenon can be visualized as the Berry curvature being constrained to move along topological “rail-tracks” in the configuration space: even when the total topological charge is tuned to zero, specific directional components remain topologically locked and cannot be erased by any smooth gauge transformation.
-In the two-component SOC BEC manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the bound $r^{\sharp}=1$ is non-degenerate since the non-vanishing Berry curvature breaks the integrability of the horizontal distribution, ensuring $\mathcal{K}=0$.
+In the two-component SOC BEC manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the bound $r^{\sharp}=1$ is non-degenerate.
+Under Assumption 2.1, the constant, non-zero Berry curvatures break the integrability of the horizontal distribution, ensuring $\mathcal{K}=0$.
 Consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$.
@@
 The extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, equipped with the natural Kaluza–Klein metric $g_{M}$ induced by synthetic gauge fields, carries a metric connection $\nabla^{C}$ with totally skew-symmetric torsion.
-The harmonic part of this torsion represents a mixed cohomology class $\omega]$ with mixed tensor rank $r=1$ for any generic configuration where the individual Berry curvature classes are non-trivial.
+Under Assumption 2.1 (constant Berry curvatures), the harmonic part of this torsion takes the explicit form given in Lemma 3.1 and represents a mixed cohomology class $\omega]$ with mixed tensor rank $r=1$ for any generic configuration where the individual Berry curvature classes are non-trivial.
 Due to the one-dimensionality of $H^{2}(T^{2}_{\mathrm{BZ}})$, these classes are necessarily proportional, forcing $\omega]$ to take the form of a single simple tensor.
@@
 The authors gratefully acknowledge stimulating discussions with colleagues in geometry and cold-atom physics that helped shape the formulation of these results.
+Appendix
+Appendix A Parallel Vertical 1-Forms
+This appendix provides the detailed proof of the condition for vertical 1-forms to be parallel, which is used in the proof of Lemma 2.5. We consider the Kaluza-Klein metric
+$g_{M}=\pi^{*}g_{\mathrm{BZ}}+\Theta^{(+)}\otimes\Theta^{(+)}+\Theta^{(-)}\otimes\Theta^{(-)},$
+with $\Theta^{(\pm)}=d\phi_{\pm}+\pi^{*}A^{(\pm)}$ and $F^{(\pm)}=dA^{(\pm)}=c^{(\pm)}\operatorname{vol}_{\mathrm{BZ}}$ (Assumption 2.1). Let $\lambda=c^{(-)}/c^{(+)}$.
+Orthonormal Frame
+Choose coordinates $(x,y)$ on $T^{2}_{\mathrm{BZ}}$ such that $g_{\mathrm{BZ}}=dx^{2}+dy^{2}$ and $\operatorname{vol}_{\mathrm{BZ}}=dx\wedge dy$. An orthonormal frame for $g_{M}$ is given by
+$e_{1}=\partial_{x},\quad e_{2}=\partial_{y},\quad e_{3}=\partial_{\phi_{+}},\quad e_{4}=\partial_{\phi_{-}},$
+with the corresponding coframe
+$e^{1}=dx,\quad e^{2}=dy,\quad e^{3}=\Theta^{(+)},\quad e^{4}=\Theta^{(-)}.$
+Connection 1-Forms
+The connection 1-forms $\omega^{a}_{\;b}$ for the Levi-Civita connection are defined by
+$\nabla^{LC}e_{a}=\omega_{a}^{\;b}\otimes e_{b},$
+where $\omega_{ab}=g_{M}(\nabla^{LC}e_{a},e_{b})$ are skew-symmetric. For our Kaluza-Klein metric with flat fibers and constant curvatures, the non-zero connection 1-forms are:
+$\omega_{13}$ $=-\omega_{31}=\frac{1}{2}c^{(+)}e^{2},$
+$\omega_{14}$ $=-\omega_{41}=\frac{1}{2}c^{(-)}e^{2},$
+$\omega_{23}$ $=-\omega_{32}=-\frac{1}{2}c^{(+)}e^{1},$
+$\omega_{24}$ $=-\omega_{42}=-\frac{1}{2}c^{(-)}e^{1}.$
+All other $\omega_{ab}=0$. This can be derived from the structure equations $de^{a}+\omega^{a}_{\;b}\wedge e^{b}=0$ and the known expressions for $de^{3}$ and $de^{4}$.
+Parallel Vertical 1-Forms
+A vertical 1-form can be written as
+$\eta=h_{+}\Theta^{(+)}+h_{-}\Theta^{(-)}=h_{+}e^{3}+h_{-}e^{4},$
+where $h_{\pm}\in C^{\infty}(M)$. The condition for $\eta$ to be parallel, $\nabla^{LC}\eta=0$, becomes
+$\nabla^{LC}\eta=(dh_{+}\otimes e^{3}+h_{+}\nabla^{LC}e^{3})+(dh_{-}\otimes e^{4}+h_{-}\nabla^{LC}e^{4})=0.$
+Using $\nabla^{LC}e^{3}=-\omega_{3}^{\;b}\otimes e_{b}=\omega_{13}e^{1}\otimes e_{1}+\omega_{23}e^{2}\otimes e_{1}+\omega_{31}e^{1}\otimes e_{2}+\omega_{32}e^{2}\otimes e_{2}$ (and similarly for $e^{4}$), we compute the components of $\nabla^{LC}\eta$.
+The components along $e^{1}\otimes e^{3}$ and $e^{2}\otimes e^{3}$ give:
+$\frac{\partial h_{+}}{\partial x}=0,\quad\frac{\partial h_{+}}{\partial y}=0,$
+so $h_{+}$ is constant along the base. Similarly, from the components along $e^{1}\otimes e^{4}$ and $e^{2}\otimes e^{4}$, we get that $h_{-}$ is also constant along the base.
+Now, consider the components along $e^{1}\otimes e^{1}$ and $e^{2}\otimes e^{2}$. These yield:
+$h_{+}\omega_{13}(e_{1})+h_{-}\omega_{14}(e_{1})=0,\quad h_{+}\omega_{23}(e_{2})+h_{-}\omega_{24}(e_{2})=0.$
+Using the expressions for $\omega_{ab}$, we obtain:
+$h_{+}\cdot\frac{1}{2}c^{(+)}+h_{-}\cdot\frac{1}{2}c^{(-)}=0,$
+which simplifies to
+$h_{+}c^{(+)}+h_{-}c^{(-)}=0\quad\text{or}\quad h_{+}+\lambda h_{-}=0.$
+Finally, the vertical derivatives (components along $e^{3}\otimes e^{3}$, etc.) imply that $h_{\pm}$ are also constant along the fibers.
+Hence, $h_{+}$ and $h_{-}$ are global constants satisfying (A.1).
+Therefore, the space of vertical parallel 1-forms is at most one-dimensional, spanned by $\lambda\Theta^{(+)}+\Theta^{(-)}$.
+However, when expressed in terms of the basis $\{d\phi_{+},d\phi_{-}\}$ of $H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, a vertical parallel 1-form corresponds to a linear combination $c_{+}d\phi_{+}+c_{-}d\phi_{-}$ such that the constants satisfy (A.1).
+This completes the derivation of condition ()(*) in the proof of Lemma 2.5.
+Appendix B Computation of the Obstruction Map $\Psi_{p}$
+This appendix provides the detailed computation of the obstruction map $\Psi_{p}$ defined in Section 4. We work under Assumption 2.1 and use the explicit harmonic representative from Lemma 3.1.
+Geometric Setup
+Recall the harmonic representative of the torsion class:
+$\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)}),\quad a_{-}=\lambda a_{+},\quad\lambda=c^{(-)}/c^{(+)}.$
+The subspace $\mathcal{V}_{2,1}$ is spanned by the simple tensor
+$\xi_{0}=\operatorname{vol}_{\mathrm{BZ}}\otimes(d\phi_{+}+\lambda d\phi_{-}).$
+Any element of $\mathcal{V}_{2,1}$ can be written as $\xi=t\xi_{0}$ for some $t\in\mathbb{R}$, which corresponds to taking $a_{+}=t$ and $a_{-}=\lambda t$.
+Definition of $\Psi_{p}$
+At a point $p=(k,\phi_{+},\phi_{-})\in M$, choose horizontal vectors $X,Y\in\mathcal{H}_{p}$ such that $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$.
+Choose a vertical vector $Z\in\mathcal{V}_{p}$ such that $(d\phi_{+}+\lambda d\phi_{-})(Z)=1$.
+For $\xi=\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\in\mathcal{V}_{2,1}$, define
+$\Psi_{p}(\xi)=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z),$
+where $(\cdot)_{\mathrm{off}}$ denotes the off-diagonal part (mapping horizontal to vertical or vice versa). As explained in the main text, the map is well-defined and linear.
+Key Computation
+We need to compute $(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z)$.
+Note that the torsion 3-form $T$ is cohomologous to $\omega_{h}$, but $T$ itself is not necessarily harmonic.
+However, the definition of $\Psi_{p}$ in the original PT theorem uses the harmonic representative to construct the map.
+Therefore, we compute using $\omega_{h}$ as the representative of the cohomology class.
+Consider the covariant derivative of $\omega_{h}$:
+$\nabla^{LC}\omega_{h}=\nabla^{LC}\bigl(\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)})\bigr).$
+Since $\operatorname{vol}_{\mathrm{BZ}}$ is parallel on the flat base, we have
+$\nabla^{LC}\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\bigl(a_{+}\nabla^{LC}\Theta^{(+)}+a_{-}\nabla^{LC}\Theta^{(-)}\bigr).$
+From the connection 1-forms computed in Appendix A, we have
+$\nabla^{LC}\Theta^{(+)}=\nabla^{LC}e^{3}=\frac{1}{2}c^{(+)}(e^{2}\otimes e^{1}-e^{1}\otimes e^{2}),$
+and similarly for $\Theta^{(-)}$ with $c^{(+)}$ replaced by $c^{(-)}$. In coordinate-free notation, this can be written as
+$\nabla^{LC}\Theta^{(\pm)}=\frac{1}{2}\pi^{*}F^{(\pm)}.$
+Thus,
+$\nabla^{LC}\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\frac{1}{2}\bigl(a_{+}\pi^{*}F^{(+)}+a_{-}\pi^{*}F^{(-)}\bigr).$
+Now, evaluate this 3-form on the vectors $X,Y,Z$. Since $X,Y$ are horizontal and $Z$ is vertical, we have
+$(\nabla^{LC}\omega_{h})(X,Y,Z)=\frac{1}{2}\bigl(a_{+}F^{(+)}(\pi_{*}X,\pi_{*}Y)+a_{-}F^{(-)}(\pi_{*}X,\pi_{*}Y)\bigr)\cdot\operatorname{vol}_{\mathrm{BZ}}(X,Y)\cdot 1,$
+where we used that $\pi^{*}F^{(\pm)}(X,Y)=F^{(\pm)}(\pi_{*}X,\pi_{*}Y)$ and that $Z$ is vertical so the contraction with the vertical part is trivial.
+By construction, $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$.
+Moreover, since $F^{(\pm)}=c^{(\pm)}\operatorname{vol}_{\mathrm{BZ}}$, we have
+$F^{(\pm)}(\pi_{*}X,\pi_{*}Y)=c^{(\pm)}.$
+Therefore,
+$(\nabla^{LC}\omega_{h})(X,Y,Z)=\frac{1}{2}(a_{+}c^{(+)}+a_{-}c^{(-)}).$
+Off-Diagonal Operator
+The off-diagonal part of the curvature operator is then given by raising the index of this 3-form.
+Specifically, there exists a fixed off-diagonal operator $J:T_{p}M\to T_{p}M$ that exchanges horizontal and vertical directions such that
+$(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z)=(a_{+}c^{(+)}+a_{-}c^{(-)})\,J.$
+The precise form of $J$ is not needed for the rank computation; it suffices that $J$ is non-zero and off-diagonal.
+Since $a_{-}=\lambda a_{+}$ and $c^{(-)}=\lambda c^{(+)}$, we have
+$a_{+}c^{(+)}+a_{-}c^{(-)}=a_{+}c^{(+)}(1+\lambda^{2}).$
+Thus,
+$\Psi_{p}(\xi)=\frac{1}{2}a_{+}c^{(+)}(1+\lambda^{2})J.$
+Kernel and Image
+The expression above shows that $\Psi_{p}(\xi)=0$ if and only if $a_{+}=0$, which implies $a_{-}=0$.
+Hence, $\ker\Psi_{p}=\{0\}$.
+Moreover, the image of $\Psi_{p}$ is one-dimensional, spanned by the off-diagonal operator $J$.
+This confirms the claim made in Section 4 that $\dim\operatorname{im}\Psi_{p}=1$.